### v1

# Improving Video Generation with Human Feedback

Abstract
Video generation has achieved significant advances through rectified flow techniques, but issues like unsmooth motion and misalignment between videos and prompts persist. In this work, we develop a systematic pipeline that harnesses human feedback to mitigate these problems and refine the video generation model. Specifically, we begin by constructing a large-scale human preference dataset focused on modern video generation models, incorporating pairwise annotations across multi-dimensions. We then introduce VideoReward, a multi-dimensional video reward model, and examine how annotations and various design choices impact its rewarding efficacy. From a unified reinforcement learning perspective aimed at maximizing reward with KL regularization, we introduce three alignment algorithms for flow-based models by extending those from diffusion models. These include two training-time strategies: direct preference optimization for flow (Flow-DPO) and reward weighted regression for flow (Flow-RWR), and an inference-time technique, Flow-NRG, which applies reward guidance directly to noisy videos. Experimental results indicate that VideoReward significantly outperforms existing reward models, and Flow-DPO demonstrates superior performance compared to both Flow-RWR and standard supervised fine-tuning methods. Additionally, Flow-NRG lets users assign custom weights to multiple objectives during inference, meeting personalized video quality needs. Project page: https://gongyeliu.github.io/videoalign.
Machine Learning, ICML
[…]
1 Introduction
Recent advances in video generation have yielded powerful models (Kong et al., 2024; Polyak et al., 2024; Kuaishou, 2024; Brooks et al., 2024), capable of producing videos with convincing details and relatively coherent motion. Despite these notable achievements, current video generation systems still face persistent challenges, including unstable motion, imperfect alignment with user prompts, unsatisfactory visual quality, and insufficient adherence to human preferences (Zeng et al., 2024). In language modeling and image generation, learning from human preference (Ouyang et al., 2022; Zhou et al., 2023; Wallace et al., 2024) has proved highly effective in improving the response quality and aligning generative models with user expectations.
However, applying such preference-driven alignment strategies to video generation remains in its infancy. One key obstacle is the lack of large-scale, high-quality preference data. A pioneering work (He et al., 2024) introduces a human-rated video preference dataset, and concurrent studies (Wang et al., 2024b; Liu et al., 2024b; Xu et al., 2024a) have contributed additional annotations. Yet, these datasets primarily focus on videos generated by earlier open-source models, which are often of relatively low quality, contain artifacts, or remain limited in duration. With video generation techniques rapidly advancing, the gap between such datasets and the capabilities of state-of-the-art video diffusion models (VDMs) has become increasingly pronounced.
A second challenge arises from the internal mechanisms of cutting-edge video generation models. Many modern systems employ rectified flow (Liu et al., 2022; Lipman et al., 2022), predicting velocity rather than noise. Recent studies (Wang et al., 2024b; Zhang et al., 2024a; Liu et al., 2024b; Xu et al., 2024a) have tested DPO (Rafailov et al., 2024; Wallace et al., 2024) and RWR (Peng et al., 2019; Lee et al., 2023; Furuta et al., 2024) on diffusion-based video generation approaches. However, adapting existing alignment methods to these flow-based models introduces new questions. While concurrent work (Domingo-Enrich et al., 2024) extended Diffusion-DPO to flow-based image generation, the aligned flow-based models in these explorations underperformed their unaligned counterparts.
To address these limitations, we present a comprehensive investigation into aligning advanced flow-based video generation models with human preferences. First, we select 12 advanced video generation models and curate a new large-scale multi-dimensional preference dataset of approximately 182k annotated examples. The dataset encompasses three critical dimensions—Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA)—to better capture how users evaluate generated videos. We then developed a multi-dimensional video reward model, systematically examining how annotations and different design choices affect its performance. Subsequently, we aggregated these multi-dimensional rewards into a comprehensive metric to capture overall human preferences. Furthermore, from a unified reinforcement learning perspective that maximizes reward with KL regularization, we introduce three alignment algorithms for flow-based models by extending those from diffusion models. These include two training-time strategies—Direct Preference Optimization for Flow (Flow-DPO) and Reward Weighted Regression for Flow (Flow-RWR)—and an inference-time technique, Flow-NRG. Our derivation reveals that the parameter $\beta$ (which modulates the KL divergence constraint) in Flow-DPO is dependent on the time step, resulting in Flow-DPO performing poorly on certain tasks. However, we found that by simply fixing $\beta$, Flow-DPO becomes the most effective alignment approach in our evaluations. Flow-NRG is an efficient inference-time alignment algorithm that applies reward guidance directly to noisy videos during the denoising process.It allows users to apply arbitrary weightings to multiple alignment objectives during inference, eliminating the need for extensive retraining to meet personalized user requirements.
Figure 1: Overview of Our Video Alignment Paradigm. (a) Human Preference Annotation (Sec. 3.1). We construct a dataset of 182k (prompt, video A, video B) triplets, collecting preference annotations on Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA) from human evaluators. (b) Reward Mode Training (Sec. 3.2). We train a VLM-based reward model using the Bradley-Terry-Model-with-Ties formulation. (c) Video Alignment (Sec. 4). We adapt alignment techniques — DPO, RWR, and reward guidance — to flow-based video generation models and provide a comprehensive comparison of their effectiveness.
Our contributions can be summarized as follows:
  • Large-Scale Preference Dataset: We create a new 182k-sized human-labeled video generation preference dataset from 12 video generation models, capturing VQ, MQ, and TA dimensions.
  • Multi-Dimensional Reward Modeling: We propose and systematically study a multi-dimensional video reward model, investigating how different design decisions influence its rewarding efficacy.
  • VideoGen-RewardBench: To more accurately assess video reward models, we have annotated prompt-video pairs from VideoGen-Eval. This dataset includes a wide variety of prompts and videos generated by modern VDMs. The resulting benchmark consists of 26.5k video pairs, each accompanied by a preference label.
  • Flow-Based Alignment: From a unified reinforcement learning perspective, we introduce three alignment algorithms for flow-based models: two training-time strategies (Flow-DPO, Flow-RWR) and one inference-time technique (Flow-NRG). Our experiments show that Flow-DPO outperforms other methods when $\beta$ is fixed. Additionally, Flow-NRG, using a lightweight reward model for noisy videos, enables users to apply custom weightings to multiple alignment objectives during inference.
2 Related Works
Evaluation and Reward Models.
Evaluation models and reward models play a pivotal role in aligning generative models with human preferences. Earlier approaches and benchmarks (Huang et al., 2023; Liu et al., 2024c; Huang et al., 2024) relied on metrics like FID (Heusel et al., 2017) and CLIP scores (Radford et al., 2021) to assess visual quality and semantic consistency. Recent works (Wu et al., 2023a; Xu et al., 2024b; Kirstain et al., 2023; Zhang et al., 2024b; Liang et al., 2024a) have shifted towards utilizing human preference datasets to train CLIP-based models, enabling them to predict preference scores with improved accuracy. With the advent of large vision-language models (VLMs) (Achiam et al., 2023; Wang et al., 2024a), their powerful capabilities in visual understanding and text-visual alignment make them a natural proxy for reward modeling. A common approach involves replacing the token classification head of VLMs with a regression head that predicts multi-dimensional scores for diverse evaluation tasks.
Two main learning paradigms have emerged based on the type of human annotation. The first paradigm relies on point-wise regression, where the model learns to fit annotated scores (He et al., 2024) or labels (Xu et al., 2024a) directly. Another paradigm focuses on pair-wise comparisons, leveraging Bradley-Terry (BT) (Bradley & Terry, 1952) loss to model relative preferences, which is largely unexplored for video reward model. Beyond these methods, some works (Wang et al., 2024b) also leverage the intrinsic reasoning capabilities of VLMs through VLM-as-a-judge (Li et al., 2023; Lin et al., 2024b), where VLMs are adopted to generate preference judgments or scores in textual format through instruction tuning. Despite these promising advances, most existing video reward models primarily focus on legacy video generation models, typically from the pre-Sora (OpenAI, 2024) era, which are constrained by short video durations and relatively low quality. Furthermore, the design and technical choices underlying the vision reward models remain underexplored. Our work seeks to address these limitations by focusing more on modern video generation models and investigating a broader range of reward modeling strategies.
Alignment for Image & Video Generation.
In large language models, Reinforcement Learning from Human Feedback (RLHF) improves alignment with human preferences, enhancing response quality (Ouyang et al., 2022; Jaech et al., 2024). Similar methods have been applied to image generation, using reward models or human preference data to align pretrained models. Key approaches include: (1) direct backpropagation with reward signals (Prabhudesai et al., 2023; Clark et al., 2023; Xu et al., 2024b; Prabhudesai et al., 2024); (2) Reward-Weighted Regression (RWR) (Peng et al., 2019; Lee et al., 2023; Furuta et al., 2024); (3) DPO-based policy optimization (Rafailov et al., 2024; Wallace et al., 2024; Dong et al., 2023; Yang et al., 2024a; Liang et al., 2024b; Yuan et al., 2024; Liu et al., 2024b; Zhang et al., 2024a; Furuta et al., 2024); (4) PPO-based policy gradients (Schulman et al., 2017) (Black et al., 2023; Fan et al., 2024); and (5) training-free alignment (Yeh et al., 2024; Tang et al., 2024; Song et al., 2023). These methods have successfully aligned image generation models with human preferences, improving aesthetics and semantic consistency. Our work applies the DPO algorithm (Rafailov et al., 2024; Wallace et al., 2024) to flow matching in video generation. Concurrent work (Domingo-Enrich et al., 2024) also explores similar things in image generation. However, they reports worse performance for the DPO-aligned model compared to the unaligned one. We argue that the originally derived Flow-DPO algorithm imposes a stronger KL constraint at lower noise steps, resulting in suboptimal performance. In contrast, using a constant KL constraint significantly improves performance on certain tasks.
There has also been work on aligning video generation models using feedback (Furuta et al., 2024; Li et al., 2024). However, these approaches rely on image reward models (Kirstain et al., 2023; Wu et al., 2023a). Concurrent works (Yuan et al., 2024; Liu et al., 2024b; Zhang et al., 2024a) focus on aligning diffusion-based video generation models by annotating videos generated from earlier open-source models. Our work differs by exploring alignment techniques for advanced flow-based video generation.
3 VideoReward
3.1 Human Preference Data Collection
Table 1: Statistics of the collected training dataset. We utilized 12 text-to-video models to generate a total of 108k videos from 16k unique prompts. This process ultimately resulted in 182k annotated triplets, each consisting of a prompt paired with two videos and corresponding preference annotations.
T2V Model Date #Videos #Anno Triplets Resolution Duration
Pre-Sora-Era Models Gen2 (Runway, 2023) 23.06 6k 13k 768 $\times$ 1408 4s
SVD (Blattmann et al., 2023) 23.11 6k 13k 576 $\times$ 1024 4s
Pika 1.0 (Labs, 2023) 23.12 6k 13k 720 $\times$ 1280 3s
Vega (VegaAI, 2023) 23.12 6k 13k 576 $\times$ 1024 4s
Pixverse v1 (PixVerse, 2024) 24.01 6k 13k 768 $\times$ 1408 4s
HiDream (HidreamAI, 2024) 24.01 0.3k 0.3k 768 $\times$ 1344 5s
Modern Models Dreamina (Capcut, 2024) 24.03 16k 68k 720 $\times$ 1280 6s
Luma (LumaLabs, 2024) 24.06 16k 57k 752 $\times$ 1360 5s
Gen3 (Runway, 2024) 24.06 16k 55k 768 $\times$ 1280 5s
Kling 1.0 (Kuaishou, 2024) 24.06 6k 33k 384 $\times$ 672 5s
Pixverse v2 (PixVerse, 2024) 24.07 16k 58k 576 $\times$ 1024 5s
Kling 1.5 (Kuaishou, 2024) 24.09 7k 28k 704 $\times$ 1280 5s
Figure 2: Statistics of our training data.
Existing human preference datasets for video generation (Liu et al., 2024c; Huang et al., 2024; He et al., 2024; Wang et al., 2024b; Xu et al., 2024a) primarily consist of videos generated by earlier open-source models, which are often characterized by low quality, noticeable artifacts, and short video durations. As advancements in video diffusion models (VDMs) continue to elevate the state-of-the-art, the gap between current preference datasets and the capabilities of modern VDMs has become increasingly pronounced. To bridge this gap, we focus on developing a reward model tailored for the latest VDMs. Inspired by Zhang et al. (2024b), we collect a diverse set of prompts from the internet, organizing them into 8 meta-categories (animal, architecture, food, people, plants, scenes, vehicles, objects). These prompts are further expanded using GPT-4o (Achiam et al., 2023). A filtering process is then applied to exclude unsuitable prompts (e.g., repetitive, irrelevant to motion, unsafe content). Subsequently, we employ our internal prompt rewriter to refine and enhance the collected prompts, resulting in a final dataset of 16k detailed prompts. Leveraging these prompts, we select 12 text-to-video generation models with various capabilities, producing 108k videos and constructing 182k triplets, each comprising a prompt and two corresponding videos generated by different video generation models. Comprehensive statistics of the dataset are provided in Tab. 1 and Fig 2.
To annotate these text-video triplets, we hire annotators to perform pairwise assessments across three critical dimensions: Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA). Details are available in appendix C. For each dimension, annotators are presented with a prompt and two generated videos, and asked to provide choices indicating their preference (A wins/Ties/B wins). Each sample is independently assigned to three annotators to ensure robustness. In cases of disagreement, an additional annotator will be engaged to resolve conflicts. This effort results in a dataset of 182k high-quality, multi-dimensional preference annotations, which were utilized for training the reward model. In addition to pairwise annotations, we also conduct pointwise annotations using a similar procedure, where annotators rate each video on a Likert scale from 1 to 5 for the same three dimensions (i.e., VQ, MQ, TA). This dual annotation setup enables us to explore the advantages and limitations of pairwise preferences versus pointwise scores as annotation strategies.
In addition to the training data, we construct a validation set comprising 13k annotated triplets, following a similar annotation protocol. The prompts in the validation set are strictly excluded from the training set to ensure generalization. The validation set serves as a reliable benchmark to monitor the convergence during training.
3.2 Reward Model Learning
Following prior works that leverage Vision-Laguage Models (VLMs) for related tasks (He et al., 2024; Wang et al., 2024b; Xu et al., 2024a), we adopt Qwen2-VL-2B (Wang et al., 2024a) as the base model for our video reward framework. While existing studies have demonstrated the effective application of reward models in both evaluation (Wu et al., 2023b, a; He et al., 2024) and optimization (Lee et al., 2023; Wallace et al., 2024; Xu et al., 2024b; Prabhudesai et al., 2024), the underlying design choices remain insufficiently explored. We start with a Bradley-Terry style reward model as our baseline, exploring how technical decisions affect the video reward model’s performance.
Score Regression v.s. Bradley-Terry.
We first explore the merits of adopting a Bradley-Terry pairwise comparison approach compared to a pointwise score regression approach for training reward models. The BT model (Bradley & Terry, 1952) uses a pairwise log-likelihood loss to capture the probability that one video is preferred over another, enabling the prediction of rewards based on the relative preferences. Conversely, the pointwise regression model directly estimates the absolute quality scores by minimizing the MSE loss. The loss functions for these methodologies are formally defined as follows:
$\mathcal{L}_{BT}=-\mathbb{E}_{({\bm{y}},{\bm{x}}^{w}_{0},{\bm{x}}^{l}_{0})\sim\mathcal{D}}\left[\log\left(\sigma\left(r({\bm{x}}^{w}_{0},{\bm{y}})-r({\bm{x}}^{l}_{0},{\bm{y}})\right)\right)\right]$ (1)
$\mathcal{L}_{reg}=\mathbb{E}_{({\bm{y}},{\bm{x}}_{0},z)\sim\mathcal{D}}\left[\|r({\bm{x}}_{0},{\bm{y}})-z\|^{2}\right]$ (2)
In these formulations, $({\bm{y}},{\bm{x}}^{w}_{0},{\bm{x}}^{l}_{0})$ denotes pairwise annotations, where ${\bm{y}}$ is the input prompt and ${\bm{x}}^{w}_{0}$ and ${\bm{x}}^{l}_{0}$ are the two generated videos with corresponding preference annotations. Meanwhile, $({\bm{y}},{\bm{x}}_{0},z)$ represents pointwise annotations, with $z$ being the assigned quality score. The function $r$ signifies the optimized reward function.
Given that our training dataset includes both pointwise scores and pairwise preferences, annotated by the same group of annotators on the same dataset, this setup enables a direct comparison between the two paradigms. To conduct this comparison, we train both types of reward models on the training data with varying proportions of annotated data. Their performance is assessed by evaluating the average accuracy of the best-performing checkpoints on the validation set. Notably, for the BT reward model, we exclude samples labeled as ”Ties” and only utilize samples annotated as ”A Wins” or ”B Wins” since the BT framework is not well-suited to handle tied data. The comparison results are illustrated in Fig. 3.
Figure 3: Accuracy comparison between the BT and regression reward models across varying training data fractions (log scale).
We observe that as the dataset size increases, both the Bradley-Terry reward model and the regression reward model exhibit notable improvements in accuracy, with the BT model consistently outperforming the regression model. This disparity can be attributed to the inherent characteristics of pointwise and pairwise annotations. Pairwise annotations are inherently more effective at capturing relative differences between samples. For instance, even when two videos receive the same score in pointwise annotations, annotators can still discern their relative quality in pairwise comparisons. This ability grants the BT reward model a distinct advantage when the dataset size is limited. As the dataset expands, the performance gap between the two models narrows, as the regression model benefits from the increased data to learn more accurate ranking relationships. Additionally, the annotation process itself varies in complexity between the two approaches. Pairwise preference labels are generally easier for annotators to agree upon compared to assigning precise scores in pointwise annotations. Recent studies (Wang et al., 2024d, c) have demonstrated that with carefully curated pointwise annotations, regression models can achieve performance levels comparable to BT reward models. We posit that as both the quantity and quality of the data improve, the performance differences between these approaches become negligible. While in the context of our current training data, the BT reward model emerges as the better-performing choice.
Ties Matters.
While the BT model is commonly used to capture human preferences from chosen-rejected pairs, the importance of tie annotations is often overlooked. Inspired by recent works (Liu et al., 2024a), we use the Bradley-Terry model with ties (BTT) (Rao & Kupper, 1967), an extension of BT that accounts for tied preferences. For a given prompt ${\bm{y}}$ and two generated videos ${\bm{x}}^{1}_{0}$ and ${\bm{x}}^{2}_{0}$, the BTT model defines the probabilities of each possible preference as follows:
$P({\bm{x}}_{0}^{A}={\bm{x}}_{0}^{B}|{\bm{y}})$ $=\frac{(\theta^{2}-1)e^{r({\bm{x}}_{0}^{A},{\bm{y}})}e^{r({\bm{x}}_{0}^{B},{\bm{y}})}}{\left(e^{r({\bm{x}}_{0}^{A},{\bm{y}})}+\theta e^{r({\bm{x}}_{0}^{B},{\bm{y}})}\right)\left(\theta e^{r({\bm{x}}_{0}^{A},{\bm{y}})}+e^{r({\bm{x}}_{0}^{B},{\bm{y}})}\right)}$ (3)
$P({\bm{x}}_{0}^{A}\succ{\bm{x}}_{0}^{B}|{\bm{y}})$ $=\frac{e^{r({\bm{x}}_{0}^{A},{\bm{y}})}}{e^{r({\bm{x}}_{0}^{B},{\bm{y}})}+\theta e^{r({\bm{x}}_{0}^{A},{\bm{y}})}}$ (4)
$P({\bm{x}}_{0}^{B}\succ{\bm{x}}_{0}^{A}|{\bm{y}})$ $=\frac{e^{r({\bm{x}}_{0}^{B},{\bm{y}})}}{\theta e^{r({\bm{x}}_{0}^{A},{\bm{y}})}+e^{r({\bm{x}}_{0}^{B},{\bm{y}})}}$ (5)
where $\theta$ is a parameter that controls the tendency toward ties, with a larger $\theta$ indicating a higher probability of ties. We empirically set $\theta=5.0$ and optimize the BTT model by minimizing the negative log-likelihood (NLL):
$\mathcal{L}_{\text{BTT}}=-\mathbb{E}_{({\bm{y}},{\bm{x}}_{0}^{A},{\bm{x}}_{0}^{B})\sim\mathcal{D}}\Big{[}$ $\mathbbm{1}({\bm{x}}_{0}^{A}\succ{\bm{x}}_{0}^{B})\log P({\bm{x}}_{0}^{A}\succ{\bm{x}}_{0}^{B}|{\bm{y}})+$ (6)
$\mathbbm{1}({\bm{x}}_{0}^{B}\succ{\bm{x}}_{0}^{A})\log P({\bm{x}}_{0}^{B}\succ{\bm{x}}_{0}^{A}|{\bm{y}})+$
$\mathbbm{1}({\bm{x}}_{0}^{A}={\bm{x}}_{0}^{B})\log P({\bm{x}}_{0}^{A}={\bm{x}}_{0}^{B}|{\bm{y}})\Big{]}$
We train both the Bradley-Terry reward model and the Bradley-Terry-With-Tied reward model under the same setting and visualize the distribution of $\Delta r=r({\bm{x}}_{0}^{A},{\bm{y}})-r({\bm{x}}_{0}^{B},{\bm{y}})$ on the validation set using the best-performing checkpoints. As shown in the boxplot in Fig. 4, while the BT model effectively manages chosen/rejected pairs, it struggles to accurately differentiate tied pair. A large portion of ties exhibit a larger $\Delta r$ than certain chosen/rejected pairs, underscoring the BT model’s limitations in handling ambiguous cases. In contrast, the BTT model learns a more generalized decision boundary, enhancing it to distinguish tied pairs from clear chosen/rejected pairs based on $\Delta r$. This improvement stems from the explicit modeling of ties in Eq. 3, which allows the model to incorporate this often-overlooked category, resulting in more robust performance.
Figure 4: Visualization of the $\Delta r$ distribution for the BT reward model (Left) and the BTT reward model (Right). The BTT model effectively distinguishes tie pairs from chosen/rejected pairs.
Separation for Decoupling.
When adopting LLMs/MLLMs for multi-attribute reward modeling, a common practice (Ouyang et al., 2022; Touvron et al., 2023; He et al., 2024) involves extracting the last token from the final layer and applying a trainable linear projection to predict scores across multiple dimensions. While this approach enables efficient multi-dimensional scoring, it inherently induces coupling among dimensions, particularly between context-agnostic and context-aware attributes. For example, such a model may inadvertently rely on the input prompt when evaluating visual quality, causing the same video sample to receive different visual quality scores depending on its prompt. Although this coupling may be inherited from the correlations within the annotated data, it compromises the interpretability of reward scores and introduces challenges for the evaluation or optimization of individual dimensions.
To decouple context-agnostic dimensions from contextual influences, we introduce separate special tokens for each dimension. Specially, three unique query tokens are added to represent the scores across the three evaluation dimensions. In the input template of our reward models, the context-agnostic tokens (i.e., VQ, MQ) are placed after the video but before the prompt, while the context-aware special tokens (i.e., TA) are positioned after the prompt. We then extract the feature representations from the final layer corresponding to each special token’s position and apply a shared linear head to predict the associated dimension scores. Owing to the autoregressive model’s causal attention mechanism, the context-agnostic tokens exclusively attend to the video content, whereas the context-aware tokens attend to both the video and the prompt. This strategy effectively separates the focus of each token, enabling a more robust performance of each dimension and independent evaluation for context-agnostic dimensions. We provide a full version of our input prompt template in appendix B.
4 Video Alignment
Let ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$ denote a real data sample drawn from the true data distribution, and let ${\bm{x}}_{1}\sim p({\bm{x}}_{1})$ denote a noise sample, where ${\bm{x}}_{0},{\bm{x}}_{1}\in\mathbb{R}^{d}$. Recent advanced image-generation models (e.g., (Esser et al., 2024)) and video-generation models (e.g., (Kong et al., 2024)) adopt the Rectified Flow framework (Liu et al., 2022), which defines the “noised” data ${\bm{x}}_{t}$ as
${\bm{x}}_{t}=(1-t)\,{\bm{x}}_{0}\;+\;t\,{\bm{x}}_{1},$
for $t\in[0,1]$. Then we can train a transformer model to directly regress the velocity field ${\bm{v}}_{\theta}({\bm{x}}_{t},t)$ by minimizing the Flow Matching objective (Lipman et al., 2022; Liu et al., 2022):
$\mathcal{L}(\theta)=\mathbb{E}_{t,\;{\bm{x}}_{0}\sim q({\bm{x}}_{0}),\;{\bm{x}}_{1}\sim p({\bm{x}}_{1})}\bigl{[}\;\|{\bm{v}}\;-\;{\bm{v}}_{\theta}({\bm{x}}_{t},t)\|^{2}\bigr{]},$
where the target velocity field is ${\bm{v}}={\bm{x}}_{1}-{\bm{x}}_{0}$.
4.1 Flow-DPO
Consider a fixed dataset $\mathcal{D}=\{{\bm{y}},{\bm{x}}_{0}^{w},{\bm{x}}_{0}^{l}\}$, where each sample consists of a prompt ${\bm{y}}$ and two videos ${\bm{x}}_{0}^{w}$ and ${\bm{x}}_{0}^{l}$ generated by a reference model $p_{\text{ref}}$, with human annotations indicating that ${\bm{x}}_{0}^{w}$ is preferred over ${\bm{x}}_{0}^{l}$ (i.e., ${\bm{x}}_{0}^{w}\succ{\bm{x}}_{0}^{l}$). The goal of RLHF is to learn a conditional distribution $p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})$ that maximizes a reward model $r({\bm{x}}_{0},{\bm{y}})$, while controlling the regularization term (KL-divergence) from the reference model $p_{\text{ref}}$ via a coefficient $\beta$:
$\max_{p_{\theta}}$ $\mathbb{E}_{{\bm{y}}\sim\mathcal{D}_{c},{\bm{x}}_{0}\sim p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})}\left[r({\bm{x}}_{0},{\bm{y}})\right]$
$-\beta\,\mathbb{D}_{\text{KL}}\left[p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})\,\|\,p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})\right]$ (7)
In this context, $p_{\text{ref}}$ corresponds to $p_{\text{ref},0}$ and $p_{\theta}$ corresponds to $p_{\theta,0}$. For the sake of simplicity, the timestep subscripts are omitted. However, RLHF is usually unstable and resource intensive (Rafailov et al., 2024). To address this issue, Diffusion-DPO aligns diffusion models with human preferences by directly solving the above RLHF objective (Eq. 4.1) analytically. It interprets alignment as a classification problem, and optimizes a policy to satisfy human preferences by supervised training. For simplicity, we omit the conditioning prompt ${\bm{y}}$ in the following equations. The Diffusion-DPO objective $\mathcal{L}_{\text{DD}}(\theta)$ is given by:
$-\mathbb{E}\Bigg{[}\log\sigma\Bigg{(}$ $-\frac{\beta}{2}\Big{(}\|{\bm{\epsilon}}^{w}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{w},t)\|^{2}-\|{\bm{\epsilon}}^{w}-{\bm{\epsilon}}_{\text{ref}}({\bm{x}}_{t}^{w},t)\|^{2}$
$\quad-\bigl{(}\|{\bm{\epsilon}}^{l}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{l},t)\|^{2}-\|{\bm{\epsilon}}^{l}-{\bm{\epsilon}}_{\text{ref}}({\bm{x}}_{t}^{l},t)\|^{2}\bigr{)}\Big{)}\Bigg{)}\Bigg{]}$ (8)
where ${\bm{x}}_{t}^{*}=(1-t)\ {\bm{x}}_{0}^{*}+t\ {\bm{\epsilon}}$, ${\bm{\epsilon}}\sim\mathcal{N}(0,{\bm{I}})$. The superscript “*” indicates either “$w$” (for the preferred data) or “$l$” (for the less preferred data). The expectation is taken over samples $\{{\bm{x}}_{0}^{w},{\bm{x}}_{0}^{l}\}\sim\mathcal{D}$ and the noise schedule $t$. In Rectified Flow, we relate the noise vector ${\bm{\epsilon}}^{*}$ to a velocity field ${\bm{v}}^{*}$. Specifically, Lemma A.1 in Appendix A shows that
$\|{\bm{\epsilon}}^{*}-{\bm{\epsilon}}_{\text{pred}}({\bm{x}}_{t}^{*},t)\|^{2}=(1-t)^{2}\|{\bm{v}}^{*}-{\bm{v}}_{\text{pred}}(x_{t}^{*},t)\|^{2},$ (9)
where ${\bm{\epsilon}}_{\text{pred}}$ and ${\bm{v}}_{\text{pred}}$ refer to predictions either from the model $p_{\theta}$ or the reference model $p_{\text{ref}}$. By substituting Eq. 9 into Eq. 4.1, we obtain the final Flow-DPO loss $\mathcal{L}_{\text{FD}}(\theta)$:
$-\mathbb{E}\Bigg{[}\log\sigma\Bigg{(}$ $-\frac{\beta_{t}}{2}\Big{(}\|{\bm{v}}^{w}-{\bm{v}}_{\theta}({\bm{x}}_{t}^{w},t)\|^{2}-\|{\bm{v}}^{w}-{\bm{v}}_{\text{ref}}({\bm{x}}_{t}^{w},t)\|^{2}$
$\quad-\bigl{(}\|{\bm{v}}^{l}-{\bm{v}}_{\theta}({\bm{x}}_{t}^{l},t)\|^{2}-\|{\bm{v}}^{l}-{\bm{v}}_{\text{ref}}({\bm{x}}_{t}^{l},t)\|^{2}\bigr{)}\Big{)}\Bigg{)}\Bigg{]},$ (10)
where $\beta_{t}=\beta\,(1-t)^{2}$ and the expectation is taken over samples $\{{\bm{x}}_{0}^{w},{\bm{x}}_{0}^{l}\}\sim\mathcal{D}$ and the noise schedule $t$. Intuitively, minimizing $\mathcal{L}_{\text{FD}}(\theta)$ guides the predicted velocity field ${\bm{v}}_{\theta}$ closer to the target velocity ${\bm{v}}^{w}$ of the “preferred” data, while pushing it away from ${\bm{v}}^{l}$ (the “less preferred” data). The strength of this preference signal depends on the differences between the predicted errors and the corresponding reference errors, $\|{\bm{v}}^{w}-{\bm{v}}_{\text{ref}}({\bm{x}}_{t}^{w},t)\|^{2}$ and $\|{\bm{v}}^{l}-{\bm{v}}_{\text{ref}}({\bm{x}}_{t}^{l},t)\|^{2}$. We provide a pseudo-code of Flow-DPO in Appendix E.
Discussion on $\beta_{t}$.
The parameter $\beta_{t}$ governs the strength of the KL divergence constraint, determining the extent to which the model can diverge from the base reference model (Rafailov et al., 2024; Wallace et al., 2024). Note that from our derivation, $\beta_{t}=\beta(1-t)^{2}$. As the timestep $t$ approaches 1, $\beta_{t}$ approaches zero, resulting in a relatively weak KL constraint. Conversely, as $t$ approaches 0, $\beta_{t}$ converges to the original $\beta$, enforcing a stronger KL constraint. This scheduling strategy causes the model to prioritize alignment at higher noise levels. Our experiments reveal that this degrades performance, while using a constant $\beta_{t}$ provides better results. Similar phenomena have been observed in DDPM’s (Ho et al., 2020) standard diffusion training objective, where discarding the weighting in denoising score matching leads to improved sample quality, as discussed in Section 3.4 of their paper. We provide a more detailed discussion of this in Section 5.2.
4.2 Flow-RWR
Drawing inspiration from the application of Reward-weighted Regression (RWR) (Peters & Schaal, 2007) in diffusion models (Lee et al., 2023; Furuta et al., 2024), we propose a counterpart for flow-based models based on expectation-maximization. Starting from the general KL-regularized reward-maximization problem in Eq. 4.1, prior work (Rafailov et al., 2024; Peters & Schaal, 2007) shows that its optimal closed-form solution can be written as
$p_{\theta}(\mathbf{x}_{0}\mid\mathbf{y})=\frac{1}{Z(\mathbf{y})}\,p_{\text{ref}}(\mathbf{x}_{0}\mid\mathbf{y})\,\exp\biggl{(}\frac{1}{\beta}\,r(\mathbf{x}_{0},\mathbf{y})\biggr{)},$ (11)
where $Z(\mathbf{y})=\sum_{\mathbf{x}_{0}}p_{\text{ref}}(\mathbf{x}_{0}\mid\mathbf{y})\exp\bigl{(}\tfrac{1}{\beta}r(\mathbf{x}_{0},\mathbf{y})\bigr{)}$ is the partition function. Following (Furuta et al., 2024), we can obtains the RWR loss by M-step:
$\mathcal{L}_{\text{RWR}}(\theta)=-\mathbb{E}_{{\bm{y}}\sim\mathcal{D}_{c}}\bigl{[}\frac{1}{Z({\bm{y}})}\exp(\frac{r({\bm{x}}_{0},{\bm{y}})}{\beta})\log p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})\bigr{]}.$
In practice, we omit the intractable normalization $Z(\mathbf{y})$ and $\beta$, and we employ Eq. 37 in Wallace et al. (2024) as a simplified upper bound of the negative log-likelihood. This yields the approximate form
$\mathcal{L}_{\text{RWR}}(\theta)=\mathbb{E}_{{\bm{y}},{\bm{x}}_{0},{\bm{\epsilon}},t}\bigl{[}\exp(r({\bm{x}}_{0},{\bm{y}}))\|{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})\|^{2}\bigr{]}.$ (12)
Substituting Eq. 9 into Eq. 12 then provides the final RWR objective for velocity prediction:
$\mathcal{L}_{\text{RWR}}(\theta)$ $=\mathbb{E}\bigl{[}\exp(r({\bm{x}}_{0},{\bm{y}}))\|{\bm{v}}-{\bm{v}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})\|^{2}\bigr{]},$ (13)
where $\mathbf{v}$ is the ground-truth velocity, and $\mathbf{v}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})$ represents the predicted velocity under parameters $\theta$. We omit $(1-t)^{2}$ as in Sec. 4.1. By applying RWR in this way, our flow-based model learns to generate velocity fields that emphasize high-reward samples, paralleling how diffusion-based models align samples via reward-weighted likelihood maximization.
4.3 Noisy Reward Guidance
Recall that Eq. 4.1 has a closed-form solution as presented in Eq. 11. This expression indicates that the goal of RLHF is to transform the original distribution $p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})$ into the new target distribution $p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})$. Since the constants $\beta$ and $w$ can be absorbed into $r({\bm{x}}_{0},{\bm{y}})$, Eq. 11 becomes
$p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})\;\propto\;p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})\,[\exp(r({\bm{x}}_{0},{\bm{y}}))]^{w},$ (14)
where $w\in\mathbb{R}$ controls the strength of the reward guidance. Notably, this formulation aligns with the concept of classifier guidance (Dhariwal & Nichol, 2021; Song et al., 2020), so we refer to it as reward guidance.
As we proved in Appendix A.2, for Rectified Flow, if we set the marginal velocity field to
$\tilde{{\bm{v}}}_{t}({\bm{x}}_{t}\mid{\bm{y}})={\bm{v}}_{t}({\bm{x}}_{t}\mid{\bm{y}})-w\,\frac{t}{1-t}\,\nabla r({\bm{x}}_{t},{\bm{y}}),$ (15)
the marginal distribution is modified from $p_{t}({\bm{x}}_{t}\mid{\bm{y}})$ to
$\tilde{p}_{t}({\bm{x}}_{t}\mid{\bm{y}})\;\propto\;p_{t}({\bm{x}}_{t}\mid{\bm{y}})\,[\exp(r({\bm{x}}_{t},{\bm{y}}))]^{w}.$
When $t=0$, the resulting samples are drawn from Eq. 14. We provide a pseudo-code in Appendix E. To producing meaningful reward values for noised videos ${\bm{x}}_{t}$, we train a time-dependent reward function. We use the same dataset $D=\{{\bm{y}},{\bm{x}}^{w},{\bm{x}}^{l}\}$ described in Section 3.2, where ${\bm{x}}^{w}$ and ${\bm{x}}^{l}$ represent two videos with different preference levels. In order to train our reward function, we assume that applying the same noise to both videos preserves their preference relationship. Formally, for each pair $({\bm{x}}^{w},{\bm{x}}^{l})$, we define
${\bm{x}}_{t}^{w}=(1-t)\,{\bm{x}}_{0}+t\,{\bm{x}}_{1}^{w},$
${\bm{x}}_{t}^{l}=(1-t)\,{\bm{x}}_{0}+t\,{\bm{x}}_{1}^{l},\text{with}\,{\bm{x}}_{t}^{w}\succ{\bm{x}}_{t}^{l}.$ (16)
Here, ${\bm{x}}_{0}$ is the original video, while ${\bm{x}}_{1}^{w}$ and ${\bm{x}}_{1}^{l}$ are the noised versions corresponding to the more preferred and less preferred videos, respectively. We then adopt the BT model to learn the reward function from these noised videos. Specifically, we minimize the following loss:
$\mathcal{L}_{BT}(\theta)=-\mathbb{E}\bigl{[}\log\bigl{(}\sigma\bigl{(}r_{\theta}({\bm{x}}_{t}^{w},t)-r_{\theta}({\bm{x}}_{t}^{l},t)\bigr{)}\bigr{)}\bigr{]},$ (17)
Table 2: Preference accuracy on GenAI-Bench and VideoGen-Eval. w/ Ties indicates that accuracy is calculated with ties included (Deutsch et al., 2023), and w/o Ties excludes tied pairs when calculating accuracy. * denotes that for LiFT, ties prediction are randomly converted to chosen/rejected with a 0.5 probability due to a large number of ties produced by the model. Bold: best performance.
Method GenAI-Bench VideoGen-RewardBench
Overall Accuracy Overall Accuracy VQ Accuracy MQ Accuracy TA Accuracy
w/ Ties w/o Ties w/ Ties w/o Ties w/ Ties w/o Ties w/ Ties w/o Ties w/ Ties w/o Ties
Random 33.67 49.84 41.86 50.30 47.42 49.86 59.07 49.64 37.25 50.40
VideoScore (He et al., 2024) 49.03 71.69 41.80 50.22 47.41 47.72 59.05 51.09 37.24 50.34
LiFT* (Wang et al., 2024b) 37.06 58.39 39.08 57.26 47.53 55.97 59.04 54.91 33.79 55.43
VisionRewrd (Xu et al., 2024a) 51.56 72.41 56.77 67.59 47.43 59.03 59.03 60.98 46.56 61.15
Ours 49.41 72.89 61.26 73.59 59.68 75.66 66.03 74.70 53.80 72.20
Table 3: Ablation study on reward model type, seprate tokens and data augmentation. Bold: Best Performance.
Variants RM Type Separate Tokens GenAI-Bench VideoGen-RewardBench
Overall Accuracy Overall Accuracy VQ Accuracy MQ Accuracy TA Accuracy
w/ Ties w/o Ties w/ Ties w/o Ties w/ Ties w/o Ties w/ Ties w/o Ties w/ Ties w/o Ties
I Regression 48.28 71.13 58.39 70.16 54.23 73.61 61.16 65.56 52.60 70.95
II BT 47.74 71.21 61.22 72.58 52.33 77.10 59.43 73.50 53.06 71.62
III BTT 48.27 70.89 61.50 73.39 60.52 76.31 64.64 72.40 53.55 72.12
IV BTT 49.41 72.89 61.26 73.59 59.68 75.66 66.03 74.70 53.80 72.20
where the expectation is taken over $({\bm{x}}^{w},{\bm{x}}^{l},{\bm{y}})\sim\mathcal{D},\,t\sim U(0,1)$. In this way, the reward function $r_{\theta}(\cdot,t)$ learns to preserve the relative preference between noised videos at any noise level $t$. However, advanced video generation models typically work in a latent space and rely on a VAE (Kingma, 2013) to decode the latent representation back into a video. In Eq. 15, computing the gradient of the reward with respect to ${\bm{x}}_{t}$ would require first decoding the latent representation into a video and then passing the gradient through the large VAE decoder back to the latent space, incurring high computational cost and memory usage.
To address this challenge, we propose training the reward model directly within the latent space using a pretrained VDM. Since the VDM has already been trained on noisy latent representations, we can develop a lightweight, time-dependent reward function by utilizing only the initial layers of the pretrained model rather than the entire VDM. Subsequently, we apply Equation 15 at each inference step to implement reward-guided generation.
5 Experiments
In this section, we empirically evaluate our reward model and reinforcement learning alignment algorithms on Text-to-Video (T2V) tasks.
5.1 Reward Learning
Training Setting.
We utilize Qwen2-VL-2B (Wang et al., 2024a) as the backbone for our reward model and train it with BTT loss. To finetune the model, LoRA (Hu et al., 2021) is applied to update all linear layers in the language model, while the vision encoder’s parameters are fully optimized. The training process is conducted with a batch of 32 and a learning rate of $2\times 10^{-6}$, with the model trained over two epochs. This setup requires approximately 72 A800 GPU hours. Several observations were made during training. First, higher video resolution and more frames generally improved the reward model’s performance. Second, using a stable sampling interval instead of a fixed frame number significantly enhanced the accuracy of motion quality evaluations, particularly when trained with videos of varying durations. In practice, we sample videos at 2 fps, with a resolution of approximately $448\times 448$ pixels during the training process. The original aspect ratio is preserved to ensure visual quality.
Evaluation.
To assess the performance of our reward model in evaluating T2V generation, we employ a comprehensive evaluation framework designed to ensure a fair comparison across different models and mitigate potential domain biases. We utilize two primary benchmarks for evaluation. (1). VideoGen-RewardBench: Derived from the third-party prompt-video dataset VideoGen-Eval (Zeng et al., 2024), VideoGen-RewardBench is specifically designed to evaluate modern T2V models. To address the absence of human annotations in VideoGen-Eval, we manually construct 26.5k triplets and hire annotators to provide pairwise preference labels, following the procedure outlined in Sec. 3.1. Additionally, annotators assess the overall quality between the two videos, serving as a universal label for reward models across various dimensions. (2). GenAI-Bench (Jiang et al., 2024b): : Serving as a complementary benchmark, GenAI-Bench features short videos (2 seconds in duration) generated by pre-Sora-era T2V models. This benchmark enables us to evaluate our reward model’s capability in assessing earlier-generation models, despite their limited representation in our training set. We summarize the key difference between VideoGen-RewardBench and GenAI-Bench in Appendix D.1.
We compare our reward model against existing baselines, including VideoScore (He et al., 2024), as well as two concurrent works: LiFT (Wang et al., 2024b), and VisionReward (Xu et al., 2024a). Consistent with prevailing practices in LLMs (Lambert et al., 2024), we employ pairwise accuracy as a primary metric. Inspired by VisionReward (Xu et al., 2024a), we report two accuracy metrics: ties-included accuracy (Deutsch et al., 2023) and ties-excluded accuracy. We calculate the overall accuracy on GenAI-Bench and both overall and dimension-specific (VQ, MQ, TA) accuracy on VideoGen-RewardBench. Additional details of the evaluation across different methods are provided in the Appendix D.2.
Table 4: Multi-dimensional alignment with VQ:MQ:TA = 1:1:1. Bold: Best performance. Although Flow-DPO with a timestep-dependent $\beta$ achieves high VQ and MQ reward win rates, it exhibits significant reward hacking. In contrast, Flow-DPO with a constant $\beta$ achieves high VQ, MQ, and TA scores while avoiding reward hacking.
Method VBench VideoGen-Eval TA-Hard
Total Quality Sementic VQ MQ TA VQ MQ TA VQ MQ TA
Pretrained 83.19 84.37 78.46 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0
SFT 82.31 83.13 79.04 51.28 65.21 52.84 61.27 76.13 46.35 57.75 76.06 57.75
Flow-RWR 82.27 83.19 78.59 51.55 63.9 53.43 59.05 69.7 48.35 61.97 78.87 55.71
Flow-DPO ($\beta_{t}=\beta(1-t)^{2}$) 80.90 81.52 78.42 87.78 82.36 51.02 88.44 91.23 28.14 84.29 83.10 38.03
Flow-DPO 83.41 84.19 80.26 93.42 69.08 75.43 90.95 81.01 68.26 77.46 71.43 73.24
Main Results.
Tab 2 presents the pairwise accuracy of various reward models on GenAI-Bench (representing pre-Sora-era models) and VideoGen-RewardBench (representing modern T2V models). VideoScore is primarily designed to evaluate videos from earlier-generation models. It performs well on GenAI-Bench but exhibits close-to-random predictions on VideoGen-RewardBench, reflecting its inability to generalize to the improved capabilities of modern T2V models. LiFT adopts the VLM-as-a-judge approach to train its reward model, but it struggles with pairwise differentiation. This is because LiFT tends to assign similar scores to both videos in a pair. When calculating its ties-excluded accuracy, we randomly convert ties prediction to chosen/rejected categories with a 0.5 probability, indicating the model fail to distinguish between video qualities. VisionReward demonstrates competitive performance on GenAI-Bench but shows diminished performance on VideoGen-RewardBench, particularly in ties-included accuracy for visual quality and motion quality. This shortfall can be attributed to the significant improvements in visual fidelity and motion smoothness exhibited by modern T2V models, which VisionReward struggles to effectively assess. In contrast, our method VideoReward outperforms all other models on VideoGen-RewardBench, showcasing its effectiveness in evaluating outputs from state-of-the-art T2V models. Moreover, despite being trained on a distinct dataset, our approach also achieves comparable performance to VisionReward on GenAI-Bench, highlighting its strong generalization across different eras of T2V models.
Ablation Study.
As discussed in Sec. 3.2, we have highlighted several key design choices for our reward model and demonstrated their effectiveness on the validation set. To further prove these findings, we conduct an ablation study on the evaluation benchmark, providing a quantitative supplement to our analysis. Results are demonstrates in Tab. 3.
Regarding the type of reward model, we compare three variants: the regression-style reward model, the Bradley-Terry reward model, and the Bradley-Terry-With-Tied reward model. The BT reward model shows a slight advantage over the regression-style reward model, likely due to inherent strengths of pairwise annotations. Pairwise annotations more accurately capture the ranking relationships within the same context and demonstrate greater tolerance to potential noise in data annotations. The BTT reward model achieves comparable performance to the BT model on ties-excluded accuracy, but exhibits a significant improvement in ties-included accuracy. This improvement stems from the BTT model’s explicit modeling to tied pairs, enabling it to learn a more robust decision boundary that effectively captures neutral relationships in ambiguous cases.
Beyond the benefits of disentangling reward’s attribute as discussed in Sec. 3.2, we observe that adopting separate tokens rather than a shared last token further enhances the model’s overall performance. This design allows the model to better represent distinct aspects of the reward, thus improving alignment with human preferences.
Figure 5: Visual comparison of videos generated by the original pretrained model and the Flow-DPO aligned model.
5.2 Video Alignment
Training Setting.
Our pretrained model $p_{\text{ref}}$ is a transformer-based latent flow model trained using rectified flow. In all alignment experiments, we applied LoRA to fine-tune the transformer models’ linear layers, as our findings indicate that full parameter fine-tuning can degrade the model’s performance or potentially lead to model collapse. For supervised fine-tuning (SFT), we utilize only the “chosen data.” In the RWR experiments, we first calculate the mean and variance of the rewards across all training samples and normalize the rewards to have a mean of 0 and a variance of 1. For the VDM reward model, we extract and concatenate the features from the first 20 transformer blocks of our VDM. These concatenated features are then fed into three fully connected heads, which output a three-dimensional reward. All parameters of the first 20 transformer blocks are trained. For the training dataset, following the approach of Rafailov et al. (Rafailov et al., 2024), we employ VideoReward as the ground-truth reward model to simulate human feedback and relabel our training dataset. Consequently, T2V models trained on this synthetically relabeled dataset can be reliably evaluated using VideoReward , ensuring a fair assessment of their performance. More experimental details, including LoRA architectures, VDM reward model and final hyperparameter settings, are provided in the Appendix F.
Figure 6: Human evaluation of Flow-DPO aligned model vs. pretrained model on VideoGen-Eval, which contains 400 prompts.
Evaluation.
We evaluate performance using both automatic and human assessments. The automatic evaluation comprises the win rate, as determined by VideoReward , and the Vbench score. The win rate is calculated by having VideoReward assign scores to outputs from both the pretrained and aligned models, then determining the proportion of instances where the aligned model’s reward exceeds that of the pretrained model. Since VideoReward generates continuous floating-point scores, ties do not occur. Vbench serves as a fine-grained benchmark for evaluating T2V models from two main perspectives: Quality and Semantic. The Quality aspect focuses on the perceptual quality of the synthesized video without considering the input condition, corresponding to our VQ and MQ metrics. The Semantic aspect assesses whether the synthesized video aligns with the user-provided guiding condition, corresponding to TA. For human evaluation, each sample is independently reviewed by two annotators to ensure reliability. In cases of disagreement, a third annotator is consulted to resolve conflicts. We employ same random seeds to generate videos for each method, ensuring a fair comparison. We employ Vbench, VideoGen-Eval and TA-Hard prompts for evaluation. We find that the standard Vbench and VideoGen-Eval prompts are relatively straightforward in terms of text alignment, as the pretrained model consistently generates well-aligned video content. To create a more rigorous assessment, we have developed additional challenging prompts (TA-Hard) that feature combinations of two subjects and incorporate uncommon actions. For example: “A talking apple with eyes and a mouth, and a singing banana with legs hosting a talent show in a vibrant theater.” We provide a prompt subset of TA-Hard in Appendix H.
Multi-dimensional Alignment.
We employ linear scalarization (Li et al., 2020), defined as $r=\frac{1}{3}(r_{\text{vq}}+r_{\text{mq}}+r_{\text{ta}})$, to consolidate multi-dimensional preferences into a single score. For each prompt ${\bm{y}}$, the reward model annotates preferences between two generated videos, such that ${\bm{x}}^{w}$ is preferred over ${\bm{x}}^{l}$ if $r({\bm{x}}^{w},{\bm{y}})>r({\bm{x}}^{l},{\bm{y}})$. This process generates a preference dataset $D=\{({\bm{y}},{\bm{x}}^{w},{\bm{x}}^{l})\}$. As presented in Tab. 4, Flow-DPO with a constant $\beta$ achieves significant improvements over the pretrained model across three dimensions and outperforms both SFT and Flow-RWR. Conversely, Flow-DPO with a timestep-dependent $\beta$ exhibits reduced performance in the TA task compared to the pretrained model. Although it maintains high reward win rates in VQ and MQ, we observe severe reward hacking issues in the generated videos. In contrast, Flow-DPO with a constant $\beta$ also achieves high VQ and MQ scores without encountering reward hacking problems. This suggests that Flow-DPO with a timestep-dependent $\beta$ struggles to effectively learn the TA task compared to VQ and MQ tasks. We will discuss this issue in detail in Section 5.25.2. Fig. 610 demonstrates that the aligned model using Flow-DPO outperforms the pretrained model in human evaluations for both VideoGen-Eval and TA-Hard prompts.
Table 5: Single-dimensional alignment with TA. Bold: Best performance. Flow-DPO with a constant $\beta$ is the most effective method, achieving best performance without reward hacking.
Method VBench VideoGen-Eval TA-Hard
Total Quality Semantic TA TA TA
Pretrained 83.19 84.37 78.46 50.00 50.00 50.00
SFT 82.71 83.48 79.62 52.88 53.81 64.79
Flow-RWR 82.40 83.36 78.58 59.66 49.50 66.20
Flow-DPO ($\beta_{t}=\beta(1-t)^{2}$) 82.35 83.00 79.75 63.67 55.95 71.83
Flow-DPO 83.38 84.28 79.80 69.09 65.49 84.51
Single-dimensional Alignment.
We also investigate the ability of different methods to align the T2V model on a specific task: TA. As shown in Tab. 5, Flow-DPO with a constant $\beta$ achieves best performance across all datasets.
Table 6: Reward guidance using multi-dimensional rewards on VideoGen-Eval prompts. The weighted reward guidence allows users to apply arbitrary weightings to multiple alignment objectives during inference to meet personalized user requirements.
VQ:MQ:TA VQ MQ TA
0:0:1 60.56 46.48 70.42
0.1:0.1:0.8 66.50 63.73 60.86
0.1:0.1:0.6 68.94 67.59 53.28
0.5:0.5:0 86.43 93.23 26.65
Reward Guidance.
We also employ linear scalarization (Li et al., 2020) to compute a weighted sum of rewards for each dimension and backpropagate the gradients to the noised latent space. Tab. 6 demonstrates that reward guidance can improve the pretrained model in the corresponding dimensions when using the respective dimension’s reward to guide the generation process. Furthermore, the weighted reward approach simultaneously enhances performance across all dimensions.
Figure 7: Accuracy of time-dependent $\beta_{t}$ vs. constant $\beta$ for TA: Flow-DPO with a constant $\beta$ consistently outperforms the timestep-dependent $\beta$ across various settings.
Ablation on $\beta$.
We meticulously adapted diffusion-DPO to flow-based models, resulting in Equation 4.1, where $\beta_{t}=\beta(1-t)^{2}$. As illustrated in Fig. 7, under various $\beta$ settings, Flow-DPO with a constant $\beta$ consistently outperforms Flow-DPO with a timestep dependent $\beta$. We note that none of the results in the figure exhibit reward hacking issues. Timestep dependent $\beta$ may cause uneven training across different timesteps (Hang et al., 2023), since T2V models utilize shared model weights for various noise levels.
Ablation on Adding Noise to the VDM Latent Reward Model.
Tab. 7 presents the results of applying reward guidance exclusively with MQ rewards on TA-Hard prompts. The reward model trained with noised latents effectively guides the generation process, whereas the reward model trained solely on cleaned latents fails to provide meaningful gradient guidance for intermediate noised latents. Consequently, the performance after guidance is even worse than without guidance. Following the methodology of Yu et al. (2023), we also attempted to directly predict the corresponding ${\bm{x}}_{0}$ from the intermediate ${\bm{x}}_{t}$ and then use $r({\bm{x}}_{0},y)$ to backpropagate gradients. However, we found that this method cannot generate normal videos.
Table 7: Reward guidance using only MQ rewards on TA-Hard. The reward model trained with noised latents guides the generation effectively, while the model trained on cleaned latents fails to provide meaningful gradient guidance for noised latents.
VQ MQ TA
VDM w/o noise 37.1 38.6 52.9
VDM w/ noise 66.2 74.6 32.4
6 Conclusion
We have constructed a large-scale preference dataset with 200k human preference annotations, encompassing visual quality, motion quality, and text alignment for modern video generation models. Building on this dataset, we develop VideoReward, a sophisticated multi-dimensional video reward model, and establish the VideoGen-RewardBench benchmark, facilitating more precise and fair evaluation of video generation models. From a unified reinforcement learning perspective, we introduce three alignment algorithms for flow-based video generation models, demonstrating their effectiveness.
Limitations & Future Work
In our experiments, excessive training with Flow-DPO led to a significant deterioration in model quality, despite improvements in alignment across specific dimensions. To prevent this decline, we employed LoRA training. Future work can explore the simultaneous use of high-quality data for supervised learning during DPO training, aiming to preserve video quality while enhancing alignment. Additionally, our algorithms have been validated on text-to-video tasks; future work can extend the validation to other conditional generation tasks, such as image-to-video generation. Moreover, our VideoReward model is vulnerable to reward hacking, wherein human assessments indicate a marked decrease in video quality, yet the reward model continues to assign high scores. This issue arises because the reward function is differentiable, making it susceptible to manipulation. Future research should focus on developing more robust reward models, potentially by incorporating uncertainty estimates and increasing data augmentation. Additionally, there is potential to apply more RLHF algorithms, such as PPO, to flow-based video generation tasks.
Appendix A Details of the Derivation
A.1 Relation beween Velocity and Noise
Lemma A.1.
Let $X_{0}\sim q$ be a real data sample drawn from the true data distribution and $X_{1}\sim p$ be a noise sample, where $X_{0},X_{1}\in\mathbb{R}^{d}$. Define ${\bm{v}}_{t}({\bm{x}}_{t}\mid X_{0},X_{1})$ to be the conditional velocity field specified by a Rectified Flow (Liu et al., 2022), and let ${\bm{v}}_{t}^{\text{pred}}({\bm{x}}_{t})$ be the predicted marginal velocity field. Then the L2 error of the noise prediction is related to the L2 error of the velocity field prediction by
$\|X_{1}-X_{1}^{\text{pred}}({\bm{x}}_{t},t)\|^{2}\;=\;(1-t)^{2}\,\bigl{\|}{\bm{v}}_{t}({\bm{x}}_{t}\mid X_{0},X_{1})-{\bm{v}}_{t}^{\text{pred}}({\bm{x}}_{t})\bigr{\|}^{2}.$ (18)
Proof.
The Rectified Flow is a time-dependent flow $\psi:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ for $t\in[0,1]$, defined by
$\psi(X_{0},X_{1})\;=\;(1-t)\,X_{0}\;+\;t\,X_{1}.$
By definition, the marginal velocity field ${\bm{v}}_{t}({\bm{x}}_{t})$ is
${\bm{v}}_{t}({\bm{x}}_{t})\;=\;$ $\mathbb{E}\bigl{[}{\bm{v}}_{t}(X_{t}\mid X_{0},X_{1})\,\bigm{|}\,X_{t}={\bm{x}}_{t}\bigr{]}$ (19)
=\;= $\mathbb{E}\bigl{[}\dot{\psi}(X_{t}\mid X_{0},X_{1})\,\bigm{|}\,X_{t}={\bm{x}}_{t}\bigr{]}$
=\;= $\mathbb{E}\bigl{[}X_{1}-X_{0}\,\bigm{|}\,X_{t}={\bm{x}}_{t}\bigr{]}$
[…]
Meanwhile, the conditional velocity field ${\bm{v}}_{t}({\bm{x}}_{t}\mid X_{0},X_{1})$ is given by
[…]
Substituting equation 20 into equation 19, we obtain
$\bigl{\|}{\bm{v}}_{t}({\bm{x}}_{t}\mid X_{0},X_{1})\;-\;{\bm{v}}_{t}({\bm{x}}_{t})\bigr{\|}^{2}\;=\;\frac{\bigl{\|}X_{1}-\mathbb{E}[\,X_{1}\mid X_{t}={\bm{x}}_{t}\,]\bigr{\|}^{2}}{(1-t)^{2}}.$
Assuming that
$X_{1}^{\text{pred}}({\bm{x}}_{t},t)\;=\;\mathbb{E}[\,X_{1}\mid X_{t}={\bm{x}}_{t}\,]\quad\text{and}\quad{\bm{v}}_{t}^{\text{pred}}({\bm{x}}_{t})\;=\;{\bm{v}}_{t}({\bm{x}}_{t}).$
Consequently,
$\bigl{\|}X_{1}-X_{1}^{\text{pred}}({\bm{x}}_{t},t)\bigr{\|}^{2}\;=\;(1-t)^{2}\,\bigl{\|}{\bm{v}}_{t}({\bm{x}}_{t}\mid X_{0},X_{1})\;-\;{\bm{v}}_{t}^{\text{pred}}({\bm{x}}_{t})\bigr{\|}^{2}.$
A.2 Reward as Classifier Guidance
We begin by citing a lemma from the Guided Flows paper (Zheng et al., 2023)
Lemma A.2.
Let $p_{t}({\bm{x}}|{\bm{y}})$ be a Gaussian Path defined by a scheduler $(\alpha_{t},\sigma_{t})$, i.e., $p_{t}({\bm{x}}|x_{0})=\mathcal{N}({\bm{x}}|\alpha_{t}x_{0},\sigma_{t}^{2}I)$ where ${\bm{y}}\in\mathbb{R}^{k}$ is a conditioning variable, then its generating velocity field ${\bm{v}}_{t}({\bm{x}}|{\bm{y}})$ is related to the score function $\nabla\log p_{t}({\bm{x}}|{\bm{y}})$ by
${\bm{v}}_{t}({\bm{x}}|{\bm{y}})=a_{t}{\bm{x}}+b_{t}\nabla\log p_{t}({\bm{x}}|{\bm{y}}),$ (21)
where
$a_{t}=\frac{\dot{\alpha}_{t}}{\alpha_{t}},\qquad b_{t}=(\dot{\alpha}_{t}\sigma_{t}-\alpha_{t}\dot{\sigma}_{t})\frac{\sigma_{t}}{\alpha_{t}}.$ (22)
Seed Appendix A.61 of the Guided Flows paper (Zheng et al., 2023) for detailed derivations.
If we define
$\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})$ $={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla r({\bm{x}}_{t},{\bm{y}})-{\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})]$ (23)
$={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))-{\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})]$ (24)
$=(1-w){\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))]$ (25)
$=a_{t}{\bm{x}}_{t}+b_{t}\nabla\left[(1-w)\log p_{t}({\bm{x}}_{t}|{\bm{y}})+w\log\exp(r({\bm{x}}_{t},{\bm{y}}))\right]$ (26)
$=a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\tilde{p_{t}}({\bm{x}}_{t}|{\bm{y}})$ (27)
where $\tilde{p_{t}}({\bm{x}}_{t}|{\bm{y}})\propto p_{t}({\bm{x}}_{t}|{\bm{y}})^{1-w}[\exp(r({\bm{x}}_{t},{\bm{y}}))]^{w}$. We change our goal from sampling from the distribution ${p_{t}}({\bm{x}}_{t}|{\bm{y}})$ to sampling from the distribution $\tilde{p_{t}}({\bm{x}}_{t}|{\bm{y}})$.
We note, however, that this analysis shows that Reward Guided Flows are guaranteed to sample from $\tilde{q}(\cdot|{\bm{y}})$ at time $t=1$ if the probability path $\tilde{p}_{t}(\cdot|{\bm{y}})$ is close to the marginal probability path $\int p_{t}(\cdot|{\bm{x}}_{1})\tilde{q}({\bm{x}}_{1}|{\bm{y}})dx_{1}$, but it is not clear to what extent this assumption holds in practice. This also mens that $\tilde{p_{t}}({\bm{x}}_{t}|{\bm{y}})$ is also a marginal gaussian path defined by $p_{t}({\bm{x}}|{\bm{x}}_{1})=\mathcal{N}({\bm{x}}|\alpha_{t}{\bm{x}}_{1},\sigma_{t}^{2}I)$.
Simlirly, if we define
$\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})$ $={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+wb_{t}\nabla r({\bm{x}}_{t},{\bm{y}})$ (28)
$={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+wb_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))$ (29)
$=a_{t}{\bm{x}}_{t}+b_{t}\nabla\left[\log p_{t}({\bm{x}}_{t}|{\bm{y}})+w\log\exp(r({\bm{x}}_{t},{\bm{y}}))\right]$ (30)
$=a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\tilde{p_{t}}({\bm{x}}_{t}|{\bm{y}})$ (31)
where $\tilde{p_{t}}({\bm{x}}_{t}|{\bm{y}})\propto p_{t}({\bm{x}}_{t}|{\bm{y}})[\exp(r({\bm{x}}_{t},{\bm{y}}))]^{w}$. We change our goal from sampling from the distribution ${p_{t}}({\bm{x}}_{t}|{\bm{y}})$ to sampling from the distribution $\tilde{p_{t}}({\bm{x}}_{t}|{\bm{y}})$.
Reward Guidance for Rectified Flow.
Rectified Flow (Liu et al., 2022) is also a Gaussian path defined by
${\bm{x}}_{t}=(1-t){\bm{x}}_{0}+t{\bm{x}}_{1}$ (32)
where ${\bm{x}}_{1}$ is from normal Gaussian distribution. Then
$p_{t}({\bm{x}}\mid{\bm{x}}_{0})=\mathcal{N}({\bm{x}}|(1-t){\bm{x}}_{0},t^{2}I)$
where $\alpha_{t}=1-t,\sigma_{t}=t$. Then we get
$a_{t}=\frac{1}{t-1},b_{t}=\frac{t}{t-1}.$ (33)
Eq. 24 becomes
$\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[\frac{1}{1-t}{\bm{x}}_{t}+\frac{t}{1-t}\nabla r({\bm{x}}_{t},{\bm{y}})+{\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})].$ (34)
Eq. 29 becomes
$\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})-w\frac{t}{1-t}\nabla r({\bm{x}}_{t},{\bm{y}}).$ (35)
We use Eq. 24 or Eq. 29 to guide inference through reward model.
Appendix B Input Template for Reward Model
Full Input Template [VIDEO] You are tasked with evaluating a generated video based on three distinct criteria: Visual Quality, Motion Quality, and Text Alignment. Please provide a rating from 0 to 10 for each of the three categories, with 0 being the worst and 10 being the best. Each evaluation should be independent of the others.
**Visual Quality:**
Evaluate the overall visual quality of the video, with a focus on static factors. The following sub-dimensions should be considered:
- **Reasonableness:** The video should not contain any significant biological or logical errors, such as abnormal body structures or nonsensical environmental setups.
- **Clarity:** Evaluate the sharpness and visibility of the video. The image should be clear and easy to interpret, with no blurring or indistinct areas.
- **Detail Richness:** Consider the level of detail in textures, materials, lighting, and other visual elements (e.g., hair, clothing, shadows).
- **Aesthetic and Creativity:** Assess the artistic aspects of the video, including the color scheme, composition, atmosphere, depth of field, and the overall creative appeal. The scene should convey a sense of harmony and balance.
- **Safety:** The video should not contain harmful or inappropriate content, such as political, violent, or adult material. If such content is present, the image quality and satisfaction score should be the lowest possible.
Please provide the ratings of Visual Quality: <|VQ_reward|>
END
**Motion Quality:**
Assess the dynamic aspects of the video, with a focus on dynamic factors. Consider the following sub-dimensions:
- **Stability:** Evaluate the continuity and stability between frames. There should be no sudden, unnatural jumps, and the video should maintain stable attributes (e.g., no fluctuating colors, textures, or missing body parts).
- **Naturalness:** The movement should align with physical laws and be realistic. For example, clothing should flow naturally with motion, and facial expressions should change appropriately (e.g., blinking, mouth movements).
- **Aesthetic Quality:** The movement should be smooth and fluid. The transitions between different motions or camera angles should be seamless, and the overall dynamic feel should be visually pleasing.
- **Fusion:** Ensure that elements in motion (e.g., edges of the subject, hair, clothing) blend naturally with the background, without obvious artifacts or the feeling of cut-and-paste effects.
- **Clarity of Motion:** The video should be clear and smooth in motion. Pay attention to any areas where the video might have blurry or unsteady sections that hinder visual continuity.
- **Amplitude:** If the video is largely static or has little movement, assign a low score for motion quality.
Please provide the ratings of Motion Quality: <|MQ_reward|>
END
**Text Alignment:**
Assess how well the video matches the textual prompt across the following sub-dimensions:
- **Subject Relevance** Evaluate how accurately the subject(s) in the video (e.g., person, animal, object) align with the textual description. The subject should match the description in terms of number, appearance, and behavior.
- **Motion Relevance:** Evaluate if the dynamic actions (e.g., gestures, posture, facial expressions like talking or blinking) align with the described prompt. The motion should match the prompt in terms of type, scale, and direction.
- **Environment Relevance:** Assess whether the background and scene fit the prompt. This includes checking if real-world locations or scenes are accurately represented, though some stylistic adaptation is acceptable.
- **Style Relevance:** If the prompt specifies a particular artistic or stylistic style, evaluate how well the video adheres to this style.
- **Camera Movement Relevance:** Check if the camera movements (e.g., following the subject, focus shifts) are consistent with the expected behavior from the prompt.
Textual prompt - [PROMPT]
Please provide the ratings of Text Alignment: <|TA_reward|>
END
Appendix C Details of Human Annotation
We provide additional details regarding the annotation process. First, annotators are provided with detailed scoring guidelines and undergo training sessions to ensure they fully understand the criteria; Tab 8 summarizes the key points for each dimension as outlined in the guidelines. Reference examples are provided to help annotators better grasp the evaluation standards. Each sample is evaluated by three independent annotators. For training and validation sets, annotators provide pairwise preference annotations and pointwise scores for Visual Quality (VQ), Motion Quality (MQ), and Tempotal Alignment (TA). For VideoGen-RewardBench, annotators evaluate the same three aspects along with an additional overall quality, using only pairwise preferences. In cases where the annotators disagree on a sample, an additional reviewer is tasked with resolving the discrepancy. The final label is determined on the basis of the reviewer’s evaluation, ensuring consistency across the dataset. Furthermore, during the annotation process, all annotators are instructed to flag any content deemed unsafe. Videos identified as unsafe are excluded from the dataset, ensuring the safety of the data used for training and evaluation.
Table 8: Key points summary outlined in annotation guidelines for each evaluation dimension.
Evaluation Dimension Key Points Summary
Visual Quality
Considering the following dimensions introducted by non-dynamic factors:
- Image Reasonableness: The image should be objectively reasonable.
- Clarity: The image should be clear and visually sharp.
- Detail Richness: The level of intricacy in the generation of details.
- Aesthetic Creativity: The generated videos should be aesthetically pleasing.
- Safety: The generated video should not contain any disturbing or uncomfortable content.
Motion Quality
Considering the following dimensions in the dynamic process of the video:
- Dynamic Stability: The continuity and stability between frames.
- Dynamic Reasonableness: The dynamic movement should align with natural physical laws.
- Motion Aesthetic Quality: The dynamic elements should be harmonious and not stiff.
- Naturalness of Dynamic Fusion: The edges should be clear during the dynamic process.
- Motion Clarity: The motion should be easy to identify.
- Dynamic Degree: The movement should be clear, avoiding still scenes.
Text Alignment
Considering the relevance to the input text prompt description.
- Subject Relevance Relevance to the described subject characteristics and subject details.
- Dynamic Information Relevance: Relevance to actions and postures as described in the text.
- Environmental Relevance: Relevance of the environment to the input text.
- Style Relevance: Relevance to the style descriptions, if exists.
- Camera Movement Relevance: Relevance to the camera descriptions, if exists.
Appendix D Details of Reward Model Evaluation
D.1 Evaluation Benchmarks
We use two benchmarks to evaluate the performance of our reward model: GenAI-Bench (Jiang et al., 2024b) and VideoGen-RewardBench. GenAI-Bench is employed to assess the accuracy of the reward model in evaluating pre-SOTA-era T2V models, while VideoGen-RewardBench is used to evaluate its performance on modern T2V models. In this subsection, we provide a detailed description of the two benchmarks, highlighting their key parameters and differences in Fig 8 and Tab. 9. We also visualize the model coverage across the training sets of different baselines and the two evaluation benchmarks, as shown in the Fig 9.
Figure 8: Video Duration and Resolution in GenAI-Bench and VideoGen-Reward Bench
Table 9: Comparison between GenAI-Bench and VideoGen-RewardBench. Eariler Models indicates that pre-Sora-era T2V models, and Modern Models indicates that T2V models after Sora.
Benchmark Prompts and Sampled Videos Human Preference Annotations
#Samples #Prompts #Earlier Models #Modern Models #Duration #Annotations #Dimensions
GenAI-Bench 3784 508 7 (Open-Source) 3 (Open-Source) 2s - 2.5s 1891 1
VideoGen-RewardBench 4923 420 0 3 (Open-Source) 9 (Close-Source) 4s - 6s 26457 4
Figure 9: The model coverage across the training sets of different baselines and the two evaluation benchmarks. VideoScore, VisionReward, and GenAI-Bench primarily focus on pre-SoRA-era models, while our training set and VideoGen-RewardBench concentrate on state-of-the-art T2V models.
GenAI-Bench
GenAI-Bench collects data from 6 pre-SOTA-era T2V models and 4 recent open-source T2V models. Human annotations for overall quality are obtained through GenAI-Arena, resulting in a benchmark consisting of 10 T2V models, 508 prompts, and 1.9k pairs. As the videos in GenAI-Bench predominantly originate from earlier video generation models, they typically have lower resolutions (most around 320x512) and shorter durations (2s-2.5s). We consider GenAI-Bench as a benchmark to assess the performance of reward models on early-generation T2V models.
VideoGen-RewardBench
VideoGen-Eval (Zeng et al., 2024) has open-sourced a dataset containing videos generated by 9 closed-source and 3 open-source models, designed to qualitatively visualize performance differences across models. Due to its high-quality data, broad coverage of the latest advancements in T2V models, and third-party nature, we leverage VideoGen-Eval to create a fair benchmark, VideoGen-RewardBench, for evaluating reward models’ performance on modern T2V models. We manually construct 26.5k video pairs and hire annotators to assess each pair’s Visual Quality, Motion Quality, Text Alignment, and Overall Quality, providing preference labels. Ultimately, VideoGen-RewardBench includes 12 T2V models, 420 prompts, and 26.5k pairs. This benchmark represents human preferences for state-of-the-art models, with videos featuring higher resolutions (480x720 - 576x1024), longer durations (4s - 6s), and improved quality. We use VideoGen-RewardBench as the primary benchmark to evaluate reward models’ performance on modern T2V models.
D.2 Comparison Methods
Random
[…]
VideoScore
VideoScore (He et al., 2024) adopts Mantis-Idefics2-8B (Jiang et al., 2024a) as its base model and trains with point-wise data using MSE loss to model human preference scores. Since VideoScore predicts scores across multiple dimensions, and its dimension definitions differ from those in VideoGen-RewardBench, we compute both the overall accuracy and the dimension-specific (VQ, MQ, TA) accuracy by averaging the scores of five dimensions when conduct evaluation GenAI-Bench and VideoGen-RewardBench, consistent with the evaluation strategy outlined in their paper. The training data for VideoScore predominantly comes from pre-SOTA-era models, which explains its relatively better performance on GenAI-Bench, while accounts for the significant performance drop on VideoGen-RewardBench.
LiFT
LiFT (Wang et al., 2024b) adopts VILA-1.5-40B (Lin et al., 2024a) as its base model and employs a VLM-as-a-judge approach. The reward model is trained through instruction tuning with inputs, preference scores along with a critic. The model generates video scores and reasons through next-token prediction. LiFT evaluates videos across three dimensions: Video Fidelity, Motion Smoothness, and Semantic Consistency, which are similar to the dimensions defined in VideoGen-RewardBench. We calculate the overall accuracy using the average scores of these three dimensions and compute the dimension-specific accuracy using the corresponding dimensional scores. LiFT predicts discrete scores on a 1-3 scale, which often leads to ties in pairwise comparisons. When calculating accuracy without ties, we randomly convert the predicted tie labels to chosen/rejected with a 50% probability, indicating that the model is unable to distinguish the relative quality between the two samples.
VisionReward
VisionReward (Xu et al., 2024a) adopts CogVLM2-Video-12B(Hong et al., 2024) as the base model and is trained to answer a set of judgment questions about the video with a binary ”yes” or ”no” response using cross-entropy loss. During inference, VisionReward evaluates 64 checklist items, providing converted into 1/0 scores. The final score is computed as the weighted average of these individual responses. We use the final score to calculate both the overall accuracy and the VQ/MQ/TA accuracy. VisionReward’s training data includes models from the pre-SOTA era models (Chen et al., 2024) as well as recent open-source T2V models (Zheng et al., 2024; Yang et al., 2024b). It performs well on GenAI-Bench and demonstrates reasonable capabilities on VideoGen-RewardBench.
Our Reward Model
We adopts QWen2-VL-2B (Wang et al., 2024a) as the base model and train it with pair-wise data using BTT loss in Eq. 6. Scores are normalized on the validation set and averaged to obtain overall scores for evaluation and optimization. When evaluating on VideoGen-RewardBench, we sample videos at 2 FPS and a resolution of 448$\times$448, consistent with the training settings. We calculate the overall accuracy by averaging the scores across the three dimensions, and compute dimension-specific accuracies using the respective scores. For GenAI-Bench, we sample videos at 2 FPS and a resolution of 256$\times$256, as the minimum resolution in GenAI-Bench is 256$\times$256. Given the significant disparities in visual quality and motion between the GenAI-Bench videos and our training data, we utilize only the predicted TA scores to calculate the overall score.
D.3 Evaluation Metrics
Similarly to VisionReward (Xu et al., 2024a), we report two accuracy metrics: ties-included accuracy (Deutsch et al., 2023) and ties-excluded accuracy. For ties-excluded accuracy, we exclude all data labeled as ”ties” and use only data labeled as ”A wins” or ”B wins” for calculation. Since all competitors predict scores based on pointwise samples, we compute the rewards for each pair, convert the relative reward relationships into binary labels, and calculate classification accuracy. For ties-included accuracy, we adopt the $O(n^{2}\log n)$ algorithm proposed in Algorithm 1 by Deutsch et al. (2023). This method traverses all possible tie thresholds, calculates three-class accuracy for each threshold, and selects the highest accuracy as the final metric.
Appendix E Pseudo-code of Flow-DPO and Flow-NRG
The Pytorch-style implementation of the Flow-DPO loss (Eq. 4.1) is shown below:
def loss(model, ref_model, x_w, x_l, c, beta):
# model: Flow model that takes prompt condition c and timestep as inputs and predicts velocity
# ref_model: Frozen initialization of the model
# x_w: Preferred Image (latents in this work)
# x_l: Non-Preferred Image (latents in this work)
# c: Conditioning (text in this work)
# beta: Regularization Parameter
# returns: DPO loss value
timestep = torch.rand(len(x_w))
noise = torch.randn_like(x_w)
noisy_x_w = (1 - timestep) * x_w + timestep * noise
noisy_x_l = (1 - timestep) * x_l + timestep * noise
velocity_w_pred = model(noisy_x_w, c, timestep)
velocity_l_pred = model(noisy_x_l, c, timestep)
velocity_ref_w_pred = ref_model(noisy_x_w, c, timestep)
velocity_ref_l_pred = ref_model(noisy_x_l, c, timestep)
velocity_w = noise - x_w
velocity_l = noise - x_l
model_w_err = (velocity_w_pred - velocity_w).norm().pow(2)
model_l_err = (velocity_l_pred - velocity_l).norm().pow(2)
ref_w_err = (velocity_ref_w_pred - velocity_w).norm().pow(2)
ref_l_err = (velocity_ref_l_pred - velocity_l).norm().pow(2)
w_diff = model_w_err - ref_w_err
l_diff = model_l_err - ref_l_err
inside_term = -0.5 * beta * (w_diff - l_diff)
loss = -1 * log(sigmoid(inside_term))
return loss
The Pytorch-style implementation of the reward guidance (Eq. 15) is shown below:
def guidance(model, reward_model, prompt_embeds, latents, timesteps, reward_weight):
# model: Flow model that predicts velocity given latents and conditions
# reward_model: Model that evaluates the quality of latents based on prompt embeddings and timestep
# prompt_embeds: Embeddings of the text prompts
# latents: Initial noise
# timesteps: Sequence of timesteps
# reward_weight: weighting coefficient of multi-dimensional rewards
# guidance_scale: scale factor for reward guidance
dts = timesteps[:-1] - timesteps[1:]
for i, t in enumerate(timesteps):
v_pred = model(latents, prompt_embeds, t)
latents = latents.detach().requires_grad_(True)
reward = reward_model(latents, prompt_embeds, t)
reward = reward * reward_weight
reward_guidance = torch.autograd.grad(reward, latents)
v_pred = v_pred - guidance_scale * t / (1 - t) * reward_guidance
latents = latents - dts[i] * v_pred
return latents
Appendix F Hyperparameters
Algorithm-agnostic hyperparameters for SFT, Flow-RWR, Flow-DPO
Training strategy LoRA (Hu et al., 2021)
LoRA alpha 128
LoRA dropout 0.0
LoRA R 64
LoRA target-modules q_proj,k_proj,v_proj,o_proj
Optimizer Adam (Kingma, 2014)
Learning rate 5e-6
Epochs 1
Batch size 64
Flow-DPO
$\beta$ 500
Table 10: Hyperparameters for Alignment Algorithms
VLM
Training strategy Full training for vision encoder LoRA for language model
LoRA alpha 128
LoRA dropout 0.0
LoRA R 64
LoRA target-modules Linear layers in language model
Optimizer Adam (Kingma, 2014)
Learning rate 2e-6
Epochs 2
Batch size 32
$\theta$ in Eq. 6 5.0
VDM
Training strategy Full training
Optimizer Adam (Kingma, 2014)
Learning rate 5e-6
Epochs 2
Batch size 144
Reward Dimension 3
Table 11: Hyperparameters for reward modeling.
Appendix G Extended Experimental Results
Figure 10: Human evaluation of Flow-DPO on TA-Hard prompt.
Appendix H Prompt Subset of TA-Hard
A rabbit and a turtle racing on a track. The rabbit is sprinting ahead, while the turtle is steadily moving along. Spectators are cheering from the sidelines, and a finish line is visible in the distance.
A lion and a zebra playing soccer on a grassy field. The lion is dribbling the ball, while the zebra is trying to block it. The field is surrounded by trees, and other animals are watching the game.
A fox and an owl stargazing together on a hilltop. The fox is lying on its back, pointing at the stars, while the owl is perched on a nearby branch, looking through a telescope. The night sky is clear, with countless stars twinkling.
A dolphin and a sea turtle exploring a coral reef. The dolphin is swimming gracefully, while the sea turtle is gliding slowly beside it. The coral reef is vibrant with colorful corals and various marine life.
A dolphin and a whale singing together in the ocean. The dolphin is leaping out of the water, while the whale is producing deep, melodic sounds. The ocean is vast and blue, with the sun setting on the horizon.
A fox and a rabbit playing a duet on a piano in a forest clearing. The fox is playing the melody, while the rabbit is accompanying with harmony. The forest is alive with the sounds of nature, and other animals are gathered to listen.
A squirrel and a chipmunk building a treehouse in a large oak tree. The squirrel is hammering nails, while the chipmunk is holding a blueprint. The tree is tall and sturdy, with branches full of leaves.
A robot with glowing blue eyes and a human with a cybernetic arm playing basketball in a futuristic gym. The robot is dribbling the ball with precision, while the human is preparing to block the shot. The gym is equipped with advanced technology and holographic scoreboards.
A knight in shining armor and a wizard with a long, flowing beard practicing archery in a medieval courtyard. The knight is aiming at a target with a longbow, while the wizard is using magic to guide the arrows. The courtyard is surrounded by stone walls and blooming flowers.
A talking apple with eyes and a mouth, and a singing banana with legs hosting a talent show in a vibrant theater. The apple is the judge, giving feedback to contestants, while the banana is the host, entertaining the audience with jokes and songs. The theater is filled with colorful lights and excited spectators.
A pirate with a wooden leg and a mermaid with a shimmering tail playing a duet on a grand piano in an underwater cave. The pirate is playing the melody, while the mermaid is accompanying with harmony. The cave is illuminated by bioluminescent sea creatures, creating a magical atmosphere.
A superhero with a cape and a detective with a magnifying glass solving a mystery in a bustling city. The superhero is flying above the streets, scanning for clues, while the detective is examining evidence on the ground. The city is alive with activity, with skyscrapers towering overhead.
A chef with a tall hat and a robot with multiple arms cooking a gourmet meal in a state-of-the-art kitchen. The chef is chopping vegetables with precision, while the robot is simultaneously stirring, frying, and baking. The kitchen is equipped with the latest culinary technology, creating a seamless cooking experience.
A painter with a beret and a poet with a quill creating art in a sunlit studio. The painter is working on a vibrant canvas, while the poet is writing verses inspired by the artwork. The studio is filled with natural light and creative energy, with art supplies scattered around.
A spider with a square face and a green-furred puppy having a playful fight in a whimsical garden. The spider is using its web to swing around, while the puppy is playfully nipping at the spider’s legs. The garden is filled with oversized flowers and colorful mushrooms.
A talking teapot with a mustache and a dancing teacup with legs performing a tea ceremony in an enchanted forest. The teapot is pouring tea, while the teacup is twirling and dancing around. The forest is magical, with glowing plants and twinkling lights.
A robot with a television screen for a head and a toaster with arms and legs having a cooking competition in a retro kitchen. The robot is displaying recipes on its screen, while the toaster is popping out perfectly toasted bread. The kitchen is styled with vintage appliances and checkered floors.
A pair of animated scissors with eyes and a mouth and a roll of tape with tiny arms and legs wrapping presents in a festive workshop. The scissors are cutting wrapping paper with precision, while the tape is sealing the packages with a smile. The workshop is decorated with holiday lights and ornaments.
A pair of animated sneakers with eyes and a mouth and a talking basketball with a face playing a game of one-on-one on an urban basketball court. The sneakers are dribbling and making quick moves, while the basketball is bouncing and trying to score. The court is surrounded by graffiti-covered walls and cheering spectators.
A paper airplane with a scarf and a paper boat with a captain’s hat racing in the rain. The airplane glides through the air while the boat sails through puddles.
A basketball with a mohawk and a soccer ball with a bandana playing hopscotch in a playground. The basketball bounces high while the soccer ball rolls smoothly.
A mechanical knight with steam-powered joints standing guard at an ancient castle gate. Gears whir softly as its head turns to scan the surroundings, while steam occasionally escapes from its armor joints.
A wandering alchemist with potion-filled vials clinking on their belt, gathering herbs in an enchanted forest where mushrooms glow and flowers whisper secrets.
A mysterious plague doctor with clockwork enhancements peeking through their dark robes, mixing herbal remedies in a medieval apothecary shop as green smoke swirls from bubbling vials.

### v2

@@
 We then introduce VideoReward, a multi-dimensional video reward model, and examine how annotations and various design choices impact its rewarding efficacy.
-From a unified reinforcement learning perspective aimed at maximizing reward with KL regularization, we introduce three alignment algorithms for flow-based models by extending those from diffusion models.
+From a unified reinforcement learning perspective aimed at maximizing reward with KL regularization, we introduce three alignment algorithms for flow-based models.
 These include two training-time strategies: direct preference optimization for flow (Flow-DPO) and reward weighted regression for flow (Flow-RWR), and an inference-time technique, Flow-NRG, which applies reward guidance directly to noisy videos.
-Experimental results indicate that VideoReward significantly outperforms existing reward models, and Flow-DPO demonstrates superior performance compared to both Flow-RWR and standard supervised fine-tuning methods.
+Experimental results indicate that VideoReward significantly outperforms existing reward models, and Flow-DPO demonstrates superior performance compared to both Flow-RWR and supervised fine-tuning methods.
 Additionally, Flow-NRG lets users assign custom weights to multiple objectives during inference, meeting personalized video quality needs.
-Project page: https://gongyeliu.github.io/videoalign.
-Machine Learning, ICML
-[…]
 1 Introduction
-Recent advances in video generation have yielded powerful models (Kong et al., 2024; Polyak et al., 2024; Kuaishou, 2024; Brooks et al., 2024), capable of producing videos with convincing details and relatively coherent motion.
-Despite these notable achievements, current video generation systems still face persistent challenges, including unstable motion, imperfect alignment with user prompts, unsatisfactory visual quality, and insufficient adherence to human preferences (Zeng et al., 2024).
-In language modeling and image generation, learning from human preference (Ouyang et al., 2022; Zhou et al., 2023; Wallace et al., 2024) has proved highly effective in improving the response quality and aligning generative models with user expectations.
-However, applying such preference-driven alignment strategies to video generation remains in its infancy.
-One key obstacle is the lack of large-scale, high-quality preference data.
-A pioneering work (He et al., 2024) introduces a human-rated video preference dataset, and concurrent studies (Wang et al., 2024b; Liu et al., 2024b; Xu et al., 2024a) have contributed additional annotations.
-Yet, these datasets primarily focus on videos generated by earlier open-source models, which are often of relatively low quality, contain artifacts, or remain limited in duration.
-With video generation techniques rapidly advancing, the gap between such datasets and the capabilities of state-of-the-art video diffusion models (VDMs) has become increasingly pronounced.
+Advancements in video generation have led to powerful models [57, 33, 32, 5] that produce realistic details and coherent motion.
+Despite this, current systems still face challenges like unstable motion, imperfect text-video alignment and insufficient alignment with human preferences [84].
+In language model and image generation, reinforcement learning from human feedback (RLHF) [52, 90, 71] has proven effective in improving response quality and aligning models with user expectations.
+However, applying RLHF to video generation is still remains in its infancy.
+A major obstacle is the lack of a reliable reward signal.
+Existing preference datasets [19, 73, 47, 76] were collected on earlier T2V models that produced short, low-resolution clips.
+Reward models trained on such data may miss fine spatial detail and long-range dynamics, while over-penalising glitches that current T2V models already suppress.
+In addition, the design space of VLM-based reward models remains under-explored, leading to sub-optimal annotation paradigms, reward hacking issues, and entangled multi-attribute scores.
+The resulting supervision is therefore noisy, biased, and easily exploited during RLHF.
 A second challenge arises from the internal mechanisms of cutting-edge video generation models.
-Many modern systems employ rectified flow (Liu et al., 2022; Lipman et al., 2022), predicting velocity rather than noise.
-Recent studies (Wang et al., 2024b; Zhang et al., 2024a; Liu et al., 2024b; Xu et al., 2024a) have tested DPO (Rafailov et al., 2024; Wallace et al., 2024) and RWR (Peng et al., 2019; Lee et al., 2023; Furuta et al., 2024) on diffusion-based video generation approaches.
-However, adapting existing alignment methods to these flow-based models introduces new questions.
-While concurrent work (Domingo-Enrich et al., 2024) extended Diffusion-DPO to flow-based image generation, the aligned flow-based models in these explorations underperformed their unaligned counterparts.
-To address these limitations, we present a comprehensive investigation into aligning advanced flow-based video generation models with human preferences.
-First, we select 12 advanced video generation models and curate a new large-scale multi-dimensional preference dataset of approximately 182k annotated examples.
-The dataset encompasses three critical dimensions—Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA)—to better capture how users evaluate generated videos.
-We then developed a multi-dimensional video reward model, systematically examining how annotations and different design choices affect its performance.
-Subsequently, we aggregated these multi-dimensional rewards into a comprehensive metric to capture overall human preferences.
-Furthermore, from a unified reinforcement learning perspective that maximizes reward with KL regularization, we introduce three alignment algorithms for flow-based models by extending those from diffusion models.
-These include two training-time strategies—Direct Preference Optimization for Flow (Flow-DPO) and Reward Weighted Regression for Flow (Flow-RWR)—and an inference-time technique, Flow-NRG.
-Our derivation reveals that the parameter $\beta$ (which modulates the KL divergence constraint) in Flow-DPO is dependent on the time step, resulting in Flow-DPO performing poorly on certain tasks.
-However, we found that by simply fixing $\beta$, Flow-DPO becomes the most effective alignment approach in our evaluations.
-Flow-NRG is an efficient inference-time alignment algorithm that applies reward guidance directly to noisy videos during the denoising process.It allows users to apply arbitrary weightings to multiple alignment objectives during inference, eliminating the need for extensive retraining to meet personalized user requirements.
-Figure 1: Overview of Our Video Alignment Paradigm.
-(a) Human Preference Annotation (Sec.
-3.1).
-We construct a dataset of 182k (prompt, video A, video B) triplets, collecting preference annotations on Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA) from human evaluators.
-(b) Reward Mode Training (Sec.
-3.2).
-We train a VLM-based reward model using the Bradley-Terry-Model-with-Ties formulation.
-(c) Video Alignment (Sec.
-4).
-We adapt alignment techniques — DPO, RWR, and reward guidance — to flow-based video generation models and provide a comprehensive comparison of their effectiveness.
+Many modern systems employ rectified flow [48, 45], predicting velocity rather than noise.
+Recent studies [73, 85, 47, 76] have tested DPO [61, 71] and RWR [54, 36, 17] on diffusion-based video generation approaches.
+However, adapting existing alignment methods to flow-based models introduces new questions.
+A recent attempts [13] for flow matching based DPO even degrade quality compared with the unaligned baseline.
+Figure 1: Overview of Video Alignment Framework.
+(a) Human Preference Annotation.
+We construct 182k prompt-video triplets, each annotated on Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA).
+(b) Reward Modeling.
+A VLM-based reward model is trained under the Bradley-Terry-Model-with-Ties formulation.
+(c) Video Alignment.
+From a unified RL perspective, we introduce three alignment algorithms for flow-based video generation: Flow-DPO, Flow-RWR, and Reward Guidance (Flow-NRG), and provide a systematic comparison.
+To address these challenges, we present a comprehensive investigation into aligning advanced flow-based video generation models with human preferences, as shown in Fig.
+1.
+We first collect 16k high-quality prompts and render them with 12 representative T2V models, producing 182k annotated examples across three key dimensions: Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA).
+We then develop a multi-dimensional video reward model, systematically analyzing how different annotations and design choices affect its performance.
+These dimensions can be aggregated into a total reward that reflects the overall preferences of humans.
+Armed with a reliable reward model, we revisit RLHF algorithms for rectified flow.
+From a unified reinforcement learning perspective that maximizes reward with KL regularization, we derive two training-time strategies—Flow-DPO and Flow-RWR, and an inference-time technique, Flow-NRG.
+We discover that a simple extension of Diffusion-DPO performs poorly because its timestep-dependent KL term pushes the policy to overfit the objective at higher noise levels, making the model vulnerable to reward hacking.
+In contrast, our Flow-DPO removes this time-dependent term and keeps the KL weight constant, showing robust performance across all dimensions.
+Flow-NRG is an efficient inference-time alignment algorithm that applies reward guidance directly to noisy videos during the denoising process.
+It allows users to apply arbitrary weightings to multiple alignment objectives during inference, eliminating the need for extensive retraining to meet personalized user requirements.
 Our contributions can be summarized as follows:
-  • Large-Scale Preference Dataset: We create a new 182k-sized human-labeled video generation preference dataset from 12 video generation models, capturing VQ, MQ, and TA dimensions.
+  • Large-Scale Preference Dataset: We create a 182k-sized, multi-dimensional, human-labeled video generation preference dataset from 12 modern video generation models.
   • Multi-Dimensional Reward Modeling: We propose and systematically study a multi-dimensional video reward model, investigating how different design decisions influence its rewarding efficacy.
-  • VideoGen-RewardBench: To more accurately assess video reward models, we have annotated prompt-video pairs from VideoGen-Eval.
-This dataset includes a wide variety of prompts and videos generated by modern VDMs.
-The resulting benchmark consists of 26.5k video pairs, each accompanied by a preference label.
-  • Flow-Based Alignment: From a unified reinforcement learning perspective, we introduce three alignment algorithms for flow-based models: two training-time strategies (Flow-DPO, Flow-RWR) and one inference-time technique (Flow-NRG).
-Our experiments show that Flow-DPO outperforms other methods when $\beta$ is fixed.
-Additionally, Flow-NRG, using a lightweight reward model for noisy videos, enables users to apply custom weightings to multiple alignment objectives during inference.
-2 Related Works
-Evaluation and Reward Models.
-Evaluation models and reward models play a pivotal role in aligning generative models with human preferences.
-Earlier approaches and benchmarks (Huang et al., 2023; Liu et al., 2024c; Huang et al., 2024) relied on metrics like FID (Heusel et al., 2017) and CLIP scores (Radford et al., 2021) to assess visual quality and semantic consistency.
-Recent works (Wu et al., 2023a; Xu et al., 2024b; Kirstain et al., 2023; Zhang et al., 2024b; Liang et al., 2024a) have shifted towards utilizing human preference datasets to train CLIP-based models, enabling them to predict preference scores with improved accuracy.
-With the advent of large vision-language models (VLMs) (Achiam et al., 2023; Wang et al., 2024a), their powerful capabilities in visual understanding and text-visual alignment make them a natural proxy for reward modeling.
-A common approach involves replacing the token classification head of VLMs with a regression head that predicts multi-dimensional scores for diverse evaluation tasks.
-Two main learning paradigms have emerged based on the type of human annotation.
-The first paradigm relies on point-wise regression, where the model learns to fit annotated scores (He et al., 2024) or labels (Xu et al., 2024a) directly.
-Another paradigm focuses on pair-wise comparisons, leveraging Bradley-Terry (BT) (Bradley & Terry, 1952) loss to model relative preferences, which is largely unexplored for video reward model.
-Beyond these methods, some works (Wang et al., 2024b) also leverage the intrinsic reasoning capabilities of VLMs through VLM-as-a-judge (Li et al., 2023; Lin et al., 2024b), where VLMs are adopted to generate preference judgments or scores in textual format through instruction tuning.
-Despite these promising advances, most existing video reward models primarily focus on legacy video generation models, typically from the pre-Sora (OpenAI, 2024) era, which are constrained by short video durations and relatively low quality.
-Furthermore, the design and technical choices underlying the vision reward models remain underexplored.
-Our work seeks to address these limitations by focusing more on modern video generation models and investigating a broader range of reward modeling strategies.
-Alignment for Image & Video Generation.
-In large language models, Reinforcement Learning from Human Feedback (RLHF) improves alignment with human preferences, enhancing response quality (Ouyang et al., 2022; Jaech et al., 2024).
-Similar methods have been applied to image generation, using reward models or human preference data to align pretrained models.
-Key approaches include: (1) direct backpropagation with reward signals (Prabhudesai et al., 2023; Clark et al., 2023; Xu et al., 2024b; Prabhudesai et al., 2024); (2) Reward-Weighted Regression (RWR) (Peng et al., 2019; Lee et al., 2023; Furuta et al., 2024); (3) DPO-based policy optimization (Rafailov et al., 2024; Wallace et al., 2024; Dong et al., 2023; Yang et al., 2024a; Liang et al., 2024b; Yuan et al., 2024; Liu et al., 2024b; Zhang et al., 2024a; Furuta et al., 2024); (4) PPO-based policy gradients (Schulman et al., 2017) (Black et al., 2023; Fan et al., 2024); and (5) training-free alignment (Yeh et al., 2024; Tang et al., 2024; Song et al., 2023).
-These methods have successfully aligned image generation models with human preferences, improving aesthetics and semantic consistency.
-Our work applies the DPO algorithm (Rafailov et al., 2024; Wallace et al., 2024) to flow matching in video generation.
-Concurrent work (Domingo-Enrich et al., 2024) also explores similar things in image generation.
-However, they reports worse performance for the DPO-aligned model compared to the unaligned one.
-We argue that the originally derived Flow-DPO algorithm imposes a stronger KL constraint at lower noise steps, resulting in suboptimal performance.
-In contrast, using a constant KL constraint significantly improves performance on certain tasks.
-There has also been work on aligning video generation models using feedback (Furuta et al., 2024; Li et al., 2024).
-However, these approaches rely on image reward models (Kirstain et al., 2023; Wu et al., 2023a).
-Concurrent works (Yuan et al., 2024; Liu et al., 2024b; Zhang et al., 2024a) focus on aligning diffusion-based video generation models by annotating videos generated from earlier open-source models.
-Our work differs by exploring alignment techniques for advanced flow-based video generation.
+  • VideoGen-RewardBench: We create a benchmark for modern reward models by annotating prompt-video pairs from VideoGen-Eval.
+This dataset, consisting of diverse prompts and videos generated by modern VDMs, results in 26.5k annotated video pairs with preference labels.
+  • Flow-Based Alignment: From a unified RL perspective, we introduce two training-time alignment strategies (Flow-DPO, Flow-RWR) and one inference-time technique (Flow-NRG).
+Experiments show that Flow-DPO with a fixed KL term outperforms other methods.
+And Flow-NRG allows custom weightings of multiple alignment objectives during inference.
+2 Related Work
+Reward modeling [75, 77, 31, 86, 41] trains CLIP-based models on human preference datasets, while newer approaches use VLMs with regression heads to predict multi-dimensional scores.
+Learning paradigms include point-wise regression [19, 76] and pair-wise comparison via Bradley-Terry loss.
+However, most video reward models focus on short, low-quality videos from the pre-Sora era [51] and lack rigorous evaluation of design choices.
+We address these limitations by targeting modern video generation and exploring broader reward modeling strategies.
+For alignment, image generation has adopted RLHF-style techniques such as reward backpropagation [58, 77], RWR [36], DPO [61, 71], PPO [2, 16], and training-free methods [80].
+Concurrent efforts [17, 47, 76, 85, 73] extend DPO/RWR to diffusion-based video models using old generation models or image-level rewards.
+We build on this by extending DPO to flow-based video generation, proposing Flow-DPO.
+More comprehensive discussions of related work can be found in Appendix A.
 3 VideoReward
+Robust RLHF begins with a reward model that faithfully mirrors human preferences, yet existing efforts are limited in two key respects: (1) Data: existing video-preference datasets were curated for earlier T2V models and mismatch with what users prefer in modern video generation models; (2) Model design: the key technique choices for VLM-based reward models remain largely uncharted.
+Fig.
+1 (a), (b) summarises our solution.
+We first build a large-scale preference dataset collected with state-of-the-art T2V models; then perform a systematic study of reward-modeling design.
 3.1 Human Preference Data Collection
 Table 1: Statistics of the collected training dataset.
-We utilized 12 text-to-video models to generate a total of 108k videos from 16k unique prompts.
-This process ultimately resulted in 182k annotated triplets, each consisting of a prompt paired with two videos and corresponding preference annotations.
+We utilize 12 T2V models to generate 108k videos from 16k unique prompts, resulting in 182k annotated triplets.
+Each triplet consists of a prompt paired with two videos, and corresponding preference annotations.
 T2V Model Date #Videos #Anno Triplets Resolution Duration
-Pre-Sora-Era Models Gen2 (Runway, 2023) 23.06 6k 13k 768 $\times$ 1408 4s
-SVD (Blattmann et al., 2023) 23.11 6k 13k 576 $\times$ 1024 4s
-Pika 1.0 (Labs, 2023) 23.12 6k 13k 720 $\times$ 1280 3s
-Vega (VegaAI, 2023) 23.12 6k 13k 576 $\times$ 1024 4s
-Pixverse v1 (PixVerse, 2024) 24.01 6k 13k 768 $\times$ 1408 4s
-HiDream (HidreamAI, 2024) 24.01 0.3k 0.3k 768 $\times$ 1344 5s
-Modern Models Dreamina (Capcut, 2024) 24.03 16k 68k 720 $\times$ 1280 6s
-Luma (LumaLabs, 2024) 24.06 16k 57k 752 $\times$ 1360 5s
-Gen3 (Runway, 2024) 24.06 16k 55k 768 $\times$ 1280 5s
-Kling 1.0 (Kuaishou, 2024) 24.06 6k 33k 384 $\times$ 672 5s
-Pixverse v2 (PixVerse, 2024) 24.07 16k 58k 576 $\times$ 1024 5s
-Kling 1.5 (Kuaishou, 2024) 24.09 7k 28k 704 $\times$ 1280 5s
-Figure 2: Statistics of our training data.
-Existing human preference datasets for video generation (Liu et al., 2024c; Huang et al., 2024; He et al., 2024; Wang et al., 2024b; Xu et al., 2024a) primarily consist of videos generated by earlier open-source models, which are often characterized by low quality, noticeable artifacts, and short video durations.
-As advancements in video diffusion models (VDMs) continue to elevate the state-of-the-art, the gap between current preference datasets and the capabilities of modern VDMs has become increasingly pronounced.
-To bridge this gap, we focus on developing a reward model tailored for the latest VDMs.
-Inspired by Zhang et al.
-(2024b), we collect a diverse set of prompts from the internet, organizing them into 8 meta-categories (animal, architecture, food, people, plants, scenes, vehicles, objects).
-These prompts are further expanded using GPT-4o (Achiam et al., 2023).
-A filtering process is then applied to exclude unsuitable prompts (e.g., repetitive, irrelevant to motion, unsafe content).
-Subsequently, we employ our internal prompt rewriter to refine and enhance the collected prompts, resulting in a final dataset of 16k detailed prompts.
-Leveraging these prompts, we select 12 text-to-video generation models with various capabilities, producing 108k videos and constructing 182k triplets, each comprising a prompt and two corresponding videos generated by different video generation models.
-Comprehensive statistics of the dataset are provided in Tab.
-1 and Fig 2.
-To annotate these text-video triplets, we hire annotators to perform pairwise assessments across three critical dimensions: Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA).
-Details are available in appendix C.
-For each dimension, annotators are presented with a prompt and two generated videos, and asked to provide choices indicating their preference (A wins/Ties/B wins).
-Each sample is independently assigned to three annotators to ensure robustness.
-In cases of disagreement, an additional annotator will be engaged to resolve conflicts.
-This effort results in a dataset of 182k high-quality, multi-dimensional preference annotations, which were utilized for training the reward model.
-In addition to pairwise annotations, we also conduct pointwise annotations using a similar procedure, where annotators rate each video on a Likert scale from 1 to 5 for the same three dimensions (i.e., VQ, MQ, TA).
-This dual annotation setup enables us to explore the advantages and limitations of pairwise preferences versus pointwise scores as annotation strategies.
-In addition to the training data, we construct a validation set comprising 13k annotated triplets, following a similar annotation protocol.
-The prompts in the validation set are strictly excluded from the training set to ensure generalization.
-The validation set serves as a reliable benchmark to monitor the convergence during training.
-3.2 Reward Model Learning
-Following prior works that leverage Vision-Laguage Models (VLMs) for related tasks (He et al., 2024; Wang et al., 2024b; Xu et al., 2024a), we adopt Qwen2-VL-2B (Wang et al., 2024a) as the base model for our video reward framework.
-While existing studies have demonstrated the effective application of reward models in both evaluation (Wu et al., 2023b, a; He et al., 2024) and optimization (Lee et al., 2023; Wallace et al., 2024; Xu et al., 2024b; Prabhudesai et al., 2024), the underlying design choices remain insufficiently explored.
-We start with a Bradley-Terry style reward model as our baseline, exploring how technical decisions affect the video reward model’s performance.
+Pre-Sora-Era Models Gen2 [63] 23.06 6k 13k 768 $\times$ 1408 4s
+SVD [3] 23.11 6k 13k 576 $\times$ 1024 4s
+Pika 1.0 [34] 23.12 6k 13k 720 $\times$ 1280 3s
+Vega [70] 23.12 6k 13k 576 $\times$ 1024 4s
+PixVerse v1 [56] 24.01 6k 13k 768 $\times$ 1408 4s
+HiDream [21] 24.01 0.3k 0.3k 768 $\times$ 1344 5s
+Modern Models Dreamina [6] 24.03 16k 68k 720 $\times$ 1280 6s
+Luma [50] 24.06 16k 57k 752 $\times$ 1360 5s
+Gen3 [64] 24.06 16k 55k 768 $\times$ 1280 5s
+Kling 1.0 [33] 24.06 6k 33k 384 $\times$ 672 5s
+PixVerse v2 [56] 24.07 16k 58k 576 $\times$ 1024 5s
+Kling 1.5 [33] 24.09 7k 28k 704 $\times$ 1280 5s
+Existing human preference datasets for video generation [49, 26, 19, 73, 76] were primarily built on early, low-resolution T2V models that produced short, artifact-laden clips.
+As VDMs continue to evolve, modern T2V models, however, generate longer, higher-fidelity videos with smoother motion.
+Consequently, legacy datasets no longer accurately reflect what users prefer today.
+Reward models trained on such collections may miss fine spatial details and long-range dynamics while over-weighting temporal glitches already mitigated by current models.
+To bridge this gap, we develop a new preference dataset expressly for state-of-the-art VDMs.
+Prompt Collection and Video Generation.
+We collect diverse prompts from the Internet, categorize them into 8 meta-categories—animal, architecture, food, people, plants, scenes, vehicles, objects—and expand them with GPT-4o.
+After removing repetitive, irrelevant, or unsafe entries, we refine them with our in-house prompt rewriter, yielding 16000 high-quality prompts.
+12 T2V models of varying capabilities then render these prompts into 108k videos, which we pair to form 182k triplets, each comprising a prompt and two corresponding videos from distinct VDMs.
+Comprehensive dataset statistics are provided in Tab.
+1 and Fig.
+8.
+Multi-dimensional Annotation.
+Professional annotators were hired to view each triplet and record pairwise preferences(A wins / Ties / B wins) separately for Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA), producing a three-label vector per triplet.
+The same annotators also assign 1–5 Likert scores to each individual video, enabling later studies that compare pointwise and pairwise supervision.
+We reserve 13 000 triplets whose prompts never appear in training as a validation set.
+Detailed annotation protocols are provided in Appendix G.
+3.2 Reward Modeling
+Prior works [19, 73, 76] on VLM-based reward models have proved effective for both evaluation [75, 19] and optimization [36, 71, 59].
+However, their core design choices remain insufficiently explored, and current methods still suffer from issues like annotation paradigms, reward hacking and entangled multi-attribute signals.
+We select the lightweight Qwen2-VL-2B [72] as our backbone, and systematically examine three key designs and demonstrate how each yield a cleaner, more reliable reward signal for RLHF.
+We also conduct a comprehensive ablation study on the evaluation benchmark in Table 6 of Appendix E to further validate the key design choices.
 Score Regression v.s. Bradley-Terry.
-We first explore the merits of adopting a Bradley-Terry pairwise comparison approach compared to a pointwise score regression approach for training reward models.
-The BT model (Bradley & Terry, 1952) uses a pairwise log-likelihood loss to capture the probability that one video is preferred over another, enabling the prediction of rewards based on the relative preferences.
-Conversely, the pointwise regression model directly estimates the absolute quality scores by minimizing the MSE loss.
-The loss functions for these methodologies are formally defined as follows:
-$\mathcal{L}_{BT}=-\mathbb{E}_{({\bm{y}},{\bm{x}}^{w}_{0},{\bm{x}}^{l}_{0})\sim\mathcal{D}}\left[\log\left(\sigma\left(r({\bm{x}}^{w}_{0},{\bm{y}})-r({\bm{x}}^{l}_{0},{\bm{y}})\right)\right)\right]$ (1)
-$\mathcal{L}_{reg}=\mathbb{E}_{({\bm{y}},{\bm{x}}_{0},z)\sim\mathcal{D}}\left[\|r({\bm{x}}_{0},{\bm{y}})-z\|^{2}\right]$ (2)
-In these formulations, $({\bm{y}},{\bm{x}}^{w}_{0},{\bm{x}}^{l}_{0})$ denotes pairwise annotations, where ${\bm{y}}$ is the input prompt and ${\bm{x}}^{w}_{0}$ and ${\bm{x}}^{l}_{0}$ are the two generated videos with corresponding preference annotations.
-Meanwhile, $({\bm{y}},{\bm{x}}_{0},z)$ represents pointwise annotations, with $z$ being the assigned quality score.
-The function $r$ signifies the optimized reward function.
-Given that our training dataset includes both pointwise scores and pairwise preferences, annotated by the same group of annotators on the same dataset, this setup enables a direct comparison between the two paradigms.
-To conduct this comparison, we train both types of reward models on the training data with varying proportions of annotated data.
-Their performance is assessed by evaluating the average accuracy of the best-performing checkpoints on the validation set.
-Notably, for the BT reward model, we exclude samples labeled as ”Ties” and only utilize samples annotated as ”A Wins” or ”B Wins” since the BT framework is not well-suited to handle tied data.
-The comparison results are illustrated in Fig.
-3.
-Figure 3: Accuracy comparison between the BT and regression reward models across varying training data fractions (log scale).
-We observe that as the dataset size increases, both the Bradley-Terry reward model and the regression reward model exhibit notable improvements in accuracy, with the BT model consistently outperforming the regression model.
-This disparity can be attributed to the inherent characteristics of pointwise and pairwise annotations.
-Pairwise annotations are inherently more effective at capturing relative differences between samples.
-For instance, even when two videos receive the same score in pointwise annotations, annotators can still discern their relative quality in pairwise comparisons.
-This ability grants the BT reward model a distinct advantage when the dataset size is limited.
-As the dataset expands, the performance gap between the two models narrows, as the regression model benefits from the increased data to learn more accurate ranking relationships.
-Additionally, the annotation process itself varies in complexity between the two approaches.
-Pairwise preference labels are generally easier for annotators to agree upon compared to assigning precise scores in pointwise annotations.
-Recent studies (Wang et al., 2024d, c) have demonstrated that with carefully curated pointwise annotations, regression models can achieve performance levels comparable to BT reward models.
-We posit that as both the quantity and quality of the data improve, the performance differences between these approaches become negligible.
-While in the context of our current training data, the BT reward model emerges as the better-performing choice.
+We first investigate two reward learning paradigms: the Bradley-Terry (BT) model [4] and pointwise score regression.
+The BT model formulates preference learning as a probabilistic ranking task.
+Given a prompt $\bm{y}$} and paired videos $({\bm{x}}^{w}_{0},{\bm{x}}^{l}_{0})$, it optimizes $\mathcal{L}_{BT}=-\mathbb{E}\left[\log\left(\sigma\left(r({\bm{x}}^{w}_{0},{\bm{y}})-r({\bm{x}}^{l}_{0},{\bm{y}})\right)\right)\right]$, where the expectation is taken over $({\bm{y}},{\bm{x}}^{w}_{0},{\bm{x}}^{l}_{0})\sim\mathcal{D}$}.
+In contrast, score regression directly predicts a scalar quality score $z\in\mathbb{R}$ using the MSE loss: $\mathcal{L}_{reg}=\mathbb{E}\left[\|r({\bm{x}}_{0},{\bm{y}})-z\|^{2}\right],$ where the expectation is taken over $({\bm{y}},{\bm{x}}_{0},z)\sim\mathcal{D}$}.
+Since our training dataset includes both pointwise scores and pairwise preferences from the same annotators, we can directly compare between the two annotation paradigms.
+We train both types of reward models on increasing subsets of the training set and report the best validation accuracy averaged over VQ, MQ, and TA.
+Fig.
+3 presents these results.
+Figure 2: BT vs. Regression. Accuracy curves across log-scaled data fractions.
+Figure 3: BT vs. BTT. Score-difference distributions ($\Delta r$) for BT (left) and BTT (right).
+As the dataset grows, both the BT and regression-style models improve in accuracy, while the BT model remains consistently superior.
+This advantage stems from the nature of pairwise annotations, which capture subtle relative distinctions more effectively.
+Even when two videos receive identical pointwise scores, annotators can still differentiate subtle quality differences.
 Ties Matters.
-While the BT model is commonly used to capture human preferences from chosen-rejected pairs, the importance of tie annotations is often overlooked.
-Inspired by recent works (Liu et al., 2024a), we use the Bradley-Terry model with ties (BTT) (Rao & Kupper, 1967), an extension of BT that accounts for tied preferences.
-For a given prompt ${\bm{y}}$ and two generated videos ${\bm{x}}^{1}_{0}$ and ${\bm{x}}^{2}_{0}$, the BTT model defines the probabilities of each possible preference as follows:
-$P({\bm{x}}_{0}^{A}={\bm{x}}_{0}^{B}|{\bm{y}})$ $=\frac{(\theta^{2}-1)e^{r({\bm{x}}_{0}^{A},{\bm{y}})}e^{r({\bm{x}}_{0}^{B},{\bm{y}})}}{\left(e^{r({\bm{x}}_{0}^{A},{\bm{y}})}+\theta e^{r({\bm{x}}_{0}^{B},{\bm{y}})}\right)\left(\theta e^{r({\bm{x}}_{0}^{A},{\bm{y}})}+e^{r({\bm{x}}_{0}^{B},{\bm{y}})}\right)}$ (3)
-$P({\bm{x}}_{0}^{A}\succ{\bm{x}}_{0}^{B}|{\bm{y}})$ $=\frac{e^{r({\bm{x}}_{0}^{A},{\bm{y}})}}{e^{r({\bm{x}}_{0}^{B},{\bm{y}})}+\theta e^{r({\bm{x}}_{0}^{A},{\bm{y}})}}$ (4)
-$P({\bm{x}}_{0}^{B}\succ{\bm{x}}_{0}^{A}|{\bm{y}})$ $=\frac{e^{r({\bm{x}}_{0}^{B},{\bm{y}})}}{\theta e^{r({\bm{x}}_{0}^{A},{\bm{y}})}+e^{r({\bm{x}}_{0}^{B},{\bm{y}})}}$ (5)
-where $\theta$ is a parameter that controls the tendency toward ties, with a larger $\theta$ indicating a higher probability of ties.
-We empirically set $\theta=5.0$ and optimize the BTT model by minimizing the negative log-likelihood (NLL):
-$\mathcal{L}_{\text{BTT}}=-\mathbb{E}_{({\bm{y}},{\bm{x}}_{0}^{A},{\bm{x}}_{0}^{B})\sim\mathcal{D}}\Big{[}$ $\mathbbm{1}({\bm{x}}_{0}^{A}\succ{\bm{x}}_{0}^{B})\log P({\bm{x}}_{0}^{A}\succ{\bm{x}}_{0}^{B}|{\bm{y}})+$ (6)
-$\mathbbm{1}({\bm{x}}_{0}^{B}\succ{\bm{x}}_{0}^{A})\log P({\bm{x}}_{0}^{B}\succ{\bm{x}}_{0}^{A}|{\bm{y}})+$
-$\mathbbm{1}({\bm{x}}_{0}^{A}={\bm{x}}_{0}^{B})\log P({\bm{x}}_{0}^{A}={\bm{x}}_{0}^{B}|{\bm{y}})\Big{]}$
-We train both the Bradley-Terry reward model and the Bradley-Terry-With-Tied reward model under the same setting and visualize the distribution of $\Delta r=r({\bm{x}}_{0}^{A},{\bm{y}})-r({\bm{x}}_{0}^{B},{\bm{y}})$ on the validation set using the best-performing checkpoints.
-As shown in the boxplot in Fig.
-4, while the BT model effectively manages chosen/rejected pairs, it struggles to accurately differentiate tied pair.
-A large portion of ties exhibit a larger $\Delta r$ than certain chosen/rejected pairs, underscoring the BT model’s limitations in handling ambiguous cases.
-In contrast, the BTT model learns a more generalized decision boundary, enhancing it to distinguish tied pairs from clear chosen/rejected pairs based on $\Delta r$.
-This improvement stems from the explicit modeling of ties in Eq.
-3, which allows the model to incorporate this often-overlooked category, resulting in more robust performance.
-Figure 4: Visualization of the $\Delta r$ distribution for the BT reward model (Left) and the BTT reward model (Right). The BTT model effectively distinguishes tie pairs from chosen/rejected pairs.
-Separation for Decoupling.
-When adopting LLMs/MLLMs for multi-attribute reward modeling, a common practice (Ouyang et al., 2022; Touvron et al., 2023; He et al., 2024) involves extracting the last token from the final layer and applying a trainable linear projection to predict scores across multiple dimensions.
-While this approach enables efficient multi-dimensional scoring, it inherently induces coupling among dimensions, particularly between context-agnostic and context-aware attributes.
-For example, such a model may inadvertently rely on the input prompt when evaluating visual quality, causing the same video sample to receive different visual quality scores depending on its prompt.
-Although this coupling may be inherited from the correlations within the annotated data, it compromises the interpretability of reward scores and introduces challenges for the evaluation or optimization of individual dimensions.
-To decouple context-agnostic dimensions from contextual influences, we introduce separate special tokens for each dimension.
-Specially, three unique query tokens are added to represent the scores across the three evaluation dimensions.
-In the input template of our reward models, the context-agnostic tokens (i.e., VQ, MQ) are placed after the video but before the prompt, while the context-aware special tokens (i.e., TA) are positioned after the prompt.
-We then extract the feature representations from the final layer corresponding to each special token’s position and apply a shared linear head to predict the associated dimension scores.
-Owing to the autoregressive model’s causal attention mechanism, the context-agnostic tokens exclusively attend to the video content, whereas the context-aware tokens attend to both the video and the prompt.
-This strategy effectively separates the focus of each token, enabling a more robust performance of each dimension and independent evaluation for context-agnostic dimensions.
-We provide a full version of our input prompt template in appendix B.
+While vanilla BT model is widely used to capture human preferences from chosen-rejected pairs, the importance of tie annotations is often overlooked.
+Inspired by recent work [46], we adopt Bradley-Terry model with ties (BTT) [62], an extension that accounts for tied preferences.
+Formally, BTT defines a tripartite preference distribution:
+$P_{\theta}(c|{\bm{y}},{\bm{x}}_{0}^{A},{\bm{x}}_{0}^{B})=\begin{cases}\displaystyle\frac{(\theta^{2}-1)\,\exp({r_{A}})\exp({r_{B}})}{(\exp({r_{A}})+\theta\exp({r_{B}}))\,(\theta\exp({r_{A}})+\exp({r_{B}}))}&\text{Tie }(x^{A}_{0}=x^{B}_{0}),\\[8.0pt] \displaystyle\frac{\exp({r_{A}})}{\exp({r_{B}})+\theta\exp({r_{A}})}&\text{A preferred }(x^{A}_{0}\succ x^{B}_{0}),\\[8.0pt] \displaystyle\frac{\exp({r_{B}})}{\theta\exp({r_{A}})+\exp({r_{B}})}&\text{B preferred }(x^{B}_{0}\succ x^{A}_{0}).\end{cases}$ (1)
+where $c$ denotes preference choice, $\theta>1$ controls the tendency toward ties, with a larger $\theta$ increasing tie probability.
+We empirically set $\theta=5.0$ and train the BTT model by minimizing negative log-likelihood:
+$\mathcal{L}_{\text{BTT}}=-\mathbb{E}_{({\bm{y}},{\bm{x}}_{0}^{A},{\bm{x}}_{0}^{B})\sim\mathcal{D}}\Big[\sum_{c\in\{\succ,\prec,=\}}\mathbbm{1}(c)\log P_{\theta}(c|{\bm{y}},{\bm{x}}_{0}^{A},{\bm{x}}_{0}^{B})\Big]$ (2)
+We train BT and BTT reward models under identical settings and visualize $\Delta r=r({\bm{x}}_{0}^{A},{\bm{y}})-r({\bm{x}}_{0}^{B},{\bm{y}})$ on the validation set (Fig.
+3).
+Although the BT model handles chosen/rejected pairs well, it struggles to handle ties—often assigning sizeable $\Delta r$ to many tie pairs, conflating them with clear preferences.
+By contrast, BTT learns a flexible decision boundary that clusters ties near zero while preserving large margins for decisive wins and losses, yielding more reliable feedback for downstream RLHF.
+Token Positioning.
+A common approach in LLM / MLLM-based reward modeling [52, 69, 19] attaches a linear projection head to the final token to predict multi-dimensional scores.
+This method forcing prompt-independent and prompt-dependent cues into one vector, causing context leakage: the same video can receive different visual-quality scores when paired with different prompts.
+We eliminate this entanglement with a simple token-positioning strategy.
+As shown in Fig.1(b), two context-agnostic tokens, [VQ], [MQ], are inserted immediately after the video and before the prompt, so they can attend only to visual content.
+A context-aware token, [TA], is placed after the full prompt, allowing it to attend to both the video and the text.
+The final-layer embeddings of these tokens are then mapped to dimension-specific scores via a shared linear layer.
+This design removes context leakage, stabilizes visual and motion assessments, and maintains parameter efficiency.
+The full input template is provided in Appendix K.
 4 Video Alignment
-Let ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$ denote a real data sample drawn from the true data distribution, and let ${\bm{x}}_{1}\sim p({\bm{x}}_{1})$ denote a noise sample, where ${\bm{x}}_{0},{\bm{x}}_{1}\in\mathbb{R}^{d}$.
-Recent advanced image-generation models (e.g., (Esser et al., 2024)) and video-generation models (e.g., (Kong et al., 2024)) adopt the Rectified Flow framework (Liu et al., 2022), which defines the “noised” data ${\bm{x}}_{t}$ as
-${\bm{x}}_{t}=(1-t)\,{\bm{x}}_{0}\;+\;t\,{\bm{x}}_{1},$
-for $t\in[0,1]$.
-Then we can train a transformer model to directly regress the velocity field ${\bm{v}}_{\theta}({\bm{x}}_{t},t)$ by minimizing the Flow Matching objective (Lipman et al., 2022; Liu et al., 2022):
+With a high-fidelity reward model established, we next introduce three alignment methods for flow-based generation models under a unified RL objective: two training-time algorithms—Flow-DPO and Flow-RWR, and one inference-time guidance technique, Flow-NRG (Fig.
+1 (c)).
+4.1 Preliminaries
+Rectified Flow.
+Let ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$ denote a data sample from the real data distribution, and ${\bm{x}}_{1}\sim p({\bm{x}}_{1})$ denote a noise sample, where ${\bm{x}}_{0},{\bm{x}}_{1}\in\mathbb{R}^{d}$.
+Recent advanced image [15] and video generation models [57, 32] adopt the Rectified Flow [48], which defines the “noisy” data ${\bm{x}}_{t}$ as ${\bm{x}}_{t}=(1-t)\,{\bm{x}}_{0}\;+\;t\,{\bm{x}}_{1},$ where $t\in[0,1]$.
+Then we can train a transformer model to regress the velocity field ${\bm{v}}_{\theta}({\bm{x}}_{t},t)$ by minimizing the Flow Matching objective [45, 48]:
 $\mathcal{L}(\theta)=\mathbb{E}_{t,\;{\bm{x}}_{0}\sim q({\bm{x}}_{0}),\;{\bm{x}}_{1}\sim p({\bm{x}}_{1})}\bigl[\;\|{\bm{v}}\;-\;{\bm{v}}_{\theta}({\bm{x}}_{t},t)\|^{2}\bigr],$
 where the target velocity field is ${\bm{v}}={\bm{x}}_{1}-{\bm{x}}_{0}$.
+RLHF.
+The goal of RLHF is to learn a conditional distribution $p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})$ that maximizes the reward $r({\bm{x}}_{0},{\bm{y}})$ while controls the KL-divergence from the reference model $p_{\text{ref}}$ via a coefficient $\beta$:
+$\max_{p_{\theta}}\ \mathbb{E}_{{\bm{y}}\sim\mathcal{D}_{c},{\bm{x}}_{0}\sim p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})}\left[r({\bm{x}}_{0},{\bm{y}})\right]-\beta\,\mathbb{D}_{\text{KL}}\left[p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})\,\|\,p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})\right].$ (3)
 4.2 Flow-DPO
-Consider a fixed dataset $\mathcal{D}=\{{\bm{y}},{\bm{x}}_{0}^{w},{\bm{x}}_{0}^{l}\}$, where each sample consists of a prompt ${\bm{y}}$ and two videos ${\bm{x}}_{0}^{w}$ and ${\bm{x}}_{0}^{l}$ generated by a reference model $p_{\text{ref}}$, with human annotations indicating that ${\bm{x}}_{0}^{w}$ is preferred over ${\bm{x}}_{0}^{l}$ (i.e., ${\bm{x}}_{0}^{w}\succ{\bm{x}}_{0}^{l}$).
-The goal of RLHF is to learn a conditional distribution $p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})$ that maximizes a reward model $r({\bm{x}}_{0},{\bm{y}})$, while controlling the regularization term (KL-divergence) from the reference model $p_{\text{ref}}$ via a coefficient $\beta$:
-$\max_{p_{\theta}}$ $\mathbb{E}_{{\bm{y}}\sim\mathcal{D}_{c},{\bm{x}}_{0}\sim p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})}\left[r({\bm{x}}_{0},{\bm{y}})\right]$
-$-\beta\,\mathbb{D}_{\text{KL}}\left[p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})\,\|\,p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})\right]$ (7)
-In this context, $p_{\text{ref}}$ corresponds to $p_{\text{ref},0}$ and $p_{\theta}$ corresponds to $p_{\theta,0}$.
-For the sake of simplicity, the timestep subscripts are omitted.
-However, RLHF is usually unstable and resource intensive (Rafailov et al., 2024).
-To address this issue, Diffusion-DPO aligns diffusion models with human preferences by directly solving the above RLHF objective (Eq.
-4.1) analytically.
-It interprets alignment as a classification problem, and optimizes a policy to satisfy human preferences by supervised training.
-For simplicity, we omit the conditioning prompt ${\bm{y}}$ in the following equations.
-The Diffusion-DPO objective $\mathcal{L}_{\text{DD}}(\theta)$ is given by:
-$-\mathbb{E}\Bigg{[}\log\sigma\Bigg{(}$ $-\frac{\beta}{2}\Big{(}\|{\bm{\epsilon}}^{w}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{w},t)\|^{2}-\|{\bm{\epsilon}}^{w}-{\bm{\epsilon}}_{\text{ref}}({\bm{x}}_{t}^{w},t)\|^{2}$
-$\quad-\bigl{(}\|{\bm{\epsilon}}^{l}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{l},t)\|^{2}-\|{\bm{\epsilon}}^{l}-{\bm{\epsilon}}_{\text{ref}}({\bm{x}}_{t}^{l},t)\|^{2}\bigr{)}\Big{)}\Bigg{)}\Bigg{]}$ (8)
-where ${\bm{x}}_{t}^{*}=(1-t)\ {\bm{x}}_{0}^{*}+t\ {\bm{\epsilon}}$, ${\bm{\epsilon}}\sim\mathcal{N}(0,{\bm{I}})$.
-The superscript “*” indicates either “$w$” (for the preferred data) or “$l$” (for the less preferred data).
+Consider a training set $\mathcal{D}=\{{\bm{y}},{\bm{x}}_{0}^{w},{\bm{x}}_{0}^{l}\}$, where $y$ is the prompt and human annotations prefer generated video ${\bm{x}}_{0}^{w}$ to ${\bm{x}}_{0}^{l}$ (i.e., ${\bm{x}}_{0}^{w}\succ{\bm{x}}_{0}^{l}$).
+Direct Preference Optimization (DPO) [61] aligns models with human preferences by analytically solving the RLHF objective (Eq.
+3) and optimizing the policy via supervised learning.
+Extending this idea to diffusion models, Diffusion-DPO [71] derives a DPO-style loss under the diffusion paradigm.
+The resulting objective $\mathcal{L}_{\text{DD}}(\theta)$ is formulated as:
+$\mathbb{E}\Bigg[\log\sigma\Bigg(-\frac{\beta}{2}\Big(\|{\bm{\epsilon}}^{w}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{w},t)\|^{2}-\|{\bm{\epsilon}}^{w}-{\bm{\epsilon}}_{\text{ref}}({\bm{x}}_{t}^{w},t)\|^{2}-\bigl(\|{\bm{\epsilon}}^{l}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{l},t)\|^{2}-\|{\bm{\epsilon}}^{l}-{\bm{\epsilon}}_{\text{ref}}({\bm{x}}_{t}^{l},t)\|^{2}\bigr)\Big)\Bigg)\Bigg]$ (4)
 The expectation is taken over samples $\{{\bm{x}}_{0}^{w},{\bm{x}}_{0}^{l}\}\sim\mathcal{D}$ and the noise schedule $t$.
 In Rectified Flow, we relate the noise vector $\bm{\epsilon}$$^{*}$ to a velocity field $\bm{v}$$^{*}$.
-Specifically, Lemma A.1 in Appendix A shows that
+Specifically, Lemma B.1 in Appendix B shows that
 $\|{\bm{\epsilon}}^{*}-{\bm{\epsilon}}_{\text{pred}}({\bm{x}}_{t}^{*},t)\|^{2}=(1-t)^{2}\|{\bm{v}}^{*}-{\bm{v}}_{\text{pred}}(x_{t}^{*},t)\|^{2},$ (5)
-where ${\bm{\epsilon}}_{\text{pred}}$ and ${\bm{v}}_{\text{pred}}$ refer to predictions either from the model $p_{\theta}$ or the reference model $p_{\text{ref}}$.
-By substituting Eq.
-9 into Eq.
-4.1, we obtain the final Flow-DPO loss $\mathcal{L}_{\text{FD}}(\theta)$:
+By substituting Eq. 5 into Eq. 4, we obtain the final Flow-DPO loss $\mathcal{L}_{\text{FD}}(\theta)$:
 $-\mathbb{E}\Bigg[\log\sigma\Bigg($ $-\frac{\beta_{t}}{2}\Big(\|{\bm{v}}^{w}-{\bm{v}}_{\theta}({\bm{x}}_{t}^{w},t)\|^{2}-\|{\bm{v}}^{w}-{\bm{v}}_{\text{ref}}({\bm{x}}_{t}^{w},t)\|^{2}$
-$\quad-\bigl{(}\|{\bm{v}}^{l}-{\bm{v}}_{\theta}({\bm{x}}_{t}^{l},t)\|^{2}-\|{\bm{v}}^{l}-{\bm{v}}_{\text{ref}}({\bm{x}}_{t}^{l},t)\|^{2}\bigr{)}\Big{)}\Bigg{)}\Bigg{]},$ (10)
-where $\beta_{t}=\beta\,(1-t)^{2}$ and the expectation is taken over samples $\{{\bm{x}}_{0}^{w},{\bm{x}}_{0}^{l}\}\sim\mathcal{D}$ and the noise schedule $t$.
+$\quad-\bigl(\|{\bm{v}}^{l}-{\bm{v}}_{\theta}({\bm{x}}_{t}^{l},t)\|^{2}-\|{\bm{v}}^{l}-{\bm{v}}_{\text{ref}}({\bm{x}}_{t}^{l},t)\|^{2}\bigr)\Big)\Bigg)\Bigg]$ (6)
+where $\beta_{t}=\beta\,(1-t)^{2}$ .
 Intuitively, minimizing $\mathcal{L}_{\text{FD}}(\theta)$ guides the predicted velocity field ${\bm{v}}_{\theta}$ closer to the target velocity ${\bm{v}}^{w}$ of the “preferred” data, while pushing it away from ${\bm{v}}^{l}$ (the “less preferred” data).
-The strength of this preference signal depends on the differences between the predicted errors and the corresponding reference errors, $\|{\bm{v}}^{w}-{\bm{v}}_{\text{ref}}({\bm{x}}_{t}^{w},t)\|^{2}$ and $\|{\bm{v}}^{l}-{\bm{v}}_{\text{ref}}({\bm{x}}_{t}^{l},t)\|^{2}$.
-We provide a pseudo-code of Flow-DPO in Appendix E.
+The strength of this preference signal depends on the differences between the predicted errors and the corresponding reference errors.
+We provide Flow-DPO pseudo-code in Appendix C.
 Discussion on $\beta_{t}$.
-The parameter $\beta_{t}$ governs the strength of the KL divergence constraint, determining the extent to which the model can diverge from the base reference model (Rafailov et al., 2024; Wallace et al., 2024).
-Note that from our derivation, $\beta_{t}=\beta(1-t)^{2}$.
-As the timestep $t$ approaches 1, $\beta_{t}$ approaches zero, resulting in a relatively weak KL constraint.
-Conversely, as $t$ approaches 0, $\beta_{t}$ converges to the original $\beta$, enforcing a stronger KL constraint.
+The KL coefficient $\beta_{t}$ controls how far the learned policy is allowed to deviate from the reference model [61, 71].
+A direct derivation yields the schedule $\beta_{t}=\beta(1-t)^{2}$.
+The penalty $\beta_{t}$ vanishes as $t$ approaches 1 and reaches $\beta$ at $t=0$.
 This scheduling strategy causes the model to prioritize alignment at higher noise levels.
-Our experiments reveal that this degrades performance, while using a constant $\beta_{t}$ provides better results.
-Similar phenomena have been observed in DDPM’s (Ho et al., 2020) standard diffusion training objective, where discarding the weighting in denoising score matching leads to improved sample quality, as discussed in Section 3.4 of their paper.
+Unlike in diffusion models, our experiments reveal that this schedule degrades alignment performance in rectified flow, leading to reward hacking and visual artifacts.
+Inspired by a similar observation in DDPM’s [22] training objective, where discarding the weighting in denoising score matching improves sample quality, we instead adopt a constant $\beta$.
+This adjustment leads to more stable training and improved alignment across all reward dimensions.
 We provide a more detailed discussion of this in Section 5.2.
 4.3 Flow-RWR
-Drawing inspiration from the application of Reward-weighted Regression (RWR) (Peters & Schaal, 2007) in diffusion models (Lee et al., 2023; Furuta et al., 2024), we propose a counterpart for flow-based models based on expectation-maximization.
+Drawing inspiration from the application of Reward-weighted Regression (RWR) [55] in diffusion models [36, 17], we propose a counterpart for flow-based models based on expectation-maximization [10].
 Starting from the general KL-regularized reward-maximization problem in Eq.
-4.1, prior work (Rafailov et al., 2024; Peters & Schaal, 2007) shows that its optimal closed-form solution can be written as
-$p_{\theta}(\mathbf{x}_{0}\mid\mathbf{y})=\frac{1}{Z(\mathbf{y})}\,p_{\text{ref}}(\mathbf{x}_{0}\mid\mathbf{y})\,\exp\biggl{(}\frac{1}{\beta}\,r(\mathbf{x}_{0},\mathbf{y})\biggr{)},$ (11)
-where $Z(\mathbf{y})=\sum_{\mathbf{x}_{0}}p_{\text{ref}}(\mathbf{x}_{0}\mid\mathbf{y})\exp\bigl{(}\tfrac{1}{\beta}r(\mathbf{x}_{0},\mathbf{y})\bigr{)}$ is the partition function.
-Following (Furuta et al., 2024), we can obtains the RWR loss by M-step:
-$\mathcal{L}_{\text{RWR}}(\theta)=-\mathbb{E}_{{\bm{y}}\sim\mathcal{D}_{c}}\bigl{[}\frac{1}{Z({\bm{y}})}\exp(\frac{r({\bm{x}}_{0},{\bm{y}})}{\beta})\log p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})\bigr{]}.$
-In practice, we omit the intractable normalization $Z(\mathbf{y})$ and $\beta$, and we employ Eq.
-37 in Wallace et al.
-(2024) as a simplified upper bound of the negative log-likelihood.
-This yields the approximate form
+3, prior work [61] shows that its optimal closed-form solution can be written as:
+$p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})=\frac{1}{Z({\bm{y}})}\,p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})\,\exp\biggl(\frac{1}{\beta}\,r({\bm{x}}_{0},{\bm{y}})\biggr),$ (7)
+where $Z({\bm{y}})=\sum_{{\bm{x}}_{0}}p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})\exp\bigl(\tfrac{1}{\beta}r({\bm{x}}_{0},{\bm{y}})\bigr)$ is the partition function.
+Following [17], we can obtains the RWR loss:
 $\mathcal{L}_{\text{RWR}}(\theta)=\mathbb{E}_{{\bm{y}},{\bm{x}}_{0},{\bm{\epsilon}},t}\bigl[\exp(r({\bm{x}}_{0},{\bm{y}}))\|{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})\|^{2}\bigr].$ (8)
-Substituting Eq. 9 into Eq. 12 then provides the final RWR objective for velocity prediction:
+For rectified-flow models, we formulate this as a reward-weighted velocity regression:
 $\mathcal{L}_{\text{RWR}}(\theta)=\mathbb{E}\bigl[\exp(r({\bm{x}}_{0},{\bm{y}}))\|{\bm{v}}-{\bm{v}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})\|^{2}\bigr],$ (9)
-where $\mathbf{v}$ is the ground-truth velocity, and $\mathbf{v}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})$ represents the predicted velocity under parameters $\theta$.
-We omit $(1-t)^{2}$ as in Sec.
-4.1.
-By applying RWR in this way, our flow-based model learns to generate velocity fields that emphasize high-reward samples, paralleling how diffusion-based models align samples via reward-weighted likelihood maximization.
+As in Flow-DPO, we omit the $(1-t)^{2}$ factor for better performance.
 4.4 Noisy Reward Guidance
-Recall that Eq.
-4.1 has a closed-form solution as presented in Eq.
-11.
-This expression indicates that the goal of RLHF is to transform the original distribution $p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})$ into the new target distribution $p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})$.
-Since the constants $\beta$ and $w$ can be absorbed into $r({\bm{x}}_{0},{\bm{y}})$, Eq.
-11 becomes
+Recall that the KL-regularized RL objective (Eq.
+3) admits a closed-form solution(Eq.
+7), which transform the original distribution $p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})$ into the new target distribution $p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})$.
+Since the constants $\beta$ and $w$ can be absorbed into $r({\bm{x}}_{0},{\bm{y}})$, the closed-form solution becomes:
 $p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})\;\propto\;p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})\,[\exp(r({\bm{x}}_{0},{\bm{y}}))]^{w},$ (10)
-where $w\in\mathbb{R}$ controls the strength of the reward guidance.
-Notably, this formulation aligns with the concept of classifier guidance (Dhariwal & Nichol, 2021; Song et al., 2020), so we refer to it as reward guidance.
-As we proved in Appendix A.2, for Rectified Flow, if we set the marginal velocity field to
+where $w\in\mathbb{R}$ controls the strength of the reward guidance. For rectified flow, as we proved in Appendix B.2, this reweighting can be achieved by shifting the velocity field:
 $\tilde{{\bm{v}}}_{t}({\bm{x}}_{t}\mid{\bm{y}})={\bm{v}}_{t}({\bm{x}}_{t}\mid{\bm{y}})-w\,\frac{t}{1-t}\,\nabla r({\bm{x}}_{t},{\bm{y}}),$ (11)
-the marginal distribution is modified from $p_{t}({\bm{x}}_{t}\mid{\bm{y}})$ to
-$\tilde{p}_{t}({\bm{x}}_{t}\mid{\bm{y}})\;\propto\;p_{t}({\bm{x}}_{t}\mid{\bm{y}})\,[\exp(r({\bm{x}}_{t},{\bm{y}}))]^{w}.$
-When $t=0$, the resulting samples are drawn from Eq.
-14.
-We provide a pseudo-code in Appendix E.
-To producing meaningful reward values for noised videos ${\bm{x}}_{t}$, we train a time-dependent reward function.
-We use the same dataset $D=\{{\bm{y}},{\bm{x}}^{w},{\bm{x}}^{l}\}$ described in Section 3.2, where ${\bm{x}}^{w}$ and ${\bm{x}}^{l}$ represent two videos with different preference levels.
-In order to train our reward function, we assume that applying the same noise to both videos preserves their preference relationship.
-Formally, for each pair $({\bm{x}}^{w},{\bm{x}}^{l})$, we define
-${\bm{x}}_{t}^{w}=(1-t)\,{\bm{x}}_{0}+t\,{\bm{x}}_{1}^{w},$
-${\bm{x}}_{t}^{l}=(1-t)\,{\bm{x}}_{0}+t\,{\bm{x}}_{1}^{l},\text{with}\,{\bm{x}}_{t}^{w}\succ{\bm{x}}_{t}^{l}.$ (16)
-Here, ${\bm{x}}_{0}$ is the original video, while ${\bm{x}}_{1}^{w}$ and ${\bm{x}}_{1}^{l}$ are the noised versions corresponding to the more preferred and less preferred videos, respectively.
-We then adopt the BT model to learn the reward function from these noised videos.
-Specifically, we minimize the following loss:
-$\mathcal{L}_{BT}(\theta)=-\mathbb{E}\bigl{[}\log\bigl{(}\sigma\bigl{(}r_{\theta}({\bm{x}}_{t}^{w},t)-r_{\theta}({\bm{x}}_{t}^{l},t)\bigr{)}\bigr{)}\bigr{]},$ (17)
-Table 2: Preference accuracy on GenAI-Bench and VideoGen-Eval.
-w/ Ties indicates that accuracy is calculated with ties included (Deutsch et al., 2023), and w/o Ties excludes tied pairs when calculating accuracy.
-* denotes that for LiFT, ties prediction are randomly converted to chosen/rejected with a 0.5 probability due to a large number of ties produced by the model.
+This modification of the marginal velocity field alters the sampling distribution to match the target form in Eq.
+equation 10.
+Since this formulation is structurally similar to classifier guidance [12, 67], we refer to it as reward guidance.
+Pseudo-code is provided in Appendix C.
+Efficient Reward on Noisy Latents.
+Computing $\nabla r$ in pixel space requires back-propagating through the full VAE decoder, which is computationally expensive.
+To address this, we propose training a lightweight, time-dependent reward model $r_{\theta}(\cdot,t)$ directly in latent space.
+For each preference pair $({\bm{x}}^{w},{\bm{x}}^{l})$, we apply identical noise to both videos, assuming their relative preference remains unchanged.
+We then adopt the Bradley–Terry loss to learn the reward function from these noised videos.
+Leveraging the fact that modern VDMs are already well trained on noisy latents, we can reuse a few early layers from the pretrained backbone to construct the reward model, avoiding the need for heavy retraining.
+We apply Eq.
+equation 11 at each inference step (except at $t=1$), enabling efficient inference-time alignment in latent space.
+5 Experiments
+Table 2: Preference accuracy on GenAI-Bench and VideoGen-RewardBench.
+For ties-excluded accuracy, we calculate accuracy using only the data labeled as “A wins” or “B wins”.
+For ties-included accuracy, we use the algorithm from Deutsch et al.
+[11], which tests various tie thresholds and selects the one that maximizes three-class accuracy.
 Bold: best performance.
@@
 Random 33.67 49.84 41.86 50.30 47.42 49.86 59.07 49.64 37.25 50.40
-VideoScore (He et al., 2024) 49.03 71.69 41.80 50.22 47.41 47.72 59.05 51.09 37.24 50.34
-LiFT* (Wang et al., 2024b) 37.06 58.39 39.08 57.26 47.53 55.97 59.04 54.91 33.79 55.43
-VisionRewrd (Xu et al., 2024a) 51.56 72.41 56.77 67.59 47.43 59.03 59.03 60.98 46.56 61.15
+VideoScore [19] 49.03 71.69 41.80 50.22 47.41 47.72 59.05 51.09 37.24 50.34
+LiFT [73] 37.06 58.39 39.08 57.26 47.53 55.97 59.04 54.91 33.79 55.43
+VisionRewrd [76] 51.56 72.41 56.77 67.59 47.43 59.03 59.03 60.98 46.56 61.15
 Ours 49.41 72.89 61.26 73.59 59.68 75.66 66.03 74.70 53.80 72.20
-Table 3: Ablation study on reward model type, seprate tokens and data augmentation. Bold: Best Performance.
-Variants RM Type Separate Tokens GenAI-Bench VideoGen-RewardBench
-Overall Accuracy Overall Accuracy VQ Accuracy MQ Accuracy TA Accuracy
-w/ Ties w/o Ties w/ Ties w/o Ties w/ Ties w/o Ties w/ Ties w/o Ties w/ Ties w/o Ties
-I Regression 48.28 71.13 58.39 70.16 54.23 73.61 61.16 65.56 52.60 70.95
-II BT 47.74 71.21 61.22 72.58 52.33 77.10 59.43 73.50 53.06 71.62
-III BTT 48.27 70.89 61.50 73.39 60.52 76.31 64.64 72.40 53.55 72.12
-IV BTT 49.41 72.89 61.26 73.59 59.68 75.66 66.03 74.70 53.80 72.20
-where the expectation is taken over $({\bm{x}}^{w},{\bm{x}}^{l},{\bm{y}})\sim\mathcal{D},\,t\sim U(0,1)$.
-In this way, the reward function $r_{\theta}(\cdot,t)$ learns to preserve the relative preference between noised videos at any noise level $t$.
-However, advanced video generation models typically work in a latent space and rely on a VAE (Kingma, 2013) to decode the latent representation back into a video.
-In Eq.
-15, computing the gradient of the reward with respect to ${\bm{x}}_{t}$ would require first decoding the latent representation into a video and then passing the gradient through the large VAE decoder back to the latent space, incurring high computational cost and memory usage.
-To address this challenge, we propose training the reward model directly within the latent space using a pretrained VDM.
-Since the VDM has already been trained on noisy latent representations, we can develop a lightweight, time-dependent reward function by utilizing only the initial layers of the pretrained model rather than the entire VDM.
-Subsequently, we apply Equation 15 at each inference step to implement reward-guided generation.
-5 Experiments
-In this section, we empirically evaluate our reward model and reinforcement learning alignment algorithms on Text-to-Video (T2V) tasks.
 5.1 Reward Learning
 Training Setting.
-We utilize Qwen2-VL-2B (Wang et al., 2024a) as the backbone for our reward model and train it with BTT loss.
-To finetune the model, LoRA (Hu et al., 2021) is applied to update all linear layers in the language model, while the vision encoder’s parameters are fully optimized.
-The training process is conducted with a batch of 32 and a learning rate of $2\times 10^{-6}$, with the model trained over two epochs.
-This setup requires approximately 72 A800 GPU hours.
+We use Qwen2-VL-2B [72] as the backbone of our reward model, trained with BTT loss.
 Several observations were made during training.
 First, higher video resolution and more frames generally improved the reward model’s performance.
-Second, using a stable sampling interval instead of a fixed frame number significantly enhanced the accuracy of motion quality evaluations, particularly when trained with videos of varying durations.
-In practice, we sample videos at 2 fps, with a resolution of approximately $448\times 448$ pixels during the training process.
-The original aspect ratio is preserved to ensure visual quality.
+Second, using a stable sampling interval instead of a fixed frame number significantly enhanced motion quality evaluations, especially for videos of varying lengths.
+In practice, we sample videos at 2 fps, with a resolution of approximately $448\times 448$ pixels while preserving the original asoect ratio.
+Hyperparameters are in Appendix I.
 Evaluation.
-To assess the performance of our reward model in evaluating T2V generation, we employ a comprehensive evaluation framework designed to ensure a fair comparison across different models and mitigate potential domain biases.
-We utilize two primary benchmarks for evaluation.
-(1).
-VideoGen-RewardBench: Derived from the third-party prompt-video dataset VideoGen-Eval (Zeng et al., 2024), VideoGen-RewardBench is specifically designed to evaluate modern T2V models.
-To address the absence of human annotations in VideoGen-Eval, we manually construct 26.5k triplets and hire annotators to provide pairwise preference labels, following the procedure outlined in Sec.
-3.1.
-Additionally, annotators assess the overall quality between the two videos, serving as a universal label for reward models across various dimensions.
-(2).
-GenAI-Bench (Jiang et al., 2024b): : Serving as a complementary benchmark, GenAI-Bench features short videos (2 seconds in duration) generated by pre-Sora-era T2V models.
-This benchmark enables us to evaluate our reward model’s capability in assessing earlier-generation models, despite their limited representation in our training set.
-We summarize the key difference between VideoGen-RewardBench and GenAI-Bench in Appendix D.1.
-We compare our reward model against existing baselines, including VideoScore (He et al., 2024), as well as two concurrent works: LiFT (Wang et al., 2024b), and VisionReward (Xu et al., 2024a).
-Consistent with prevailing practices in LLMs (Lambert et al., 2024), we employ pairwise accuracy as a primary metric.
-Inspired by VisionReward (Xu et al., 2024a), we report two accuracy metrics: ties-included accuracy (Deutsch et al., 2023) and ties-excluded accuracy.
-We calculate the overall accuracy on GenAI-Bench and both overall and dimension-specific (VQ, MQ, TA) accuracy on VideoGen-RewardBench.
-Additional details of the evaluation across different methods are provided in the Appendix D.2.
+We evaluate our reward model on two benchmarks targeting different generations of T2V models: (1) VideoGen-RewardBench: Built upon the third-party prompt-video dataset VideoGen-Eval [84], this benchmark targets modern T2V models.
+We address the lack of human annotations in VideoGen-Eval by manually constructing 26.5k triplets and hiring annotators to provide pairwise preference labels.
+Annotators also assess overall video quality, serving as a universal label across all dimensions.
+(2) GenAI-Bench [29]: GenAI-Bench features short (2-seconds) videos generated by pre-Sora-era T2V models, enabling evaluation on earlier-generation outputs.
+A detailed comparison between the two benchmarks is provided in Appendix H.1.
+We evaluate our reward model against existing baselines, including VideoScore [19], as well as two concurrent works: LiFT [73] and VisionReward [76].
+Consistent with practices in LLM evaluation [35], we use pairwise accuracy, reporting both ties-included [11] and ties-excluded accuracy.
+We calculate overall accuracy on GenAI-Bench and dimension-specific (VQ, MQ, TA) accuracy on VideoGen-RewardBench.
+Additional evaluation details are in Appendix H.2 and Appendix H.3
+Main Results.
+Tab.
+2 presents the pairwise accuracy across both benchmarks.
+VideoScore performs well on GenAI-Bench but fails on VideoGen-RewardBench, indicating poor generalization to modern T2V models.
+LiFT improves over VideoScore on modern videos but remains below 60% accuracy, showing limited pairwise discrimination ability.
+VisionReward demonstrates competitive performance on GenAI-Bench but underperforms on VideoGen-RewardBench, especially on visual and motion dimensions under ties-included settings.
+This drop stems from its difficulty in assessing the improved fidelity and motion smoothness of modern outputs.
+In contrast, our method VideoReward outperforms all other models on VideoGen-RewardBench, showcasing its strong alignment with human preferences on modern T2V generations.
+Moreover, despite being trained on a disjoint dataset (see Fig.
+10), it still achieves comparable performance on GenAI-Bench, indicating robust generalization across different eras of T2V models.
+Ablation studies are provided in Appendix E.
+5.2 Video Alignment
+Training Setting.
+Our pretrained model $p_{\text{ref}}$ is an internal, research-purpose video generation model based on Transformer architecture [53], which is trained using rectified flow (see Appendix 7 for details).
+Following SD3 [15], all alignment experiments fine-tune the Transformer using LoRA [24].
+For training-based alignment methods, including SFT, Flow-DPO, and Flow-RWR, we adopt VideoReward to provide the reward signals.
+For reward guidance, we employ the latent reward model to generate rewards.
+For supervised fine-tuning (SFT), we utilize only the “chosen data”.
+Following Rafailov et al.
+[61], we employ VideoReward as the ground-truth reward model to simulate human feedback and relabel our training dataset, ensuring that models optimised on these synthetic labels can be evaluated fairly by the same reward function.
+Hyperparameter settings are provided in the Appendix I.
 Table 3: Multi-dimensional alignment with VQ:MQ:TA = 1:1:1.
@@
 Flow-DPO 83.41 84.19 80.26 93.42 69.08 75.43 90.95 81.01 68.26 77.46 71.43 73.24
-Main Results.
-Tab 2 presents the pairwise accuracy of various reward models on GenAI-Bench (representing pre-Sora-era models) and VideoGen-RewardBench (representing modern T2V models).
-VideoScore is primarily designed to evaluate videos from earlier-generation models.
-It performs well on GenAI-Bench but exhibits close-to-random predictions on VideoGen-RewardBench, reflecting its inability to generalize to the improved capabilities of modern T2V models.
-LiFT adopts the VLM-as-a-judge approach to train its reward model, but it struggles with pairwise differentiation.
-This is because LiFT tends to assign similar scores to both videos in a pair.
-When calculating its ties-excluded accuracy, we randomly convert ties prediction to chosen/rejected categories with a 0.5 probability, indicating the model fail to distinguish between video qualities.
-VisionReward demonstrates competitive performance on GenAI-Bench but shows diminished performance on VideoGen-RewardBench, particularly in ties-included accuracy for visual quality and motion quality.
-This shortfall can be attributed to the significant improvements in visual fidelity and motion smoothness exhibited by modern T2V models, which VisionReward struggles to effectively assess.
-In contrast, our method VideoReward outperforms all other models on VideoGen-RewardBench, showcasing its effectiveness in evaluating outputs from state-of-the-art T2V models.
-Moreover, despite being trained on a distinct dataset, our approach also achieves comparable performance to VisionReward on GenAI-Bench, highlighting its strong generalization across different eras of T2V models.
-Ablation Study.
-As discussed in Sec.
-3.2, we have highlighted several key design choices for our reward model and demonstrated their effectiveness on the validation set.
-To further prove these findings, we conduct an ablation study on the evaluation benchmark, providing a quantitative supplement to our analysis.
-Results are demonstrates in Tab.
-3.
-Regarding the type of reward model, we compare three variants: the regression-style reward model, the Bradley-Terry reward model, and the Bradley-Terry-With-Tied reward model.
-The BT reward model shows a slight advantage over the regression-style reward model, likely due to inherent strengths of pairwise annotations.
-Pairwise annotations more accurately capture the ranking relationships within the same context and demonstrate greater tolerance to potential noise in data annotations.
-The BTT reward model achieves comparable performance to the BT model on ties-excluded accuracy, but exhibits a significant improvement in ties-included accuracy.
-This improvement stems from the BTT model’s explicit modeling to tied pairs, enabling it to learn a more robust decision boundary that effectively captures neutral relationships in ambiguous cases.
-Beyond the benefits of disentangling reward’s attribute as discussed in Sec.
-3.2, we observe that adopting separate tokens rather than a shared last token further enhances the model’s overall performance.
-This design allows the model to better represent distinct aspects of the reward, thus improving alignment with human preferences.
-Figure 5: Visual comparison of videos generated by the original pretrained model and the Flow-DPO aligned model.
-5.2 Video Alignment
-Training Setting.
-Our pretrained model $p_{\text{ref}}$ is a transformer-based latent flow model trained using rectified flow.
-In all alignment experiments, we applied LoRA to fine-tune the transformer models’ linear layers, as our findings indicate that full parameter fine-tuning can degrade the model’s performance or potentially lead to model collapse.
-For supervised fine-tuning (SFT), we utilize only the “chosen data.” In the RWR experiments, we first calculate the mean and variance of the rewards across all training samples and normalize the rewards to have a mean of 0 and a variance of 1.
-For the VDM reward model, we extract and concatenate the features from the first 20 transformer blocks of our VDM.
-These concatenated features are then fed into three fully connected heads, which output a three-dimensional reward.
-All parameters of the first 20 transformer blocks are trained.
-For the training dataset, following the approach of Rafailov et al.
-(Rafailov et al., 2024), we employ VideoReward as the ground-truth reward model to simulate human feedback and relabel our training dataset.
-Consequently, T2V models trained on this synthetically relabeled dataset can be reliably evaluated using VideoReward , ensuring a fair assessment of their performance.
-More experimental details, including LoRA architectures, VDM reward model and final hyperparameter settings, are provided in the Appendix F.
-Figure 6: Human evaluation of Flow-DPO aligned model vs. pretrained model on VideoGen-Eval, which contains 400 prompts.
-Evaluation.
-We evaluate performance using both automatic and human assessments.
-The automatic evaluation comprises the win rate, as determined by VideoReward , and the Vbench score.
-The win rate is calculated by having VideoReward assign scores to outputs from both the pretrained and aligned models, then determining the proportion of instances where the aligned model’s reward exceeds that of the pretrained model.
-Since VideoReward generates continuous floating-point scores, ties do not occur.
-Vbench serves as a fine-grained benchmark for evaluating T2V models from two main perspectives: Quality and Semantic.
-The Quality aspect focuses on the perceptual quality of the synthesized video without considering the input condition, corresponding to our VQ and MQ metrics.
-The Semantic aspect assesses whether the synthesized video aligns with the user-provided guiding condition, corresponding to TA.
-For human evaluation, each sample is independently reviewed by two annotators to ensure reliability.
-In cases of disagreement, a third annotator is consulted to resolve conflicts.
-We employ same random seeds to generate videos for each method, ensuring a fair comparison.
-We employ Vbench, VideoGen-Eval and TA-Hard prompts for evaluation.
-We find that the standard Vbench and VideoGen-Eval prompts are relatively straightforward in terms of text alignment, as the pretrained model consistently generates well-aligned video content.
-To create a more rigorous assessment, we have developed additional challenging prompts (TA-Hard) that feature combinations of two subjects and incorporate uncommon actions.
-For example: “A talking apple with eyes and a mouth, and a singing banana with legs hosting a talent show in a vibrant theater.” We provide a prompt subset of TA-Hard in Appendix H.
-Multi-dimensional Alignment.
-We employ linear scalarization (Li et al., 2020), defined as $r=\frac{1}{3}(r_{\text{vq}}+r_{\text{mq}}+r_{\text{ta}})$, to consolidate multi-dimensional preferences into a single score.
-For each prompt ${\bm{y}}$, the reward model annotates preferences between two generated videos, such that ${\bm{x}}^{w}$ is preferred over ${\bm{x}}^{l}$ if $r({\bm{x}}^{w},{\bm{y}})>r({\bm{x}}^{l},{\bm{y}})$.
-This process generates a preference dataset $D=\{({\bm{y}},{\bm{x}}^{w},{\bm{x}}^{l})\}$.
-As presented in Tab.
-4, Flow-DPO with a constant $\beta$ achieves significant improvements over the pretrained model across three dimensions and outperforms both SFT and Flow-RWR.
-Conversely, Flow-DPO with a timestep-dependent $\beta$ exhibits reduced performance in the TA task compared to the pretrained model.
-Although it maintains high reward win rates in VQ and MQ, we observe severe reward hacking issues in the generated videos.
-In contrast, Flow-DPO with a constant $\beta$ also achieves high VQ and MQ scores without encountering reward hacking problems.
-This suggests that Flow-DPO with a timestep-dependent $\beta$ struggles to effectively learn the TA task compared to VQ and MQ tasks.
-We will discuss this issue in detail in Section 5.25.2.
-Fig.
-610 demonstrates that the aligned model using Flow-DPO outperforms the pretrained model in human evaluations for both VideoGen-Eval and TA-Hard prompts.
 Table 4: Single-dimensional alignment with TA. Bold: Best performance. Flow-DPO with a constant $\beta$ is the most effective method, achieving best performance without reward hacking.
@@
 Flow-DPO 83.38 84.28 79.80 69.09 65.49 84.51
+Figure 4: Comparison of videos generated by the original model and the Flow-DPO aligned model.
+Evaluation.
+We evaluate performance using both automatic and human assessments.
+For automatic evaluation, we measure the win rate (via VideoReward) and the Vbench score.
+The win rate is the proportion of cases where VideoReward assigns a higher reward to the aligned model than to the pretrained model.
+Vbench is a fine-grained T2V benchmark that assesses Quality and Semantic alignment.
+For human evaluation, each sample is reviewed by two annotators, with a third annotator resolving disagreements.
+Identical random seeds are used across methods for strict comparability.
+Our evaluation uses prompts from Vbench, VideoGen-Eval, and a new TA-Hard set that stresses complex semantics, since we notice that the Vbench and VideoGen-Eval prompts are relatively straightforward in terms of text alignment.
+Appendix L lists a subset of TA-Hard prompts.
+Figure 5: Human evaluation of Flow-DPO aligned model vs. pretrained model on VideoGen-Eval.
+Figure 6: Accuracy of time-dependent $\beta_{t}$ v.s. constant $\beta$ for TA.
+Multi-dimensional Alignment.
+We use linear scalarization [40], $r=\frac{1}{3}(r_{\text{vq}}+r_{\text{mq}}+r_{\text{ta}})$, to aggregate multi-dimensional preferences into a single score, and forming a dataset $D=\{({\bm{y}},{\bm{x}}^{w},{\bm{x}}^{l})\}$ where $r({\bm{x}}^{w},{\bm{y}})>r({\bm{x}}^{l},{\bm{y}})$.
+Table 3 shows that Flow-DPO with a constant $\beta$ significantly improves over the pretrained model across all three dimensions and outperforms both SFT and Flow-RWR.
+In contrast, Flow-DPO with timestep-dependent $\beta$ underperforms the pretrained model on TA, despite high VQ/MQ reward win rates due to reward hacking.
+Meanwhile, the constant-$\beta$ variant achieves high VQ/MQ scores without such hacking issues, suggesting it learns TA more effectively.
+We discuss this further in Sec.
+5.2.
+Figures 6 confirms these findings in human studies.
 Single-dimensional Alignment.
-We also investigate the ability of different methods to align the T2V model on a specific task: TA. As shown in Tab. 5, Flow-DPO with a constant $\beta$ achieves best performance across all datasets.
-Table 6: Reward guidance using multi-dimensional rewards on VideoGen-Eval prompts.
-The weighted reward guidence allows users to apply arbitrary weightings to multiple alignment objectives during inference to meet personalized user requirements.
+We also investigate the ability of different methods on a specific task: TA. Tab. 4 shows that Flow-DPO with a constant $\beta$ achieves best performance across all datasets.
+Reward Guidance.
+We apply linear scalarization to combine dimension rewards into a weighted sum and backpropagate gradients to the noised latent with guidance strength $w=100$.
+Table 5 shows that users can steer generation toward custom trade-offs by simply choosing custom weights.
+Table 5: Reward guidance on VideoGen-Eval.
 VQ:MQ:TA VQ MQ TA
@@
 0.5:0.5:0 86.43 93.23 26.65
-Reward Guidance.
-We also employ linear scalarization (Li et al., 2020) to compute a weighted sum of rewards for each dimension and backpropagate the gradients to the noised latent space.
-Tab.
-6 demonstrates that reward guidance can improve the pretrained model in the corresponding dimensions when using the respective dimension’s reward to guide the generation process.
-Furthermore, the weighted reward approach simultaneously enhances performance across all dimensions.
-Figure 7: Accuracy of time-dependent $\beta_{t}$ vs. constant $\beta$ for TA: Flow-DPO with a constant $\beta$ consistently outperforms the timestep-dependent $\beta$ across various settings.
 Ablation on $\beta$.
 We meticulously adapted diffusion-DPO to flow-based models, resulting in Equation 4.2, where $\beta_{t}=\beta(1-t)^{2}$.
-As illustrated in Fig.
-7, under various $\beta$ settings, Flow-DPO with a constant $\beta$ consistently outperforms Flow-DPO with a timestep dependent $\beta$.
-We note that none of the results in the figure exhibit reward hacking issues.
-Timestep dependent $\beta$ may cause uneven training across different timesteps (Hang et al., 2023), since T2V models utilize shared model weights for various noise levels.
-Ablation on Adding Noise to the VDM Latent Reward Model.
-Tab.
-7 presents the results of applying reward guidance exclusively with MQ rewards on TA-Hard prompts.
-The reward model trained with noised latents effectively guides the generation process, whereas the reward model trained solely on cleaned latents fails to provide meaningful gradient guidance for intermediate noised latents.
-Consequently, the performance after guidance is even worse than without guidance.
-Following the methodology of Yu et al.
-(2023), we also attempted to directly predict the corresponding ${\bm{x}}_{0}$ from the intermediate ${\bm{x}}_{t}$ and then use $r({\bm{x}}_{0},y)$ to backpropagate gradients.
-However, we found that this method cannot generate normal videos.
-Table 7: Reward guidance using only MQ rewards on TA-Hard.
-The reward model trained with noised latents guides the generation effectively, while the model trained on cleaned latents fails to provide meaningful gradient guidance for noised latents.
-VQ MQ TA
-VDM w/o noise 37.1 38.6 52.9
-VDM w/ noise 66.2 74.6 32.4
+Figure 6 shows that under various $\beta$ values, a constant $\beta$ consistently outperforms the timestep-dependent variant.
+This is likely because varying $\beta$ across timesteps leads to uneven training [18], given that T2V models use shared weights for different noise levels.
 6 Conclusion
-We have constructed a large-scale preference dataset with 200k human preference annotations, encompassing visual quality, motion quality, and text alignment for modern video generation models.
-Building on this dataset, we develop VideoReward, a sophisticated multi-dimensional video reward model, and establish the VideoGen-RewardBench benchmark, facilitating more precise and fair evaluation of video generation models.
-From a unified reinforcement learning perspective, we introduce three alignment algorithms for flow-based video generation models, demonstrating their effectiveness.
-Limitations & Future Work
+We present a large-scale preference dataset of 182k human annotations covering visual quality, motion quality, and text alignment for modern video generation models.
+Building on this dataset, we introduce VideoReward, a multi-dimensional video reward model, and establish the VideoGen-RewardBench benchmark for more accurate and fair evaluations.
+From a unified reinforcement learning perspective, we further propose three alignment algorithms tailored to flow-based video generation, demonstrating their effectiveness in practice.
+7 Limitations & Future Work
 In our experiments, excessive training with Flow-DPO led to a significant deterioration in model quality, despite improvements in alignment across specific dimensions.
@@
 Additionally, our algorithms have been validated on text-to-video tasks; future work can extend the validation to other conditional generation tasks, such as image-to-video generation.
-Moreover, our VideoReward model is vulnerable to reward hacking, wherein human assessments indicate a marked decrease in video quality, yet the reward model continues to assign high scores.
+Moreover, despite our efforts like incorporating tie annotations into the modeling, our VideoReward model still carries a potential risk of reward hacking, where human assessments indicate a marked decrease in video quality, yet the reward model continues to assign high scores.
 This issue arises because the reward function is differentiable, making it susceptible to manipulation.
@@
 Additionally, there is potential to apply more RLHF algorithms, such as PPO, to flow-based video generation tasks.
-Appendix A Details of the Derivation
-A.1 Relation beween Velocity and Noise
-Lemma A.1.
+Acknowledgements
+This work was supported by the JC STEM Lab of AI for Science and Engineering, which is funded by The Hong Kong Jockey Club Charities Trust and the Research Grants Council of Hong Kong (Project No.
+CUHK14213224).
+It was also supported by the National Natural Science Foundation of China (Grant No.
+62576191).
+Appendix A Related Works
+Evaluation and Reward Models.
+Evaluation models and reward models play a pivotal role in aligning generative models with human preferences.
+Earlier approaches and benchmarks Huang et al.
+[2023], Liu et al.
+[2024c], Huang et al.
+[2024] relied on metrics like FID Heusel et al.
+[2017] and CLIP scores Radford et al.
+[2021] to assess visual quality and semantic consistency.
+Recent works Wu et al.
+[2023], Xu et al.
+[2024b], Kirstain et al.
+[2023], Zhang et al.
+[2024b], Liang et al.
+[2024a] have shifted towards utilizing human preference datasets to train CLIP-based models, enabling them to predict preference scores with improved accuracy.
+With the advent of large vision-language models (VLMs) Achiam et al.
+[2023], Wang et al.
+[2024a], their powerful capabilities in visual understanding and text-visual alignment make them a natural proxy for reward modeling.
+A common approach involves replacing the token classification head of VLMs with a regression head that predicts multi-dimensional scores for diverse evaluation tasks.
+Two main learning paradigms have emerged based on the type of human annotation.
+The first paradigm relies on point-wise regression, where the model learns to fit annotated scores He et al.
+[2024] or labels Xu et al.
+[2024a] directly.
+Another paradigm focuses on pair-wise comparisons Wu et al.
+[2025], leveraging Bradley-Terry (BT) Bradley and Terry [1952] loss or rank loss to model relative preferences, which is largely unexplored for video reward model.
+Beyond these methods, some works Wang et al.
+[2024b] also leverage the intrinsic reasoning capabilities of VLMs through VLM-as-a-judge Li et al.
+[2023], Lin et al.
+[2024b], where VLMs are adopted to generate preference judgments or scores in textual format through instruction tuning.
+Despite these promising advances, most existing video reward models primarily focus on legacy video generation models, typically from the pre-Sora OpenAI [2024] era, which are constrained by short video durations and relatively low quality.
+Furthermore, the technical choices underlying the vision reward models remain underexplored.
+Our work seeks to address these limitations by focusing more on modern video generation models and investigating a broader range of reward modeling strategies.
+Alignment for Image & Video Generation.
+In large language models, Reinforcement Learning from Human Feedback (RLHF) improves alignment with human preferences, enhancing response quality Ouyang et al.
+[2022], Jaech et al.
+[2024].
+Similar methods have been applied to image generation, using reward models or human preference data to align pretrained models.
+Key approaches include: (1) direct backpropagation with reward signals Prabhudesai et al.
+[2023], Clark et al.
+[2023], Xu et al.
+[2024b], Prabhudesai et al.
+[2024]; (2) Reward-Weighted Regression (RWR) Peng et al.
+[2019], Lee et al.
+[2023], Furuta et al.
+[2024]; (3) DPO-based policy optimization Rafailov et al.
+[2024], Wallace et al.
+[2024], Liang et al.
+[2024b], Dong et al.
+[2023], Yang et al.
+[2024a], Liang et al.
+[2024b], Yuan et al.
+[2024b], Liu et al.
+[2024b], Zhang et al.
+[2024a], Furuta et al.
+[2024]; (4) PPO-based policy gradients Schulman et al.
+[2017] Black et al.
+[2023], Fan et al.
+[2024], Zhang et al.
+[2024c], Chen et al.
+[2024a]; and (5) training-free alignment Yeh et al.
+[2024], Tang et al.
+[2024], Song et al.
+[2023].
+These methods have successfully aligned image generation models with human preferences, improving aesthetics and semantic consistency.
+They focus on improving the accuracy of rewards on noised images using reward models trained on clean images, whereas our Flow-NRG directly trains noise-aware reward models to obtain accurate gradients.
+Our work applies the DPO algorithm Rafailov et al.
+[2024], Wallace et al.
+[2024] to flow matching in video generation.
+Concurrent work Domingo-Enrich et al.
+[2024] also explores similar things in image generation.
+However, they reports worse performance for the DPO-aligned model compared to the unaligned one.
+We argue that the originally derived Flow-DPO algorithm imposes a stronger KL constraint at lower noise steps, resulting in suboptimal performance.
+In contrast, using a constant KL constraint significantly improves performance on certain tasks.
+Some prior work explores aligning video generation models using direct backpropagation with differentiable rewards Yuan et al.
+[2024a], Li et al.
+[2024b, a], Prabhudesai et al.
+[2024], often relying on image reward models Kirstain et al.
+[2023], Wu et al.
+[2023].
+However, these approaches cannot be directly applied to modern T2V generation with a VLM-based video reward and large VAE decoders, as they exceed the memory limits of existing GPU setups, requiring specialized engineering techniques to handle.
+Discussion with Concurrent Video Alignment Works.
+Several concurrent works also explore aligning video generation models using feedback.
+Furuta et al.
+[2024] derives a unified probabilistic objective for offline RL fine-tuning of text-to-video models for DPO and RWR.
+VideoDPO Liu et al.
+[2024b] introduces a re-weighting factor on the Diffusion-DPO loss to adjust the impact of each pair, encouraging the model to learn more effectively from pairs with clearer distinctions.
+VisionReward Xu et al.
+[2024a] ensures that all dimensions of the chosen data outperform those of the rejected data, addressing the issue of confounding factors in preference data.
+OnlineVPO Zhang et al.
+[2024a] presents an online DPO algorithm to tackle off-policy optimization.
+LIFT Wang et al.
+[2024b] proposes applying RWR on synthesized datasets while simultaneously performing supervised learning on real video-text datasets, as synthesized videos often suffer from low temporal consistency.
+All of these works use Diffusion-DPO or Diffusion-RWR and focus on aligning diffusion-based video generation models, where the videos in the preference datasets are either generated by earlier open-source models or use image reward models directly.
+In contrast, our work explores alignment techniques for advanced flow-based video generation.
+We extend Diffusion-DPO into Flow-DPO, but our derivation reveals that the parameter $\beta$ (the KL divergence constraint) in Flow-DPO is timestep-dependent, which leads to suboptimal performance on certain tasks.
+However, fixing $\beta$ resolves this issue.
+SPO Liang et al.
+[2024b] also assumes that the preference order between pairs of images remains consistent when adding the same noise, and constructs win-lose pairs for noised images, proposing step-aware preference optimization.
+Our work differs from SPO in that while SPO focuses on improving the training method DPO, our Flow-NRG specifically targets training noise-ware reward model on noised videos.
+Appendix B Details of the Derivation
+B.1 Relation beween Velocity and Noise
+Lemma B.1.
 Let $X_{0}\sim q$ be a real data sample drawn from the true data distribution and $X_{1}\sim p$ be a noise sample, where $X_{0},X_{1}\in\mathbb{R}^{d}$.
-Define ${\bm{v}}_{t}({\bm{x}}_{t}\mid X_{0},X_{1})$ to be the conditional velocity field specified by a Rectified Flow (Liu et al., 2022), and let ${\bm{v}}_{t}^{\text{pred}}({\bm{x}}_{t})$ be the predicted marginal velocity field.
+Define ${\bm{v}}_{t}({\bm{x}}_{t}\mid X_{0},X_{1})$ to be the conditional velocity field specified by a Rectified Flow Liu et al.
+[2022], and let ${\bm{v}}_{t}^{\text{pred}}({\bm{x}}_{t})$ be the predicted marginal velocity field.
 Then the L2 error of the noise prediction is related to the L2 error of the velocity field prediction by
@@
 ${\bm{v}}_{t}({\bm{x}}_{t})\;=\;$ $\mathbb{E}\bigl[{\bm{v}}_{t}(X_{t}\mid X_{0},X_{1})\,\bigm|\,X_{t}={\bm{x}}_{t}\bigr]$ (13)
-=\;= $\mathbb{E}\bigl{[}\dot{\psi}(X_{t}\mid X_{0},X_{1})\,\bigm{|}\,X_{t}={\bm{x}}_{t}\bigr{]}$
-=\;= $\mathbb{E}\bigl{[}X_{1}-X_{0}\,\bigm{|}\,X_{t}={\bm{x}}_{t}\bigr{]}$
-[…]
-[…]
-[…]
+$=\;$ $\mathbb{E}\bigl[\dot{\psi}(X_{t}\mid X_{0},X_{1})\,\bigm|\,X_{t}={\bm{x}}_{t}\bigr]$
+$=\;$ $\mathbb{E}\bigl[X_{1}-X_{0}\,\bigm|\,X_{t}={\bm{x}}_{t}\bigr]$
+$=\;$ $\mathbb{E}\Bigl[\frac{\,X_{1}-\bigl((1-t)\,X_{0}+t\,X_{1}\bigr)\,}{\,1-t\,}\,\Bigm|\,X_{t}={\bm{x}}_{t}\Bigr]$
+$=\;$ $\mathbb{E}\Bigl[\frac{\,X_{1}-{\bm{x}}_{t}\,}{\,1-t\,}\,\Bigm|\,X_{t}={\bm{x}}_{t}\Bigr]$
+$=\;$ $\frac{\mathbb{E}[\,X_{1}\mid X_{t}={\bm{x}}_{t}\,]-{\bm{x}}_{t}}{\,1-t\,}.$
 Meanwhile, the conditional velocity field ${\bm{v}}_{t}({\bm{x}}_{t}\mid X_{0},X_{1})$ is given by
-[…]
+${\bm{v}}_{t}({\bm{x}}_{t}\mid X_{0},X_{1})\;=\;\frac{\,X_{1}-{\bm{x}}_{t}\,}{\,1-t\,}.$ (14)
 Substituting equation 14 into equation 13, we obtain
@@
 $\bigl\|X_{1}-X_{1}^{\text{pred}}({\bm{x}}_{t},t)\bigr\|^{2}\;=\;(1-t)^{2}\,\bigl\|{\bm{v}}_{t}({\bm{x}}_{t}\mid X_{0},X_{1})\;-\;{\bm{v}}_{t}^{\text{pred}}({\bm{x}}_{t})\bigr\|^{2}.$
-A.2 Reward as Classifier Guidance
-We begin by citing a lemma from the Guided Flows paper (Zheng et al., 2023)
-Lemma A.2.
+B.2 Reward as Classifier Guidance
+We begin by citing a lemma from the Guided Flows paper Zheng et al. [2023].
+Lemma B.2.
 Let $p_{t}({\bm{x}}|{\bm{y}})$ be a Gaussian Path defined by a scheduler $(\alpha_{t},\sigma_{t})$, i.e., $p_{t}({\bm{x}}|x_{0})=\mathcal{N}({\bm{x}}|\alpha_{t}x_{0},\sigma_{t}^{2}I)$ where ${\bm{y}}\in\mathbb{R}^{k}$ is a conditioning variable, then its generating velocity field ${\bm{v}}_{t}({\bm{x}}|{\bm{y}})$ is related to the score function $\nabla\log p_{t}({\bm{x}}|{\bm{y}})$ by
@@
 $a_{t}=\frac{\dot{\alpha}_{t}}{\alpha_{t}},\qquad b_{t}=(\dot{\alpha}_{t}\sigma_{t}-\alpha_{t}\dot{\sigma}_{t})\frac{\sigma_{t}}{\alpha_{t}}.$ (16)
-Seed Appendix A.61 of the Guided Flows paper (Zheng et al., 2023) for detailed derivations.
+Seed Appendix A.61 of the Guided Flows paper Zheng et al. [2023] for detailed derivations.
 If we define
-$\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})$ $={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla r({\bm{x}}_{t},{\bm{y}})-{\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})]$ (23)
-$={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))-{\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})]$ (24)
-$=(1-w){\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))]$ (25)
-$=a_{t}{\bm{x}}_{t}+b_{t}\nabla\left[(1-w)\log p_{t}({\bm{x}}_{t}|{\bm{y}})+w\log\exp(r({\bm{x}}_{t},{\bm{y}}))\right]$ (26)
+$\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})$ $={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla r({\bm{x}}_{t},{\bm{y}})-{\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})]$
+$={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))-{\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})]$
+$=(1-w){\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))]$
+$=a_{t}{\bm{x}}_{t}+b_{t}\nabla\left[(1-w)\log p_{t}({\bm{x}}_{t}|{\bm{y}})+w\log\exp(r({\bm{x}}_{t},{\bm{y}}))\right]$
 $=a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\tilde{p_{t}}({\bm{x}}_{t}|{\bm{y}})$ (17)
@@
 We change our goal from sampling from the distribution ${p_{t}}({\bm{x}}_{t}|{\bm{y}})$ to sampling from the distribution $\tilde{p_{t}}({\bm{x}}_{t}|{\bm{y}})$.
-We note, however, that this analysis shows that Reward Guided Flows are guaranteed to sample from $\tilde{q}(\cdot|{\bm{y}})$ at time $t=1$ if the probability path $\tilde{p}_{t}(\cdot|{\bm{y}})$ is close to the marginal probability path $\int p_{t}(\cdot|{\bm{x}}_{1})\tilde{q}({\bm{x}}_{1}|{\bm{y}})dx_{1}$, but it is not clear to what extent this assumption holds in practice.
+We note that this analysis shows that Reward Guided Flows are guaranteed to sample from $\tilde{q}(\cdot|{\bm{y}})$ at time $t=1$ if the probability path $\tilde{p}_{t}(\cdot|{\bm{y}})$ is close to the marginal probability path $\int p_{t}(\cdot|{\bm{x}}_{1})\tilde{q}({\bm{x}}_{1}|{\bm{y}})dx_{1}$, but it is not clear to what extent this assumption holds in practice.
 This also mens that $\tilde{p_{t}}({\bm{x}}_{t}|{\bm{y}})$ is also a marginal gaussian path defined by $p_{t}({\bm{x}}|{\bm{x}}_{1})=\mathcal{N}({\bm{x}}|\alpha_{t}{\bm{x}}_{1},\sigma_{t}^{2}I)$.
 Simlirly, if we define
-$\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})$ $={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+wb_{t}\nabla r({\bm{x}}_{t},{\bm{y}})$ (28)
-$={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+wb_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))$ (29)
-$=a_{t}{\bm{x}}_{t}+b_{t}\nabla\left[\log p_{t}({\bm{x}}_{t}|{\bm{y}})+w\log\exp(r({\bm{x}}_{t},{\bm{y}}))\right]$ (30)
+$\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})$ $={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+wb_{t}\nabla r({\bm{x}}_{t},{\bm{y}})$
+$={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+wb_{t}\nabla\log\exp(r({\bm{x}}_{t},{\bm{y}}))$
+$=a_{t}{\bm{x}}_{t}+b_{t}\nabla\left[\log p_{t}({\bm{x}}_{t}|{\bm{y}})+w\log\exp(r({\bm{x}}_{t},{\bm{y}}))\right]$
 $=a_{t}{\bm{x}}_{t}+b_{t}\nabla\log\tilde{p_{t}}({\bm{x}}_{t}|{\bm{y}})$ (18)
@@
 Reward Guidance for Rectified Flow.
-Rectified Flow (Liu et al., 2022) is also a Gaussian path defined by
+Rectified Flow Liu et al. [2022] is also a Gaussian path defined by
 ${\bm{x}}_{t}=(1-t){\bm{x}}_{0}+t{\bm{x}}_{1}$ (19)
@@
 $\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})+w[\frac{1}{1-t}{\bm{x}}_{t}+\frac{t}{1-t}\nabla r({\bm{x}}_{t},{\bm{y}})+{\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})].$ (21)
-Eq. 29 becomes
+Eq. B.2 becomes
 $\tilde{v}_{t}({\bm{x}}_{t}|{\bm{y}})={\bm{v}}_{t}({\bm{x}}_{t}|{\bm{y}})-w\frac{t}{1-t}\nabla r({\bm{x}}_{t},{\bm{y}}).$ (22)
-We use Eq. 24 or Eq. 29 to guide inference through reward model.
-Appendix B Input Template for Reward Model
-Full Input Template [VIDEO] You are tasked with evaluating a generated video based on three distinct criteria: Visual Quality, Motion Quality, and Text Alignment.
-Please provide a rating from 0 to 10 for each of the three categories, with 0 being the worst and 10 being the best.
-Each evaluation should be independent of the others.
-**Visual Quality:**
-Evaluate the overall visual quality of the video, with a focus on static factors. The following sub-dimensions should be considered:
-- **Reasonableness:** The video should not contain any significant biological or logical errors, such as abnormal body structures or nonsensical environmental setups.
-- **Clarity:** Evaluate the sharpness and visibility of the video. The image should be clear and easy to interpret, with no blurring or indistinct areas.
-- **Detail Richness:** Consider the level of detail in textures, materials, lighting, and other visual elements (e.g., hair, clothing, shadows).
-- **Aesthetic and Creativity:** Assess the artistic aspects of the video, including the color scheme, composition, atmosphere, depth of field, and the overall creative appeal.
-The scene should convey a sense of harmony and balance.
-- **Safety:** The video should not contain harmful or inappropriate content, such as political, violent, or adult material.
-If such content is present, the image quality and satisfaction score should be the lowest possible.
-Please provide the ratings of Visual Quality: <|VQ_reward|>
-END
-**Motion Quality:**
-Assess the dynamic aspects of the video, with a focus on dynamic factors. Consider the following sub-dimensions:
-- **Stability:** Evaluate the continuity and stability between frames.
-There should be no sudden, unnatural jumps, and the video should maintain stable attributes (e.g., no fluctuating colors, textures, or missing body parts).
-- **Naturalness:** The movement should align with physical laws and be realistic.
-For example, clothing should flow naturally with motion, and facial expressions should change appropriately (e.g., blinking, mouth movements).
-- **Aesthetic Quality:** The movement should be smooth and fluid.
-The transitions between different motions or camera angles should be seamless, and the overall dynamic feel should be visually pleasing.
-- **Fusion:** Ensure that elements in motion (e.g., edges of the subject, hair, clothing) blend naturally with the background, without obvious artifacts or the feeling of cut-and-paste effects.
-- **Clarity of Motion:** The video should be clear and smooth in motion. Pay attention to any areas where the video might have blurry or unsteady sections that hinder visual continuity.
-- **Amplitude:** If the video is largely static or has little movement, assign a low score for motion quality.
-Please provide the ratings of Motion Quality: <|MQ_reward|>
-END
-**Text Alignment:**
-Assess how well the video matches the textual prompt across the following sub-dimensions:
-- **Subject Relevance** Evaluate how accurately the subject(s) in the video (e.g., person, animal, object) align with the textual description.
-The subject should match the description in terms of number, appearance, and behavior.
-- **Motion Relevance:** Evaluate if the dynamic actions (e.g., gestures, posture, facial expressions like talking or blinking) align with the described prompt.
-The motion should match the prompt in terms of type, scale, and direction.
-- **Environment Relevance:** Assess whether the background and scene fit the prompt.
-This includes checking if real-world locations or scenes are accurately represented, though some stylistic adaptation is acceptable.
-- **Style Relevance:** If the prompt specifies a particular artistic or stylistic style, evaluate how well the video adheres to this style.
-- **Camera Movement Relevance:** Check if the camera movements (e.g., following the subject, focus shifts) are consistent with the expected behavior from the prompt.
-Textual prompt - [PROMPT]
-Please provide the ratings of Text Alignment: <|TA_reward|>
-END
-Appendix C Details of Human Annotation
+We use Eq. 17 or Eq. B.2 to guide inference through reward model.
+Appendix C Pseudo-code of Flow-DPO and Flow-NRG
+The Pytorch-style implementation of the Flow-DPO loss (Eq. 4.2) is shown below:
+def loss(model, ref_model, x_w, x_l, c, beta):
+# model: Flow model that takes prompt condition c and timestep as inputs and predicts velocity
+# ref_model: Frozen initialization of the model
+# x_w: Preferred Image (latents in this work)
+# x_l: Non-Preferred Image (latents in this work)
+# c: Conditioning (text in this work)
+# beta: Regularization Parameter
+timestep = torch.rand(len(x_w))
+noise = torch.randn_like(x_w)
+noisy_x_w = (1 - timestep) * x_w + timestep * noise
+noisy_x_l = (1 - timestep) * x_l + timestep * noise
+velocity_w_pred = model(noisy_x_w, c, timestep)
+velocity_l_pred = model(noisy_x_l, c, timestep)
+velocity_ref_w_pred = ref_model(noisy_x_w, c, timestep)
+velocity_ref_l_pred = ref_model(noisy_x_l, c, timestep)
+velocity_w = noise - x_w
+velocity_l = noise - x_l
+model_w_err = (velocity_w_pred - velocity_w).norm().pow(2)
+model_l_err = (velocity_l_pred - velocity_l).norm().pow(2)
+ref_w_err = (velocity_ref_w_pred - velocity_w).norm().pow(2)
+ref_l_err = (velocity_ref_l_pred - velocity_l).norm().pow(2)
+w_diff = model_w_err - ref_w_err
+l_diff = model_l_err - ref_l_err
+inside_term = -0.5 * beta * (w_diff - l_diff)
+loss = -1 * log(sigmoid(inside_term))
+return loss
+The Pytorch-style implementation of the reward guidance (Eq. 11) is shown below:
+def reward_guidance(model, reward_model, prompt_embeds, latents, timesteps, reward_weight, rg_scale, cfg_scale):
+# model: Flow model that predicts velocity given latents and conditions
+# reward_model: Model that evaluates the quality of latents based on prompt embeddings and timestep
+# prompt_embeds: Embeddings of the text prompts
+# latents: Initial noise
+# timesteps: Sequence of timesteps
+# reward_weight: weighting coefficient of multi-dimensional rewards
+# rg_scale: scale factor for reward guidance
+# cfg_scale: scale factor for classifier free guidance
+dts = timesteps[:-1] - timesteps[1:]
+for i, t in enumerate(timesteps):
+v_pred = model(latents, prompt_embeds, t)
+if cfg_scale != 1.0:
+v_pred_uncond = model(latents, None, t) # unconditional prediction
+v_pred = v_pred_uncond + cfg_scale * (v_pred - v_pred_uncond)
+latents = latents.detach().requires_grad_(True)
+reward = reward_model(latents, prompt_embeds, t)
+reward = (reward * reward_weight).sum()
+reward_guidance = torch.autograd.grad(reward, latents)
+if t != 1:
+v_pred = v_pred - rg_scale * t / (1 - t) * reward_guidance
+latents = latents - dts[i] * v_pred
+return latents
+Appendix D Architecture of Internal Video Diffusion Model
+Our work employs an internal text-to-video foundation model, which is a Transformer-based latent diffusion model Peebles and Xie [2023] with 1B parameters.
+The model integrates a 3D VAE to encode video data into a latent space, alongside a Transformer-based video diffusion model.
+Each Transformer block is instantiated as a sequence of 2D spatial-attention, 3D self-attention, cross-attention, and FFN modules.
+Importantly, the model is trained under Flow Matching framework.
+An illustration of the model architecture is provided in Fig 7.
+Figure 7: Architecture of Our Internal Video Latent Diffusion Model Backbone.
+Appendix E Ablation Study of Our Reward Model
+Ablation of VideoReward
+To further validate the key design choices of our reward model discussed in Sec.3.2, we conduct an ablation study on the evaluation benchmark, providing a quantitative supplement to our analysis in Tab.
+6.
+We compare three reward model variants: regression-style, Bradley-Terry, and Bradley-Terry-With-Tied.
+The BT model slightly outperforms the regression-style model, likely due to the advantages of pairwise annotations, which better capture ranking relationships and are more robust to annotation noise.
+The BTT model matches the BT model on ties-excluded accuracy but significantly improves ties-included accuracy, as its explicit handling of tied pairs helps it learn a more robust decision boundary, capturing neutral relationships in ambiguous cases.
+Additionally, we find that using separate tokens for each reward attribute, instead of a shared last token further improves performance.
+This design better represents distinct reward aspects, enhancing alignment with human preferences.
+Table 6: Ablation study on reward model type and seprate tokens. Bold: Best Performance.
+Variants RM Type Separate Tokens GenAI-Bench VideoGen-RewardBench
+Overall Accuracy Overall Accuracy VQ Accuracy MQ Accuracy TA Accuracy
+w/ Ties w/o Ties w/ Ties w/o Ties w/ Ties w/o Ties w/ Ties w/o Ties w/ Ties w/o Ties
+I Regression 48.28 71.13 58.39 70.16 54.23 73.61 61.16 65.56 52.60 70.95
+II BT 47.74 71.21 61.22 72.58 52.33 77.10 59.43 73.50 53.06 71.62
+III BTT 48.27 70.89 61.50 73.39 60.52 76.31 64.64 72.40 53.55 72.12
+IV BTT 49.41 72.89 61.26 73.59 59.68 75.66 66.03 74.70 53.80 72.20
+Ablation on Noisy-Latent Reward Training.
+We apply reward guidance using only MQ rewards on TA-Hard prompts.
+When trained with noised latents, the reward model successfully guides generation, yielding an MQ win rate of 74.6.
+By contrast, the model trained on clean latents fails to offer meaningful guidance for intermediate noised latents, achieving only a 38.6 win rate and underperforming even the unguided baseline.
+Following Yu et al.
+[2023], we also try directly predicting ${\bm{x}}_{0}$ from ${\bm{x}}_{t}$ and then backpropagating via $r({\bm{x}}_{0},y)$.
+However, this method produces unusable videos without meaningful content.
+Appendix F Statistics of Training Data
+In Tab. 1 and Fig. 8, we provide a comprehensive statistics of our 182k-sized training dataset.
+Figure 8: Statistics of our training data.
+Previous reward models (e.g., VideoScore He et al.
+[2024], LIFT Wang et al.
+[2024b]) were trained on datasets with mostly low resolutions ($\leq$480p) and short durations (2s), whereas ours include recent T2V models with higher resolutions ($\geq$720p) and longer clips (3–6s).
+The dataset statistics are shown in Tab.
+7.
+Table 7: Comparison of video preference datasets in terms of spatial resolution and temporal duration.
+Dataset Resolution Duration
+Existing Preference Datasets
+VideoFeedback He et al. [2024] $256\times 256$$576\times 1024$ 1s – 3s
+VideoDPO Liu et al. [2024b] $320\times 512$ 2s
+LiFT-HRA Wang et al. [2024b] $480\times 720$ 6s
+Our Datasets
+Ours $384\times 672$$768\times 1408$ 3s – 6s
+Appendix G Details of Human Annotation
 We provide additional details regarding the annotation process.
@@
 - Camera Movement Relevance: Relevance to the camera descriptions, if exists.
-Appendix D Details of Reward Model Evaluation
-D.1 Evaluation Benchmarks
-We use two benchmarks to evaluate the performance of our reward model: GenAI-Bench (Jiang et al., 2024b) and VideoGen-RewardBench.
+Appendix H Details of Reward Model Evaluation
+H.1 Evaluation Benchmarks
+We evaluate our reward model using two benchmarks: GenAI-Bench Jiang et al.
+[2024b] and VideoGen-RewardBench.
 GenAI-Bench is employed to assess the accuracy of the reward model in evaluating pre-SOTA-era T2V models, while VideoGen-RewardBench is used to evaluate its performance on modern T2V models.
-In this subsection, we provide a detailed description of the two benchmarks, highlighting their key parameters and differences in Fig 8 and Tab.
+In this subsection, we describe both benchmarks, highlighting key parameters and differences in Fig 9 and Tab.
 9.
@@
 VideoGen-RewardBench
-VideoGen-Eval (Zeng et al., 2024) has open-sourced a dataset containing videos generated by 9 closed-source and 3 open-source models, designed to qualitatively visualize performance differences across models.
+VideoGen-Eval Zeng et al.
+[2024] has open-sourced a dataset containing videos generated by 9 closed-source and 3 open-source models, designed to qualitatively visualize performance differences across models.
 Due to its high-quality data, broad coverage of the latest advancements in T2V models, and third-party nature, we leverage VideoGen-Eval to create a fair benchmark, VideoGen-RewardBench, for evaluating reward models’ performance on modern T2V models.
@@
 We use VideoGen-RewardBench as the primary benchmark to evaluate reward models’ performance on modern T2V models.
-D.2 Comparison Methods
+H.2 Comparison Methods
 Random
-[…]
+To eliminate the influence of metric calculations and benchmark distributions on our evaluation results, we introduce a special baseline: random scores.
+Specifically, for each triplet (prompt, video A, video B), we randomly sample $r_{A}$ and $r_{B}$ from a standard normal distribution, denoted as $r_{A},r_{B}\sim\mathcal{N}(0,1)$.
+We then calculate accuracy in the same manner as for the other models.
+The mathematical expectation of random scores for ties-excluded accuracy is $\mathbb{E}(acc)=\frac{1}{2}$, and the mathematical expectation of ties-included accuracy is $\mathbb{E}(acc)=\max(\frac{1}{3},p(c=\text{"Ties"}))$.
 VideoScore
-VideoScore (He et al., 2024) adopts Mantis-Idefics2-8B (Jiang et al., 2024a) as its base model and trains with point-wise data using MSE loss to model human preference scores.
+VideoScore He et al.
+[2024] adopts Mantis-Idefics2-8B Jiang et al.
+[2024a] as its base model and trains with point-wise data using MSE loss to model human preference scores.
 Since VideoScore predicts scores across multiple dimensions, and its dimension definitions differ from those in VideoGen-RewardBench, we compute both the overall accuracy and the dimension-specific (VQ, MQ, TA) accuracy by averaging the scores of five dimensions when conduct evaluation GenAI-Bench and VideoGen-RewardBench, consistent with the evaluation strategy outlined in their paper.
@@
 LiFT
-LiFT (Wang et al., 2024b) adopts VILA-1.5-40B (Lin et al., 2024a) as its base model and employs a VLM-as-a-judge approach.
+LiFT Wang et al.
+[2024b] adopts VILA-1.5-40B Lin et al.
+[2024a] as its base model and employs a VLM-as-a-judge approach.
 The reward model is trained through instruction tuning with inputs, preference scores along with a critic.
@@
 VisionReward
-VisionReward (Xu et al., 2024a) adopts CogVLM2-Video-12B(Hong et al., 2024) as the base model and is trained to answer a set of judgment questions about the video with a binary ”yes” or ”no” response using cross-entropy loss.
+VisionReward Xu et al.
+[2024a] adopts CogVLM2-Video-12BHong et al.
+[2024] as the base model and is trained to answer a set of judgment questions about the video with a binary ”yes” or ”no” response using cross-entropy loss.
 During inference, VisionReward evaluates 64 checklist items, providing converted into 1/0 scores.
@@
 We use the final score to calculate both the overall accuracy and the VQ/MQ/TA accuracy.
-VisionReward’s training data includes models from the pre-SOTA era models (Chen et al., 2024) as well as recent open-source T2V models (Zheng et al., 2024; Yang et al., 2024b).
+VisionReward’s training data includes models from the pre-SOTA era models Chen et al.
+[2024b] as well as recent open-source T2V models Zheng et al.
+[2024], Yang et al.
+[2024b].
 It performs well on GenAI-Bench and demonstrates reasonable capabilities on VideoGen-RewardBench.
 Our Reward Model
-We adopts QWen2-VL-2B (Wang et al., 2024a) as the base model and train it with pair-wise data using BTT loss in Eq.
-6.
+We adopts QWen2-VL-2B Wang et al.
+[2024a] as the base model and train it with pair-wise data using BTT loss in Eq.
+2.
 Scores are normalized on the validation set and averaged to obtain overall scores for evaluation and optimization.
-When evaluating on VideoGen-RewardBench, we sample videos at 2 FPS and a resolution of 448$\times$448, consistent with the training settings.
+When evaluating on VideoGen-RewardBench, we sample videos at 2 FPS and a resolution of $\times448$, consistent with the training settings.
 We calculate the overall accuracy by averaging the scores across the three dimensions, and compute dimension-specific accuracies using the respective scores.
-For GenAI-Bench, we sample videos at 2 FPS and a resolution of 256$\times$256, as the minimum resolution in GenAI-Bench is 256$\times$256.
+For GenAI-Bench, we sample videos at 2 FPS and a resolution of $\times256$, as the minimum resolution in GenAI-Bench is $\times256$.
 Given the significant disparities in visual quality and motion between the GenAI-Bench videos and our training data, we utilize only the predicted TA scores to calculate the overall score.
-D.3 Evaluation Metrics
-Similarly to VisionReward (Xu et al., 2024a), we report two accuracy metrics: ties-included accuracy (Deutsch et al., 2023) and ties-excluded accuracy.
+H.3 Evaluation Metrics
+Similarly to VisionReward Xu et al.
+[2024a], we report two accuracy metrics: ties-included accuracy Deutsch et al.
+[2023] and ties-excluded accuracy.
 For ties-excluded accuracy, we exclude all data labeled as ”ties” and use only data labeled as ”A wins” or ”B wins” for calculation.
 Since all competitors predict scores based on pointwise samples, we compute the rewards for each pair, convert the relative reward relationships into binary labels, and calculate classification accuracy.
-For ties-included accuracy, we adopt the $O(n^{2}\log n)$ algorithm proposed in Algorithm 1 by Deutsch et al.
-(2023).
+For ties-included accuracy, we adopt the tie calibration algorithm proposed in Algorithm 1 by Deutsch et al.
+[2023].
 This method traverses all possible tie thresholds, calculates three-class accuracy for each threshold, and selects the highest accuracy as the final metric.
-Appendix E Pseudo-code of Flow-DPO and Flow-NRG
-The Pytorch-style implementation of the Flow-DPO loss (Eq. 4.1) is shown below:
-def loss(model, ref_model, x_w, x_l, c, beta):
-# model: Flow model that takes prompt condition c and timestep as inputs and predicts velocity
-# ref_model: Frozen initialization of the model
-# x_w: Preferred Image (latents in this work)
-# x_l: Non-Preferred Image (latents in this work)
-# c: Conditioning (text in this work)
-# beta: Regularization Parameter
-# returns: DPO loss value
-timestep = torch.rand(len(x_w))
-noise = torch.randn_like(x_w)
-noisy_x_w = (1 - timestep) * x_w + timestep * noise
-noisy_x_l = (1 - timestep) * x_l + timestep * noise
-velocity_w_pred = model(noisy_x_w, c, timestep)
-velocity_l_pred = model(noisy_x_l, c, timestep)
-velocity_ref_w_pred = ref_model(noisy_x_w, c, timestep)
-velocity_ref_l_pred = ref_model(noisy_x_l, c, timestep)
-velocity_w = noise - x_w
-velocity_l = noise - x_l
-model_w_err = (velocity_w_pred - velocity_w).norm().pow(2)
-model_l_err = (velocity_l_pred - velocity_l).norm().pow(2)
-ref_w_err = (velocity_ref_w_pred - velocity_w).norm().pow(2)
-ref_l_err = (velocity_ref_l_pred - velocity_l).norm().pow(2)
-w_diff = model_w_err - ref_w_err
-l_diff = model_l_err - ref_l_err
-inside_term = -0.5 * beta * (w_diff - l_diff)
-loss = -1 * log(sigmoid(inside_term))
-return loss
-The Pytorch-style implementation of the reward guidance (Eq. 15) is shown below:
-def guidance(model, reward_model, prompt_embeds, latents, timesteps, reward_weight):
-# model: Flow model that predicts velocity given latents and conditions
-# reward_model: Model that evaluates the quality of latents based on prompt embeddings and timestep
-# prompt_embeds: Embeddings of the text prompts
-# latents: Initial noise
-# timesteps: Sequence of timesteps
-# reward_weight: weighting coefficient of multi-dimensional rewards
-# guidance_scale: scale factor for reward guidance
-dts = timesteps[:-1] - timesteps[1:]
-for i, t in enumerate(timesteps):
-v_pred = model(latents, prompt_embeds, t)
-latents = latents.detach().requires_grad_(True)
-reward = reward_model(latents, prompt_embeds, t)
-reward = reward * reward_weight
-reward_guidance = torch.autograd.grad(reward, latents)
-v_pred = v_pred - guidance_scale * t / (1 - t) * reward_guidance
-latents = latents - dts[i] * v_pred
-return latents
-Appendix F Hyperparameters
+Appendix I Hyperparameters
+In all alignment experiments, we applied LoRA to fine-tune the transformer models’ linear layers, as our findings indicate that full parameter fine-tuning can degrade the model’s performance or potentially lead to model collapse.
+Table 10: Hyperparameters for alignment algorithms
 Algorithm-agnostic hyperparameters for SFT, Flow-RWR, Flow-DPO
-Training strategy LoRA (Hu et al., 2021)
+Training strategy LoRA [Hu et al., 2021]
 LoRA alpha 128
@@
 LoRA target-modules q_proj,k_proj,v_proj,o_proj
-Optimizer Adam (Kingma, 2014)
+Optimizer Adam [Kingma, 2014]
 Learning rate 5e-6
@@
 Batch size 64
+GPUs 16 NVIDIA A800@80G
 Flow-DPO
 $\beta$ 500
-Table 10: Hyperparameters for Alignment Algorithms
+Table 11: Hyperparameters for reward modeling.
 VLM
@@
 LoRA target-modules Linear layers in language model
-Optimizer Adam (Kingma, 2014)
+Optimizer Adam [Kingma, 2014]
 Learning rate 2e-6
@@
 Batch size 32
+GPUs 8 NVIDIA A800@80G
 $\theta$ in Eq. 2 5.0
@@
 Training strategy Full training
-Optimizer Adam (Kingma, 2014)
+Optimizer Adam [Kingma, 2014]
 Learning rate 5e-6
@@
 Reward Dimension 3
-Table 11: Hyperparameters for reward modeling.
-Appendix G Extended Experimental Results
-Figure 10: Human evaluation of Flow-DPO on TA-Hard prompt.
-Appendix H Prompt Subset of TA-Hard
+GPUs 8 NVIDIA A800@80G
+Appendix J Additional Qualitative Results
+We present additional qualitative results generated by both the original model and the Flow-DPO aligned model, as shown in Fig.11 and Fig.12.
+Figure 11: Additional visual comparison of videos generated by the original model and the Flow-DPO aligned model.
+Figure 12: Additional visual comparison of videos generated by the original model and the Flow-DPO aligned model.
+Appendix K Input Template for Reward Model
+Full Input Template [VIDEO] You are tasked with evaluating a generated video based on three distinct criteria: Visual Quality, Motion Quality, and Text Alignment.
+Please provide a rating from 0 to 10 for each of the three categories, with 0 being the worst and 10 being the best.
+Each evaluation should be independent of the others.
+**Visual Quality:** Evaluate the overall visual quality of the video, with a focus on static factors.
+The following sub-dimensions should be considered: - **Reasonableness:** The video should not contain any significant biological or logical errors, such as abnormal body structures or nonsensical environmental setups.
+- **Clarity:** Evaluate the sharpness and visibility of the video.
+The image should be clear and easy to interpret, with no blurring or indistinct areas.
+- **Detail Richness:** Consider the level of detail in textures, materials, lighting, and other visual elements (e.g., hair, clothing, shadows).
+- **Aesthetic and Creativity:** Assess the artistic aspects of the video, including the color scheme, composition, atmosphere, depth of field, and the overall creative appeal.
+The scene should convey a sense of harmony and balance.
+- **Safety:** The video should not contain harmful or inappropriate content, such as political, violent, or adult material.
+If such content is present, the image quality and satisfaction score should be the lowest possible.
+Please provide the ratings of Visual Quality: <|VQ_reward|> END **Motion Quality:** Assess the dynamic aspects of the video, with a focus on dynamic factors.
+Consider the following sub-dimensions: - **Stability:** Evaluate the continuity and stability between frames.
+There should be no sudden, unnatural jumps, and the video should maintain stable attributes (e.g., no fluctuating colors, textures, or missing body parts).
+- **Naturalness:** The movement should align with physical laws and be realistic.
+For example, clothing should flow naturally with motion, and facial expressions should change appropriately (e.g., blinking, mouth movements).
+- **Aesthetic Quality:** The movement should be smooth and fluid.
+The transitions between different motions or camera angles should be seamless, and the overall dynamic feel should be visually pleasing.
+- **Fusion:** Ensure that elements in motion (e.g., edges of the subject, hair, clothing) blend naturally with the background, without obvious artifacts or the feeling of cut-and-paste effects.
+- **Clarity of Motion:** The video should be clear and smooth in motion.
+Pay attention to any areas where the video might have blurry or unsteady sections that hinder visual continuity.
+- **Amplitude:** If the video is largely static or has little movement, assign a low score for motion quality.
+Please provide the ratings of Motion Quality: <|MQ_reward|> END **Text Alignment:** Assess how well the video matches the textual prompt across the following sub-dimensions: - **Subject Relevance** Evaluate how accurately the subject(s) in the video (e.g., person, animal, object) align with the textual description.
+The subject should match the description in terms of number, appearance, and behavior.
+- **Motion Relevance:** Evaluate if the dynamic actions (e.g., gestures, posture, facial expressions like talking or blinking) align with the described prompt.
+The motion should match the prompt in terms of type, scale, and direction.
+- **Environment Relevance:** Assess whether the background and scene fit the prompt.
+This includes checking if real-world locations or scenes are accurately represented, though some stylistic adaptation is acceptable.
+- **Style Relevance:** If the prompt specifies a particular artistic or stylistic style, evaluate how well the video adheres to this style.
+- **Camera Movement Relevance:** Check if the camera movements (e.g., following the subject, focus shifts) are consistent with the expected behavior from the prompt.
+Textual prompt - [PROMPT] Please provide the ratings of Text Alignment: <|TA_reward|> END
+Appendix L Prompt Subset of TA-Hard
 A rabbit and a turtle racing on a track.
@@
 Spectators are cheering from the sidelines, and a finish line is visible in the distance.
-A lion and a zebra playing soccer on a grassy field. The lion is dribbling the ball, while the zebra is trying to block it. The field is surrounded by trees, and other animals are watching the game.
+A lion and a zebra playing soccer on a grassy field.
+The lion is dribbling the ball, while the zebra is trying to block it.
+The field is surrounded by trees, and other animals are watching the game.
 A fox and an owl stargazing together on a hilltop.
@@
 The court is surrounded by graffiti-covered walls and cheering spectators.
-A paper airplane with a scarf and a paper boat with a captain’s hat racing in the rain. The airplane glides through the air while the boat sails through puddles.
-A basketball with a mohawk and a soccer ball with a bandana playing hopscotch in a playground. The basketball bounces high while the soccer ball rolls smoothly.
+A paper airplane with a scarf and a paper boat with a captain’s hat racing in the rain.
+The airplane glides through the air while the boat sails through puddles.
+A basketball with a mohawk and a soccer ball with a bandana playing hopscotch in a playground.
+The basketball bounces high while the soccer ball rolls smoothly.
 A mechanical knight with steam-powered joints standing guard at an ancient castle gate.